A Complete Bibliography of the Journal of Statistical Physics: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

22 March 2023 Version 1.11

Title word cross-reference

(∇ + Δ) [CDH21]. 0 [MPDH20]. 1 [DLDG21, DY21, KZ20, Mol20]. 2 [AHP20, AOdLHT20, BC20, Sam20a]. 3 [DY21, Iin21]. _{κ} [FLM22]. A + B → 2A [Jun20a]. α [GAD20, ZD21]. β < 3 [KM21]. CβE [Ass22]. d [TK21]. D_{l} [Hik22]. Δ [BdBBR23]. d ≥ 3 [CCM20]. d ≥ 5 [CN23]. Γ [BBC^{+}20, Mor20a]. H [Oto23]. K [Cao21, FP21, GL23a, Li21]. L^{2} [ADT21]. L^{2} ∩ L^{∞} [AMSY23]. (d + 1) [MGPCA22]. Z^{2} [JRA20]. Z^{d} [Mas22, Yak21]. U [Bla21]. S^{1} [LS22b]. N [MW20b]. ∇ϕ [AT21]. O(n) [AS21a]. p [BCS22, BH20, Eld20, Fac21a, Fac21b, HB21, Hik22, KV20]. Φ [JK23]. ϕ^{4} [AS21a]. Φ_{3}^{4} [HS22]. R [DSS20, FGR20]. σ [Cra21]. SU(n) [TK21]. Z [DSS20].

-Averaging [Cao21]. -Cell [MPDH20]. -Contest [KV20]. -Convergence [BBC^{+}20, Mor20a]. -D [KZ20, Mol20]. -Dimensional [DLDG21, TK21]. -Entropies [JK23]. -FPUT [GAD20]. -Hypocoercivity [ADT21]. -Ising [Bla21]. -Mixing [GL23a]. -Model [CDH21]. -Models [Cra21]. -Particle [MW20b]. -Partite [Li21]. -Positivity [FGR20]. -Processes [FP21]. -Regular [BdBBR23]. -Spin [BCS22, BH20, Eld20, Fac21a, Fac21b, HB21, Hik22]. -Stable [ZD21]. -State [Iin21]. -Symmetric [AS21a]. -Theorem [Oto23]. -Voter [Bla21].

2-Spin [SS21]. 2D [FPV22, Gao22, KR22b].

63 [GHL20].

Abelian [EN21, KW20]. Abrupt [MQ20]. Absence [WOK20]. Absolute [DGZ20]. Absorbing [GN21]. Absorption [LS20d]. Abstract [BTV21, Fia20, XX20]. Acceleration [GL22b]. Accommodation [KBK23]. Actions [RJS22, TYM23, ZM21]. Activated [LS21]. Activity [SND^{+}20]. Additive [LLV22]. Adiabatic [JM23]. Adsorption [Mad23, PD22, WLB^{+}20]. Advection [DHL20]. Adversarial [BKN22]. Affine [FK21]. Affinity [RCF20, YQ21]. Agent [Cao21]. Agent-Based [Cao21]. Agents [MMR20]. Aggregates [CG22, Cha23]. Aggregation [HP20b, TR22]. Algorithm [Bód20]. Algorithmic [ES22]. Algorithms [MGM23]. Alignment [ABT23]. Almeida [MW22]. Almost [DHS23, LLV22, PV23]. Along [IOVW20]. Alternating [BKN20]. Ampère [CDW21]. Analysis [AKL20, FKSS20, GMW20, HJ21, LJ20, NP21, SN22, Wan22, WZN20]. Analytical [GMW20]. Ancestor [DM22]. Andersen [Sha20]. Anderson [AC21, Chu23, CSZ21, DK23a, GRM20, SM23a, dHW22]. Andrea [Pel21]. Angelo [Pel21]. Angular [JKNV23, WFKM^{+}20]. Anharmonic [Men20]. Animals [FvdH21, FvdH22]. Anisotropic [BN22]. Anisotropically [BPR22]. Annealing [ABCM20, GLV21]. Annihilation [KZ20]. Ant [EFR23]. Anti [BdBBR23]. Anti-ferromagnetic [BdBBR23]. Antiquadrupolar [Lim20]. Antiquadrupolar-Disordered [Lim20]. Application [CCS20, CB20, GJ23, TCND20]. Applications [AT21, KZ20, LT21, Li21, RJS22]. Approach [ABvSY21, AKL20, BBC^{+}20, BGM20, CWZ20, CLM20, Eld20, Iin21, KG23, KR22a, Lau22, LS20a, MPS21, NPT20, WM22]. Approximate [GL23b, NPV23]. Approximating [GGM21]. Approximation [CWZ20, DMO20, FW21, HJZ20, HS20a, Kut23, LS20a, Son20, TD20]. Arctic [DDG20]. Area [GR22]. Argument [AMSY23]. Arnold [MSG^{+}20]. ASEP [LT23, XU22]. Askin [WM22]. Assignment [CES21]. Assimilation [CCH^{+}20b, TCVB20]. Assist [DTZ22]. Assisted [Mou23]. Associated [PR20, TT21]. Association [JKG21]. Associative [YK20]. Asymmetric [BBK^{+}23]. Asymmetry [Pen20]. Asymptotic [CD20b, FLM22, GMV20, Haf20, HLN22, KdS21, KD20]. Asymptotics [ADT21, BN22, BEL22, BS20d, HDS22, LP22, MJB23]. Athermal [DAR22]. Attachment [GLY21, RCF20]. Attracting [Ara21]. Attraction [Sam20a]. Attractor [Ara21, PMACF20]. Automata [DAT21, Har22]. Autonomous [CLM20]. Autoregressive [AFKH^{+}20]. Avalanche [HXX23]. Average [DHS23]. Averaged [AOdLHT20, DHL20, RB20]. Averages [BHL23, KL20a]. Averaging [Cao21, CWZ20, CCHS22, Gao20, Gao22, GJJ21, RV21, XX20]. avoiding [LP22]. Aw [DT20, DTZ22].

Background [FL23]. Backward [Ngu22]. Bak [GS22, KV21, Vol22]. Balance [MM20a]. Balancing [RF21, WBC21]. Balescu [FB22]. Ball [KL20b]. Ballistic [KZ20]. Ballot [GL22c]. Band [FM21, HK22, Jan22, Shc20, Tan20]. Bands [TK21]. Barotropic [FF20]. Barrier [Bód20]. Barycentric [KV21]. Based [Cao21, GLY21, GT20, Meo23, SN22, YS21]. Bayes [LLV22]. BCS [HL22]. Becker [PN23]. BEG [Lim20]. Behavior [AAVF23, DFP20, Dem23, GZ20, KdS21, Phi20, Sak21, TH22, WM22]. Behaviors [GK22b, HP20a, HPRS20]. Behaviour [AM21, KBK23, MJB23]. Below [FB21, NIB21]. Bernoulli [MPR20]. Bernstein [CDW21]. Berry [NPV23, PV20, TD20]. Beta [NT20, TT21]. Bethe [DK23a, KTT22]. Between [BS23, BPP20, Der23, Fac21a, Fac21b, JPR21, JR21, Wre22, YK20]. Beyond [EGN20, Pel21]. BGK [BY20, HHW21, KLY21, LS20b, MM20b, MW20b, jSY23]. Bhatnagar [Oto23]. Biased [MMR20]. Bibliometric [WZN20]. Bidirectional [JVG21]. Bifurcation [TCDN20]. Bifurcations [Rod21]. Big [GH20a]. Big-Jump [GH20a]. Billiards [Bru21, DLDG21]. Bimodal [KC21]. Binary [Các21, Zam22]. Biorthogonal [CW22]. Bipartite [Liu20, Liu21]. Bipolar [TTZ22]. Birth [MBL20]. Bivectorial [YQ21]. Black [Pel21]. Blast [KR22c, SCDK23]. Bloch [Wat20, Wat19]. Block [JLS22, KLSS20]. Blume [AAVF23, Kim21, MS22a]. Body [CLLL23, DP21b, EM22, KTT23, SEG22]. Bohm [Pig23]. Bohmian [DGT^{+}20]. Boltzmann [AS21b, ABCM20, ALT22, AMSY23, Bob23a, BGSRS20, Bou20, CDK23, CKL20, CCHS22, CS20, Den22, GM21a, HJZ20, HJ23, JKNV23, Kep23, LL23b, LSW22, MM20b, Oto23, PR20, Qi21, TTZ22, YS21]. Boltzmann-Grad [BGSRS20]. Bond [XZ21]. Boolean [CMG20]. Bootstrap [AK21, Har22]. Bordered [BEL22]. Borel [Abh23]. Bose [CCS21, FKSS20, NPT20, Pec22, Tas20]. Bosonic [FM21]. Bosons [BY20, BPPS20, CCS21, RS22]. Bound [CG20, TD20, Vol22]. Boundaries [DDG20, DHS21, KG23]. Boundary [BGJOS22, BL22, CKL20, DGT^{+}20, EGN20, FGR22, GN21, Iin21, LYZ23, NPT20, TH22]. Bounded [BY20, CCHS22, DSS20, GOS20, LLY22]. Bounds [Cra21, Har23, JLM23, KD20]. Boussinesq [AOdLHT20]. Box [CCS20, KL20b, SM23a]. Braided [Fen21]. Brain [You20]. Bramson [Der23]. Branching [BR21, BK21, EGvdHN20, FGR20, GS22, Kut23, LS20d, PKTM23]. Breaking [AAAB22, BC22, IISS23, Tas20]. Bridges [GMS21, GMS23]. Broglie [Pig23]. Brownian [GMS23, ACR21, EKN20, GS20, GMS21, LLX20, LS20d, NIB21, PPR22, PT20]. BTW [SSS21]. Bulk [CCH20a]. Burgers [Fac21a, Fac21b]. Burnett [Bob22].

Cahn [Mor20a]. Calculation [MGPCA22]. Calculus [CM20a, CM20b]. Campaign [Nei21]. Cane [TCND20]. Canonical [BB22, Hab23, SM23b, ZCD22]. Capacities [BGM20]. Capacity [TYM23, ZM21]. Capel [AAVF23, Kim21, MS22a]. Captive [MM23]. Carbon [SN22]. Carlo [ADS^{+}22, FK21, MGM23, WM22]. Case [CLM20, CW22, Pie20]. Casimir [NPT20]. Catalan [MSU20]. Categories [Fen21, Hep20]. Cauchy [HJ23]. Causation [AFKH^{+}20]. Cayley [JRS20, RRA23, Roz22]. Cell [KS21, MPDH20, PN23, Tan20]. Cell-Construction [Tan20]. Cells [FK20]. Cellular [DAT21, Har22]. Cencini [Pel21]. Central [ABC23, CGH22, Czu22, EM22]. Certain [TYM23, ZM21]. CFT [Iin21]. Chain [ADS^{+}22, BCD22, DDD^{+}20, GAD20, Gol23, GKM21, LS22a, Men20, MGM23, Spo20]. Chains [BKN20, CK22, DHO20, FGK20, GH20b, GL20, HNK22, IOS21, Jon22, LS23, MB21, NR21, SBB^{+}21]. Change [FvMST23, JLR23, MQ20]. Channels [LMAC20, ZL21]. Chaos [Cao21, CT21, Cri23, Eld20, Gol23, GM21b]. Chaotic [Gal21]. Chapman [GPRR20]. Characterisation [MGM23]. Characteristic [Shc20]. Characteristics [FN20]. Characterization [TR22]. Characters [LS23]. Charge [CN21, HKR22]. Charged [Sam20a]. Cheaper [CGP20a]. Chemical [GL22a, Kim21, MM20a, MC23]. Chemical-Reaction [MC23]. Chemically [KLY21]. Chernikov [PR20]. Chimeras [MM22]. Choice [GLY21, PR20]. Circuit [HSS20]. Circular [WLB^{+}20]. Circulation [CCH^{+}20b]. Circulation-Preserving [CCH^{+}20b]. Class [Bau20, CD20a, GH20a, HKR17, HKR23b, LT23, LXZ23a, LXZ23b, RNA23, XU22]. Classical [DDD^{+}20, GV20, JLM23, LMAC20, Luc20b, NF20b, PR20, Rue20, Spo20, ZW22, vEKM20, MB21]. Clausius [WQ20]. Climate [AOdLHT20, AvdH20, Luc20a, Pie20, RB20, TBD^{+}20, WFKM^{+}20]. Cliques [HDS22]. Close [DLDG21]. Closure [CLM20]. CLT [Jan22]. Cluster [JK22, KK21, NF20b, Sco21]. Clustering [BS20d, FL18, vdHvdHM23, FL20a]. Clusters [BDNS20, KD20, MM21]. Coagulation [FLNV23]. Cocycles [DHS23]. Coefficient [Urr22]. Collapsed [BO21, LP22]. Collective [ABT23]. Collision [ABT23, HHW21]. Collisions [HKR23a]. Color [FL18, FL20a]. Combined [CLL21]. Combustion [JMPR23]. Comment [GMS23]. Comments [Các21]. Common [CLTC15]. Community [Li21]. commutative [CM20b, CM20a]. Compact [Pap20, vGR20, FGK20, TYM23, ZM21]. Comparison [JPR21, JR21, Jun20b]. Competing [RRA23]. Competition [BS23]. Competitive [AM21]. Complete [EGvdHN20, MGPCA22, SLZ23]. Completely [HKP21]. Complexity [DZW^{+}20, SEG22, VLC23]. Component [WLB^{+}20]. Compressible [FH22]. Computation [BR22c, CFVY20, CGT21, RB20, Unt22]. Computational [SN22]. Computing [BPC22]. Concavity [Wre22]. Concentration [BCM20, CMRU20, CCR21]. Condensates [Tas20]. Condensation [ACR21, CCS21, God21, NPT20, Pec22, Yua20]. Condensed [JCG20]. Condition [CKL20, GN21, Hut22]. Conditional [SD21]. Conditioned [BR21, CLV22]. Conditions [DS20, DGT^{+}20, JK22, LYZ23, NPT20]. Conductance [AT21]. conduction [GPRR20]. Conductivities [WLO20]. Conductivity [GMP20]. Confetti [GR22]. Configuration [LT23]. Configurations [BGN21]. Confined [DG20, FPT21]. Conformal [KO20]. Conjectures [DSS20]. Connected [Mol22]. Consciousness [Hep20]. Consensus [LL22]. Conservation [BPP20, EN21, Oto23]. Conservative [Lep23, RF21, EN21]. Consistent [HHW21]. Constant [ZXD22, BC20]. Constraint [BS20d, CDS21]. Constraints [KY23]. Construction [LYZ23, Tan20]. Contact [BL22, CDLP23, XU22]. Contest [KV20]. Continued [AS21a, Abh23]. Continuity [Moh22]. Continuous [HM23]. Continuous-Time [HM23]. Continuum [FK21]. Contracting [Ara21]. Contraction [SS21]. Control [ZL22]. Convective [YS21]. Conventional [BS23]. Convergence [AS21a, AESW21, BBC^{+}20, Cao21, DG23, FP21, FLM22, JK22, KV20, LS22a, LL23b, LLV22, Men20, Mor20a, NF20a, NF20b, NPV23, Son20, Sub22, Yua20]. Convex [CCHS22, GM21b]. Cooperative [JVG21]. Core [KK20b]. Corners [PT20]. Corrected [EK20a]. Correction [BS20b, BR22a, Các20a, CM20a, CLTC23, Fac21a, FvdH22, FL20a, HKR23b, HR23, HKR23a, HKR23c, JT21, Jan21, LXZ23a, MC21, RZ22, TN21, Wat20]. Corrections [BPPS20, Wre21]. Correlation [AFKH^{+}20, Ciu22, GKM21, ORD22, Rue20, SS21]. Correlations [AHP20, BEL22, BGH20, Cra21, DAR22, GG22, Sam20b, Sos23, XTL20]. Correspondence [KKS23]. Corridored [MM23]. Cost [BPC22]. Coulomb [ABE23, CHR22, FL23]. Countable [Pin23, Tak20, Tak23]. Counterexamples [Czu22]. Counterions [Sam20a]. Counterpart [Luc20b]. Counting [CGT21, DIM23, THFH20]. Coupled [EGK21, ST22, Tas20, TCVB20, TZQY20]. Coupling [AS21a, CDS21, CK20, KL20a, LS20c, MS22b, WY22]. Couplings [HP21]. Coverage [Mou23]. Covered [GR22]. Cox [CDD20]. Crack [BDR23]. Creating [YK20]. Criterion [Fia20]. Critical [BN22, BGN21, CJN20, EGN20, GP23, GMP20, HM20, Hut22, Iin21, Jon22, Jun20a, KR22b, LS20d]. Criticality [GR22]. Cross [Der23, ORD22]. Cross-Correlation [ORD22]. Cross-Overs [Der23]. Crosscap [CK22]. Crystal [Gat23, GM21a, Rue20]. Crystallization [BKN20, FS20b]. Crystals [BS20a, SND^{+}20]. Cubes [Jan20, Jan21]. Cucker [BM20]. CUE [Sos23]. Cumulant [GGM21]. Curie [Ban20, CJN22, KM20, KM21, Lee22, LW20]. Current [CGP20b]. Currents [KR22b]. Curvature [BS20c, FvMST23]. Curves [BH20, DDG20, HB21, Hik22]. Cutoff [BHP21, Den22, HJZ20, HJ23, AMSY23, Kep23]. Cycle [CQ21, Xu21, YQ21]. Cycles [HLP^{+}20b, Kar20, Pig23]. Cyclic [FL18, FL20a]. Cylinder [GK22b, Rya21, Sam20a].

D [AHP20, AOdLHT20, BC20, DY21, KZ20, Mol20, Sam20a]. Damage [AAVF23]. Damping [HLW20]. Data [Bód20, CD20b, CCH^{+}20b, DSS20, HJ23, LJ20, LD22, TCVB20]. Davide [Pel21]. Dealer [KTT23]. Dean [KLvR20]. Debye [FB21]. Decay [AHP20, CB20, DE22, LS20b, SS21]. Decaying [LL23a]. Decoherence [ZL21]. decomposability [MSU20]. Decomposition [DK23b, GG20, iSHNY22]. Deep [ABCM20]. Defined [PYZ20]. Deformed [LL23a, SS22]. Degeneracy [BJN21]. Degenerate [BS13, BS20b, MGPCA22, Pap20]. Degree [GLY21]. Delay [AM21, XX20]. Delay-Induced [AM21]. Democratic [GZ20]. Dense [AAAB22]. Densities [BC20]. Density [CLLL23, DK23a, EK20b, GHO21, HL22, JLM23, Mor20b, Rya21, ZW23]. Dependence [KY23, PPR22]. Dependent [DS20, DGZ20, HHW21, LT20, LS20a, XX20, Zam22]. Depth [MQ20]. Derivation [Ben23, BR22c, CS20, DT20, DTZ22, Hut22, YS21]. Deriving [CLLL23]. Descendants [Kut23]. Describing [Dem23]. Description [BPPS20, Pen20, ZL22]. Detailed [DHP20, MM20a]. Detecting [GG20]. Detection [Li21, Pie20]. Determinants [BEL22]. Determination [Mou23]. Deviation [BHL23, Bha20, Bou20, BR22c, CH20, GT20, RS22, RB20, Tak23]. Deviations [AK21, BBC^{+}20, BBO21, CGH22, CGJ21, DHS21, FB21, FB22, GBE22, KL20b, SGG23, Sco21, XZ21]. Diagonal [BEL22]. Diagonalizability [MS21]. Diagram [DK23b, MW20a]. Dichotomous [TZQY20]. Differentiability [MSG^{+}20]. Differential [CT22, Cri23, Son20, ZD21, ZXJL23]. Diffuse [CKL20]. Diffusion [ABC23, Các20a, Các20b, CGJ21, CGV23, Doy22, FGR22, FK20, FW21, GJJ21, GN21, HOiS22, KTT22, LJ20, LXZ23a, LXZ23b, OMM22]. Diffusions [ALP21, CCS20, GS22, LS22a]. Diffusive [BL22, TTZ22, WPK21]. Dilute [BCM20]. Diluted [HLP^{+}20b]. Dimension [FPV22, FL18, HM20, Urr22, FL20a]. Dimensional [AAVF23, ABE23, BS20a, BH22, BOP23, Bla21, Bro22b, CDN21, CCS21, DAT21, DHO20, DLDG21, FS20b, FL20b, GL22c, KK22, MGPCA20, NP21, Pec22, Sam20b, SKM20, SH23, SSvB21, SLZ23, TK21, WM22, XU22, ZW23]. Dimensions [CN23, Cra21, JR21, KR22c, PMACF20]. Dimer [Ciu22]. Dipolar [ZRW20]. Dirac [ABL21, LL23b]. Direct [CGT21]. Directed [EFR23, GS20, GK22b, JRA20, LP22, WY22, Wu20]. Dirichlet [NPT20]. Disasters [GH20b]. Discrete [ABV22, Chu23, GGGM23, GL22c, KP20, MM20b, Vol22]. Discs [WLB^{+}20]. Disk [HNK22]. Disorder [Các21, CCM20, DS20, DK23a, GH20a, LRRS21, Mas22, NPV23, Wu20]. Disordered [DAR22, GS20, Lim20, NR21, SGNS20]. dispersion [WPK21]. Displacement [Ciu22, DAR22]. Displacements [Chu23]. Disruptive [RF21]. Dissecting [Rod21]. Dissipation [JK23, KKMS20, MGPCA20, Oto23, THFH20, WBC21, Tan22]. Dissipative [CM20a, CM20b, MGPCA22]. Distance [BHP21, CDD20, MPDH20, Sam20b, Wir22]. Distinguishing [Ste22]. Distortion [MC16, MC21]. Distribution [Ber22, GLY21, Hiu21, JLY20, KK21, LT23, Wen23, Wre21, YS21, ZXJL23]. Distributions [BGM20, CH22, MPS21, MSU20]. Divergence [LS20c, NT23]. Diverging [AS21a]. Divisible [CH22]. DLA [PYZ20]. DNA [Maj23]. do [LS21]. Does [Nis23]. Domain [CCHS22, DDG20, GL23b]. Domains [LLY22]. Döring [PN23]. Double [SCDK23]. Doubling [CMG20]. Drainage [PPR22]. Drift [KLvR20]. Drifted [PT20]. Driven [AJR21, BGJOS22, BPR22, BL22, DIM23, EKN20, FIS20, THFH20, TZQY20]. Driver [DTZ22]. Driver-Assist [DTZ22]. Droplet [Wil23]. Droplets [BN22]. Drude [Mas22]. Dual [Wir22]. Duality [BBK^{+}23, Fac21a, Fac21b]. Duffing [DS21]. Dugesia [CLTC15]. During [CDK23]. Dyadic [BM21]. Dynamic [GG20, ZRW20]. Dynamical [ABvSY21, ACM22, CDS21, CLM20, DS21, FB21, FB22, FFMV20, GR20, GL23a, GPS22, GHO21, Haf20, Hiu22, KK20b, KZ20, KM21, LS20a, LD22, PKG20]. Dynamically [ABB20]. Dynamics [ADS^{+}22, ABT23, BN22, BJ23, BOP23, Bla21, BPPS20, BR22c, BM20, CG23, CCR21, CLLL23, CB20, DS20, DG23, DGZ20, EGvdHN20, EGN20, FGM23, FvMST23, Fur23, GL22a, GZ20, HKR17, HKR18, HP20b, HKR23b, HKR23c, JVG21, JKG21, Jun20a, KS21, KLvR20, KP20, MMR20, NF20a, NIB21, RCF20, SKM20, SS23, VSG20, YS20, ZXD22, WPK21]. Dynein [NP21]. Dyson [LLX20].

Each [CLTC15, CLTC23]. Easy [ADS^{+}22]. Eden [KD20]. Edge [EFR23, FIS20, Seo20]. Edge-Transitive [FIS20]. Edges [LL23a]. Edgeworth [Haf20]. Effect [BJN21, BPR22, BO21, Ciu22, GH20a, HXX23, NPT20]. Effective [CLV22, Gol23, LS20c]. Effects [DS21, Phi20]. Efficient [Bód20, GT20, MB21, Nei21]. Ehrenfests [ABB20]. Eigenstate [SHRE23]. Eigenvalue [Due22, HK22, KK22, SM23a]. Eigenvalues [CCH20a]. Eight [FNS21]. Eight-Vertex [FNS21]. Einstein [CCS21, NPT20, Pec22, Tas20]. Elastic [BS20a]. Electric [HSS20, Li21]. Electrical [BR22a, BR22b]. Electrodynamics [Wre21]. Elephant [Ber22, Lau22, MT20]. Ellipsoidal [LS20b]. Elliptic [Due22]. Emergence [DZW^{+}20, DGZ20, HR16, HR23, SY20]. Emergent [HKR17, HP20a, HPRS20, HP20b, HKR23b]. Empirical [LLX20]. Emptying [CCS20]. End [FIS20]. Energy [Ban20, BJN21, Bob23b, FL20b, HL22, KY23, KO20, Lee22, LRRS21, Luc20b, MW20a, MM20b, jSY23, Sub22]. Engine [Xu21]. Enhanced [ZW22]. Ensemble [CW22, HPRS20, Hab23, SS22, SM23b, TCVB20]. Ensembles [FZ20, MPS21, NT20, Pie20, Seo20, Spo20, TT21]. Enskog [DT20, GPRR20]. Enskog-Type [DT20]. Entanglement [CG20, GOS20, Jon22]. Entropic [BB20]. Entropies [JK23, SN22]. Entropy [ABL21, BHL21, BCJP21, CGT21, CDK23, CLW23, DP21b, FGM23, GPS22, GOS20, JL22, Jon22, Mor20b, RJS22, TYM23, Tan22, ZM21]. Environment [FPT21, Haf20, Unt22, Jun20b]. Environments [WY22]. Epidemics [FIS20]. Epitaxial [Gat23]. Equality [SD21]. Equation [AOdLHT20, ABL21, ALT22, AMSY23, ABV22, AESW21, BKL20, Ben23, Bob23b, Bob23a, Bou20, CKL20, CCHS22, CLLL23, CS20, CCM20, CSZ21, DAN20, Dem23, Den22, DRB20, DSS20, FB21, FB22, FLM22, HJZ20, HJ23, JKNV23, Kep23, LL23b, MM20b, MW20b, NS22, Oto23, PR20, Qi21, NS21]. Equations [ABvSY21, AC21, BW20, Ben23, Bob22, BR22c, BDD23, CT22, Cop22, Cri23, EK20a, FPV22, Gao22, KR22c, LJ20, LYZ23, LY21, LSW22, Son20, TTZ22, XX20, YS21, ZD21, ZXJL23, ZV20]. Equi [CW22]. Equi-spaced [CW22]. Equidistribution [Tak20]. Equilibrium [AJR21, Cao21, CMRU20, DHP20, DMO20, GMV20, LL23b, Mor20b, NF20a, Oto23, BGJOS22, BR22c, EM20, FLNV23, GL22a, Men20, dZGFBC20]. Equipartition [Luc20b]. Erased [HS20b]. Erasing [WBC21]. Erdös [GGM21, AK21, CCH20a]. Ergodic [BHL23, LW20, Mor20a]. Ergodicity [JRA20, ZW23]. Error [WBC21]. Escape [Bód20]. Esseen [TD20]. Estimate [BS13, BS20b, Bód20, CLLL23]. Estimates [Bha20, EKN20, GM21b, RS22]. Estimation [CT22, GT20, KL20a, ZW22]. Euclidean [BC20, CES21, FK21]. Euler [AOdLHT20, Ben23, CB20, FPV22, KR22c]. Evaluation [FK21, MB21]. Even [Urr22]. Event [MGM23]. Event-Chain [MGM23]. Events [MQ20]. Evolution [BDR23, BB22, CCR21, FW21, TTV23, XX20]. Exact [GL23a, KTT23, Li21, Sam20a, Unt22]. Examples [BCJP21, DSS20]. Excess [NB22]. Excitable [Pie20]. Excitations [DS21]. Excited [Ngu22]. Exclusion [BBK^{+}23, DHS21, JKG21, MJB23, Mou23, NP21, vGR20]. Excursion [GK22a]. Existence [Den22, FP21, JLY20, RZ20, RZ22]. Existing [YK20]. Exit [Bha20]. Expanded [BTV21]. Expanding [DLDG21, DHS23, KS21]. Expansion [CDK23, NB22]. Expansions [DK23b, Gre22, Haf20, JK22, Sco21, NF20b]. Explicit [DRB20]. Explosion [JPR21]. Exponent [Mol20, Tex20]. Exponential [AHP20, GMV20, JT20, LS20b, JT21]. Exponents [FZ20, GHL20, Hut22, Lin21, Mol22]. Extension [Tan20]. Extensions [GL23a, LL20]. External [BCS22, BGN21, CW22, Kim21, RRA23, ZXD22]. Extinction [RNA23]. Extracting [LD22]. Extrapolation [Abh23]. Extremal [CFVY20, CKL23]. Extreme [AvdH20, MQ20, PMACF20, RB20]. Extremes [God21].

Facilitated [Sha20]. Facilitation [GLPP20]. Factorial [iSHNY22]. Faithful [BPC22]. Families [Ara21]. Far [LS21]. Fast [BPC22, EKN20, EGK21, FS20a, LL23a]. Fastest [DE22]. FBM [Son20]. Feature [Fur23]. Features [GH20b]. Fechner [Mae21]. Feller [GN21]. Feller-type [GN21]. Fermi [ABL21, Gol23, LL23b, MGPCA20]. Fermion [Jon22]. Fermionic [AJR21, CLL21]. Fermions [BY20, Mas22]. Fernández [Fia20]. Ferromagnet [AAVF23]. Ferromagnetic [BCM20, BdBBR23]. Ferromagnetism [TK21, Tan20]. Ferromagnets [Fac21a, Fac21b, SGNS20].

Feynman [KG23]. Fibrous [HXX23]. Field [AS21a, AHP20, ACCM21, Ass22, AC21, BW20, BCS22, Ben23, BGN21, BCD22, BPPS20, CK22, CG20, CDS21, CWZ20, CLL21, DFP20, FL23, FK21, GMV20, GL22c, GM21b, HLP20a, Hut22, IISS23, JP22, KK20a, Kim21, KL20a, Kos23, KO20, McD20, NP21, NP20, NS20, RRA23, Sak21, SGNS20, ZXD22, ZXJL23, Abh23]. Field-Theoretic [NS20]. Fields [ACM22, KM20, LT20, Mol22, MS22a]. Films [Gat23]. Filter [TCVB20]. Financial [KTT23]. Find [Xu21]. Finding [DE22]. Finger [McD20]. Finite [BDNS20, Các20b, CKL23, DIM23, GR20, LL20, LSW22, NRS20, Các20a]. Finite-Spin [CKL23]. Finite-Velocity [Các20b, DIM23, Các20a]. Finiteness [HKR23a]. First [DM22, EGvdHN20, FW21, Fre22, TR22]. First-Order [TR22]. Fit [PKG20]. Fitness [PKTM23]. Fixation [Bla21, PR21]. Fixed [CGT21, KR22a]. Flat [Bru21, FM21, GMS21, GMS23, MGPCA22, TK21, Tan20]. Flat-Band [FM21, Tan20]. Flat-to-Flat [GMS21, GMS23]. Flexible [JVG21]. Flickering [AFKH^{+}20]. Flights [DMS23]. Flip [CCR21]. Flocks [HP20a]. Flooding [WPK21]. Flow [GV20, HP21, SS23]. Flows [DG23, FH22, GR20]. Fluctuating [BS20a, BBC^{+}20, DDD^{+}20, God21, TZQY20, Unt22]. Fluctuation [Ban20, BGSRS20, BC22, DHP20, DW21, FGM23, Gal20, Gal21, KLSS20, Oto23, Sos23, THFH20, Tan22]. Fluctuation-dissipation [Tan22]. Fluctuation-Symmetry [BC22]. Fluctuations [BPP20, BB20, DMO20, Due22, Hiu22, JLS22, Liu21, Tex20, Yak21, dZGFBC20]. Fluid [FH22, TN18, TN21]. Fluids [Ben23, DHL20]. Fluxes [DHP20]. Fokker [ADT21, BW20, HLP20a, LLY22, Cop22, FPT21, LY21, MM20b]. Following [JPR21]. Football [Wan22]. Foraging [Wie20]. Force [ZXD22]. Forces [BM20, NPT20]. Forcing [GV20]. Forests [GS20, KMN22]. Form [Hiu21]. Formal [DK23b]. Formalism [BH22, LLV22]. Formation [CB20, FM21]. Formed [SPL20]. Formula [Mor20b, Wir22]. Formulas [NT23]. Formulation [CB20, HP21, JP22]. Foundation [Mae21]. Foundations [EK20a]. Four [FL18, FL20a]. Fourier [ALT22]. Fourth [Urr22]. FPUT [GAD20]. Fractal [CG22, Cha23, EN21, SPL20]. Fraction [GR22]. Fractional [ACM22, BDS20, DAN20, EKN20, Gao20, GRM20, HLW20, LJ20, MBL20, ORD22]. Framework [ZCD22]. Fréchet [PKTM23]. Fredrickson [Sha20]. Free [Ban20, CDK23, FK21, Fur23, GMRS20, GL22c, HDS22, Jon22, LRRS21, MSU20, Pig23, RJS22, Sak21, Sub22, TYM23, ZM21]. Free-Fermion [Jon22]. Freely [BS20a]. Freeness [SS21]. Frequency [HHW21, OMM22, TZQY20]. Frog [GP23]. Fronts [CB20, Der23]. Full [THFH20]. Fully [LY21]. Function [Ass22, CLY21, FK21, GGM21, LP22, SM23a, YS21, ZCD22, vdHvdHM23]. Functional [CM20a, CM20b, GHO21, JLM23, NPV23]. Functionals [BBC^{+}20, GJJ21]. Functions [AS21a, CH20, GKM21, HKR22, LLV22, Rue20, Wre22, Abh23, Jun20b]. Fuss [MSU20].

Galton [DM22, MPR20, dHW22]. Gap [BS13, BS20b, Ciu22, HL22, Sha20]. Gaps [Ciu22]. Gas [ABB20, CDK23, CHR22, Dem23, Fia20, HHW21, JR21, KLY21, KS22, NPT20, SKM20, SSvB21, SCDK23, TH22, Wen23, vEKM20]. Gases [ABE23, ALT22, BPR22, FKSS20, JV22, MM20b, MGPCA20, MGPCA22, NF20b, Pec22]. Gauge [DGZ20]. Gaussian [CJN20, CMRU20, GJ23, GL22c, HKR22, HS22, LD22, MPS21, Mol22, MS22a, PV20, Sak21, Sos23, Yak21, ZL21]. General [GJJ21, KLSS20, LW20, SS22, WLO20]. Generalisations [DT20]. Generalised [DRB20]. Generalization [EK20a, PKG20]. Generalized [Bob22, CGH22, CDLP23, CKL20, CB20, DP21a, GH20a, HKR17, HKR23b, HLW20, KG23, LLX20, Spo20, Tex20, TR22, ZCD22, ZV20]. Generated [FFMV20, PV20]. Generating [GGM21, YK20]. Generative [BKN22]. Generic [GPRR20, SSvB21, PKG20]. Gennes [SS23]. Geometric [ABV22, BS20c, CG22, DIM23, GP23, Meo23]. Geometry [BCS22, GMS21, GMS23, Sam20a]. Geostrophic [CCH^{+}20b]. Ghatak [AC21]. Gibbs [BdBBR23, CKL23, FX21, Hiu21, KK20b, Kos23, KM21, LS22a, Mor20b, RRA23, RZ20, RZ22, Roz22, iSHNY22, Spo20, SM23b, Tak23, WOK20]. Gibbsian [CDHM23]. Ginibre [SS22]. Ginzburg [AT21]. Giorgi [AMSY23]. Glass [ACCM21, AC21, BS23, BKN22, LRRS21]. Glasses [DS20, MW22, Sub22]. Glassy [GMRS20]. Glauber [FvMST23]. Global [CD20b, CD20a, Den22, FL20b, LY21, EGvdHN20]. Globally [ST22]. Goldstein [BS20b, AESW21, BS13, DE22]. Governing [LD22]. Grad [BGSRS20, GM21a]. Gradient [BL21, HP21, MM20a, Roz22, SS23, dZGFBC20]. Grand [BB22]. Granular [ALT22, GHO21, KS22]. Graph [AK21, EGvdHN20, Fre22, GGM21, Li21, SPL20]. Graphene [CN21]. Graphical [KMN22]. Graphons [Cop22]. Graphs [ABC23, CCH20a, CMG20, FIS20, GR20, HDS22, Har23, MM21, Ste22, YK20, vdHvdHM23]. Gravity [KTT22]. GREM [FP21, NF20a]. GREM-Like [FP21]. Gross [Oto23, CCS21]. Ground [CG20, RRA23, vEKM20, vEM20]. Group [AKL20, CDS21, KK20a]. Groups [HKR17, HKR23b]. Growth [DZW^{+}20, FGR20, Gat23, HOiS22, KD20, McD20, RF21, Unt22, ZL22]. Growth-Rate [Unt22]. GUE [PT20]. Guernsey [FB22]. Guerra [Fac21a, Fac21b].

Hadamard [KK22]. Haldane [SY20]. Half [BKL20, Lin21, Pig23, Wu20]. Half-Line [Lin21]. Half-Space [BKL20, Wu20]. Hall [GMP20]. Hamiltonian [CG23, DHO20, GL22a, NP20, SSvB21]. Hamiltonians [BCS22, BTV21, CLV22, DGT^{+}20]. Hanoio [SPL20]. Hard [GG22, HNK22, HG20, JKNV23, JR21, KK20b, Seo20, SCDK23, Urr22]. Hard-Core [KK20b]. Hard-Disk [HNK22]. Hard-Sphere [GG22, JKNV23]. Harmonic [BCD22, DHP20, Lep23, TZQY20]. Heat [BPP20, BCD22, DHP20, HOiS22, Xu21, GPRR20]. Heavy [HLN22].

Hebbian [AAAB22]. Hegselmann [LL22]. Height [Bód20, Roz22]. Height-Periodic [Roz22]. Heisenberg [HKR22]. Hermitian [Jan22]. Hessian [FL20b]. Heteroclinic [Rod21]. Hexagonal [BJ23]. Hide [Ber22]. Hierarchical [Jan20, Jan21, MW22]. Hierarchy [Fac21a, Fac21b, HP20b]. High [Bob23b, DK23a, EM22, FL20b, GK22b, HP21, HL22, HM20, KR22a, NT20, SLZ23]. High-Dimensional [FL20b, SLZ23]. High-Order [HP21]. High-Temperature [KR22a]. Higher [BH22, BPPS20, Bro22b]. Higher-Dimensional [BH22, Bro22b]. Hilbert [TTV23]. Hilliard [Mor20a]. HMF [LLM20]. Holes [Pel21]. Holevo [LMAC20]. Holocene [AFKH^{+}20]. Holonomic [KY23]. Homeo [LS22b]. Homoenergetic [Kep23]. Homogeneity [Wre22]. Homogeneous [ABT23, CGH22, FB22, GP23, LL23b, AESW21]. Homogenization [Due22, EGK21]. Honor [ACF^{+}20]. Hopf [TCDN20]. Hops [NP21]. Horns [Bru21]. Hubbard [FM21, TK21]. Hungry [Wie20]. Hydrodynamic [Doy22, GS22, vGR20]. Hydrodynamical [CN21]. Hydrodynamics [BBC^{+}20, Bob22, BdPGN20, DDD^{+}20, EGN20, JPR21, JR21, XU22, Oto23]. Hyperbola [PRD23]. Hyperbolic [CLY21, Cra21, DLDG21, LJ20]. Hypercube [ZW23]. hypergeometric [Abh23]. Hyperspheres [PMACF20]. Hyperuniformity [HKR22]. Hypocoercivity [ADT21, FS20a]. Hypothesis [Gal21, SHRE23]. Hysteresis [Phi20].

Ideal [NPT20, MC23]. Identities [Rue20]. II [BCJP21, BTV21, EGvdHN20, HB21, LS22a, TYM23, TCDN20, Yua20]. III [Hik22, TCND20]. Immigration [MBL20]. Implications [PMACF20]. Implied [SM23b]. Importance [GT20]. Improved [Har23, KD20]. Impulsive [GV20]. Inclusion [JCG20]. Indefinite [LL20]. Independence [Har23, Lyc20]. Independent [God21]. Index [CFVY20, SN22]. Index-Based [SN22]. Individual [ZL22]. Individuals [BM20]. Induce [Nis23]. Induced [AM21, Bód20, FW21, HSS20, KK22, SND^{+}20]. Inductive [Fia20]. Inelastic [BPR22, KS22, Qi21, GHO21]. Inequalities [CM20a, CM20b, JLR23]. Inequality [Pap20]. Infection [Phi20]. Infinite [BdBBR23, CGP20b, CLY21, FPV22, FP21, JRS20, Kos23, MC16, MC21, MB21, Nas22, jSY23, TH22, XX20, ZW23]. Infinite-Dimensional [ZW23]. Infinite-Order [MC16, MC21]. Infinite-Range [TH22]. Infinitely [BM20, CH22, HM23]. Information [BPC22, WBC21]. Inhomogeneity [BO21]. Inhomogeneous [CCH20a, CKL23, GN21, LLM20, LXZ23a, LXZ23b, Ste22, vdHvdHM23]. Initial [CD20b, DS20, DSS20, Gat23, HJ23, LT23]. Injection [FLNV23]. Insights [Lau22]. Inspired [NP21]. instabilities [GPRR20]. Instability [LLM20]. Integer [BHL21]. Integrability [Fen21]. Integrable [BVG21, BOP23, CK22, FGK20, GGGM23, KP20]. Integral [BR22c, Fre22, FKSS20, Gre22]. Integrals [FW21, Lyc20]. Integrated [DK23a, GL22b]. Integration [BDS20]. Intense [JPR21]. Interacting [BPPS20, CGV23, CDK23, Cri23, DFP20, Dem23, FGR22, FKSS20, GLPP20, GMP20, Gou23, JK23, JV22, LP22, Mas22, RS22, VSG20]. Interaction [Ban20, BB20, CDHM23, DM22, HJZ20, RZ20, RZ22]. Interactions [BN22, CG23, DHO20, FB22, GM21a, KK20b, MS22b, RRA23, SY20, ZV20, vEM20]. Interface [AT21, CK20, DMO20, IOVW20]. Interfaces [GL23b]. Interior [DGT^{+}20]. Interior-Boundary [DGT^{+}20]. Intermediate [EGK21, Wu20]. Internal [Pie20]. Interpolated [Fac21a, Fac21b]. Interpolation [BG20]. intersecting [GMS21, GMS23, JO20]. Interval [BY20]. Introducing [YS21]. Introduction [ACF^{+}20, KMN22, Luc20a]. Invariance [BRO22a, DHS23, IOVW20]. Invariant [ABE23, CLY21, FPV22, FNS21, LS22a, SHRE23]. Involving [DAN20, GH20b]. Ionic [ZRW20]. Ionization [HS21]. Irregular [GPS22]. Irreversibility [Bou20, FGM23, Pen20, WQ20]. ISAW [BO21]. Ising [AHP20, BEL22, BGN21, Bla21, CJN20, CJN22, CG20, DAT21, GMRS20, GMV20, GL23b, GOS20, KR22b, KLSS20, RRA23]. Isolated [Lep23]. Isometric [VPG20]. Isotropic [FZ20, SND^{+}20]. Issue [ACF^{+}20, Luc20a]. Iterated [CLY21]. Iteration [YS21].

Jacobi [TT21]. Jamming [DP21a, NIB21]. japonica [CLTC15]. Jarzynski [SD21]. Jellium [Sam20b]. Joel [ACF^{+}20]. Joint [BRO22a, Moh22]. Jump [DE22, GH20a, Pap20, Shi21]. Jumps [BGJOS22, CGJ21, Pap20]. Junctions [SN22].

Kac [BBO21, BB22, CT21, Hiu22, KG23]. Kac-Like [BBO21]. Kalman [TCVB20]. Kant [Hep20]. Kardar [BKL20, CCM20, KP20]. Kawasaki [BN22, BJ23, FvMST23, KLvR20]. KdV [SM23b]. Kelvin [WQ20]. Kernel [MGPCA22]. Kinetic [ABT23, Bob23b, Dem23, DT20, DTZ22, EK20a, EM20, FS20a, GMV20, GM21b, GBE22, KTT23, TN18, TN21, XTL20, ZL22]. Kinetics [BBC^{+}20]. Kingman [Yua20]. Kirkpatrick [ABvSY21, Ban20, DW21, IISS23, LL23a, LRRS21, Liu21]. Knudsen [KBK23]. Korteweg [Gao20]. KPZ [DP21a, FNS21, NS21, NS22]. KPZ-Equation [NS22]. Kramers [CWZ20, Son20, TD20]. Krause [LL22]. Krook [Oto23]. Kuramoto [BCS20, BW21, HPRS20, HKR23a, MM21, SLZ23].

Lack [PKG20]. Lack-of-Fit [PKG20]. Lagrangian [AOdLHT20, DHL20]. Lagrangians [CLV22]. Lamperti [Mol22]. Landau [ABL21, AT21, Bob23b, CDW21, DY21, FL23, FB21, LSW22, SS23]. Landauer [BC22, WBC21]. Landscape [FS20b, FL20b, Lee22, SM23a]. Landscapes [GM21b]. Langevin [DAN20, DRB20, DG23, GJ23, ZV20]. Lapses [GNH21]. Large [ADT21, AK21, AESW21, BHL23, BBC^{+}20, BBO21, BdBBR23, Bou20, BR22c, CG20, CGH22, CGJ21, CLV22, CLL21, CH20, DHS21, FB21, FB22, GBE22, GT20, HLN22, HM23, KL20b, LLX20, LS20d, RS22, RB20, SM23a, SGG23, Sco21, Tak23, vEM20]. Last [Bha20, CCS20]. Last-Passage [CCS20]. Lattice [BJ23, BO21, CMRU20, CCR21, CHR22, DK23a, FvdH21, FvdH22, GS20, HPRS20, HOiS22, IT20, JVG21, KK20b, KL20a, KP20, Mad23, Sak21, SKM20, SSS21, VSG20, Wat19, Wat20, Yak21, YS21, vEKM20, SSvB21]. Lattice-Gas [vEKM20]. Lattices [KTT22, LL20, SH23]. Law [CH22, GMV20, LLX20, LS20d, MBL20, Ste22, GPRR20, Hiu21, WQ20]. Laws [LD22, Mae21, Oto23]. Lax [GGGM23]. Layer [AS21b, FGR22, KBK23]. Layers [Gat23]. Laziest [MT20]. Leadership [Mol20]. Leakage [dSL23]. Learning [AS21b, BL21, Fur23, Li21]. Least [SS22]. Lebowitz [ACF^{+}20]. Lee [HJN23]. Leffler [Ber22]. Lemma [CCHS22]. Lenard [FB22]. Length [FB21, GK22a, VSG20]. Lengths [NPV23]. Leroy [Abh23]. Level [CGJ21, CN23, FP21, Tak23]. Level-2 [Tak23]. Level-Set [CN23]. Levels [MM20b]. Lévy [BHP21, BDS20, DMS23, ZD21, ZXD22]. Lie [DHL20, HKR17, HKR23b, KK20a]. Lifetime [GP23]. Lifshitz [GRM20]. Light [Ser20]. Light-Speed [Ser20]. Like [BBO21, Các21, FP21, Sam20a]. Like-Charged [Sam20a]. Likelihood [YS20]. Limit [AS21a, ABC23, AT21, BGSRS20, CDN21, CGH22, CDLP23, CQ21, CDH21, CP20, Czu22, EM22, FK21, FX21, FS20a, FPV22, FNS21, GM21a, GS22, HM23, HLP20a, HLP^{+}20b, JKNV23, KK20a, KL20b, LS20c, Liu20, Lyc20, MT20, MW20b, NP20, PR20, ST22, Seo20, TTZ22, vGR20]. Limit-Cycle [CQ21]. Limited [VSG20]. Limits [CLL21, CS22, FR20, FGR22, JO20, LS22b, Pin23, WY22]. Line [ACCM21, CES21, CDD20, Lim20, Lin21, MW22, NT20, PR21]. Linear [BM21, BB20, BDD23, CCS20, FS20a, HM23, Yak21, ZD21, ABV22, WLO20]. Linearized [ADT21, BW20, LS20b]. Links [YK20]. Liouville [KO20]. Liquid [SND^{+}20]. Local [AT21, ALP21, CGL21, CKL20, DY21, EN21, EM22, GAD20, HK22, HNK22, JKG21, KV21, KL20a, Sco21, SGNS20, Sos23, Wu21, ZV20, CD20a, Oto23]. Localised [FM21]. Localization [ABV22, Chu23, CSZ21, GH20a, HK22, KTT22, BTV21]. Location [GLY21]. Location-Based [GLY21]. Locked [BW21, HR16, HR23, HKR23a]. Loewner [FLM22]. Log [Pap20]. Log-Sobolev [Pap20]. Logarithmic [Hik22, JLR23]. Lohe [HR16, HKR17, HKR18, HP20a, HP20b, HKP21, HP21, HKR23b, HR23, HKR23c, KK20a]. Long [BGJOS22, CG23, CSZ21, EGvdHN20, EM22, FB22, GM21b, HJZ20, Ngu22, SS23, WOK20, XTL20]. Long-Range [CG23, EM22, HJZ20, WOK20, XTL20]. Long-Time [CSZ21, GM21b]. Longer [Wie20]. Looking [Pen20]. Loop [FKSS20, HS20b]. Loop-Erased [HS20b]. Lorentz [ABB20, Rya21, SKM20, TH22, Wen23]. Lorentz-Gas [TH22]. Lorentzian [Wre21]. Lorenz [Ara21, GHL20, GV20]. Loss [BKN22, LLV22]. Low [KC21, Rya21]. Lower [Moh22]. LSW [CD20b]. LTE [JV22]. Lucky [Wil23]. Lyapunov [FZ20, Gal21, GHL20, Lin21, Moh22, Tex20].

Machines [AS21b, ABCM20, Zha20]. Macroscopic [GL22a, Hiu22, MJB23, TH22, TTV23]. Magnetic [BW20, Ben23, BCD22, CG20, CWZ20, FL23, GMV20]. Magnetization [CGT21, JLS22]. Magnetized [CDW21]. Mainly [Dem23]. Majority [MMR20]. Maki [JRS20]. Manhattan [CDD20, Rya21]. Manifold [vGR20]. Manifolds [FL20b, HKR18, HKR23c]. Many [BM20, CLLL23, DP21b, SEG22]. Many-Body [CLLL23, DP21b, SEG22]. Maps [DLDG21, KC21, MSG^{+}20, ST22]. Marcenko [EK20b]. Margin [AFKH^{+}20]. Marked [RZ20, RZ22]. Market [Wan22]. Markov [MC21, ADS^{+}22, CLV22, GL23b, GH20b, GL20, HNK22, JL22, LS22a, MC16, MS21, Pap20, Pin23, Shi21, Tak20, Tak23]. Markovian [CGJ21, DFP20, RV21]. Martingale [GJJ21, Lau22]. Mass [CJN20, HLW20, KR22b, LS20c]. Massimo [Pel21]. Matched [MJB23]. Matching [BC20]. Matchings [CES21]. Materials [HXX23]. Mathematical [WQ20, You20]. Mathematics [DT21]. Matrices [CW22, GGGM23, HK22, Jan22, Shc20, Tex20]. Matrix [BR22c, BH20, BCS20, CLLL23, FZ20, HP21, HB21, Hik22, KK20a, LL23a, Seo20]. Matrix-Valued [BCS20]. Maximizing [CP20]. Maximum [CDN21, CDHR23, YS20, ZCD22]. Maxwell [FL23, BPR22, EK20a, Kep23, KS22, YS21]. Maxwellian [LY21]. Maxwellians [DY21]. Mean [ACCM21, AC21, BPPS20, CLL21, DFP20, FvMST23, GM21b, HLP20a, Hut22, JP22, KK20a, KL20a, Kos23, NP21, NP20, SGNS20, ZXJL23]. Mean-Field [ACCM21, BPPS20, CLL21, HLP20a, Hut22, KK20a, KL20a, Kos23, SGNS20, ZXJL23]. Meander [Kar20]. Measure [BdBBR23, CLY21, HS22, JLR23, KK22, LS22a, LLX20, Qi21]. Measured [VLC23]. Measurement [BPP20, Wre23, ZW22]. Measurements [BCJP21]. Measures [CB20, FPV22, FNS21, JO20, LT21, NT23, PV20, RRA23, Roz22, Wu21]. Measuring [PMACF20]. Mechanical [Mae21, IOS21]. Mechanics [BS20a, BVG21, DG20, Fac21a, Fac21b, Hep20, Luc20a, Rue20]. Media [Các20b, GHO21, Mor20a, ZXD22, Các20a]. mediterranea [CLTC15]. Medium [BdPGN20, Tan22]. Melting [Maj23]. Membrane [CN23, CDHR23]. Membranes [FK20]. Memory [Bau20, CGL21, GNH21]. Merge [IT20]. Merging [FN20]. Metals [BR22b, BR22a]. Metastability [BJ23, BGM20, Kim21, dSL23, Lee22]. Metastable [BJN21, GMV20, KM20, LS22a]. Metastates [Kos23]. Method [BG20, DK23b, MJB23, Tan20, Xu21]. Metric [Hab23]. Metrics [LL20]. Metropolis [NF20a]. Micro [FPT21]. Micro-organism [FPT21]. Microcanonical [ABV22, CGT21, KL20a, SM23b, TR22]. Microscopic [Ciu22, EK20a, Pen20, Xu21]. Mild [BL22, Cri23]. Mill [EFR23]. Minima [SGNS20]. Minimal [MT20]. Minimization [DP21b]. Minimizer [Tak20]. Minimum [FL20b]. Mirror [Roz22, Rya21]. Mirrors [Rya21]. Mitigation [FGM23]. Mittag [Ber22]. Mittag-Leffler [Ber22]. Mixed [BCS22, CG23, CLLL23, WM22]. Mixed-Norm [CLLL23]. Mixing [GL23a, MP22]. Mixture [Jan20, Jan21, WLB^{+}20, dZGFBC20]. Mixtures [HHW21]. Mode [CDS21, GG20]. Mode-Coupling [CDS21]. Model [ABvSY21, AHP20, AT21, ABB20, AC21, ABT23, BY20, BDR23, Ban20, BKN22, BB22, BdBBR23, BOP23, BS13, BM21, BdPGN20, BCS20, BW21, CJN20, CJN22, CN21, CG20, Cao21, CN23, CDH21, CDHR23, CD20b, CT21, CCH^{+}20b, DFP20, DDG20, DW21, DE22, DT20, DK23a, DZW^{+}20, DK23b, EN21, FR20, FNS21, GP23, GMRS20, GMV20, GRM20, Gol23, Gou23, GOS20, HR16, HKR17, HKP21, HP21, HKR23b, HR23, HKR23a, HHW21, HOiS22, Hiu22, IISS23, JLS22, JM23, JMPR23, JRS20, KTT23, KV21, KK20a, KW20, Kim21, KR22b, KTT22, Kos23, KM20, KM21, LL22, Lee22, LL23a, LLM20, LRRS21, LS20b, Lim20, Liu20, Liu21, MW20a, MM21, MGPCA20, MT20, MS22a, PR21, PN23, RRA23, RF21, Roz22, SM23a, SSvB21, Sha20, SLZ23, TN18, TN21, TH22, TCND20, TR22]. Model [Vol22, Wil23, WM22, Wre21, You20, Yua20, Zam22, dHW22, Iin21, BS20b]. Modeling [AFKH^{+}20, MM20a]. Modelling [AOdLHT20, GHO21, KBK23]. Models [AKL20, BDR23, BCM20, BGN21, Bob23b, BH20, Chu23, CH22, CKL23, Cra21, CS22, DAT21, DTZ22, Eld20, FM21, Gat23, GH20a, HJ21, HB21, Hik22, KLY21, KK20b, KLSS20, KL20a, LW20, Li21, LMAC20, MM20b, MS22b, RB20, RV21, Rya21, TK21, Wat19, vEKM20, Wat20]. Moderate [Bha20, Sco21, XZ21]. Modes [AS21b, SGG23]. Modified [ALT22, Bob23a, LS23, Pap20]. Molecular [BOP23]. Molecules [Kep23, jSY23]. Moment [LYZ23]. Moments [Ass22, BK21, Haf20, JT20, JT21, PR20]. Momentum [KBK23, WFKM^{+}20]. Monolayer [KS21]. Monotonicity [LT21, MSG^{+}20]. Monte [ADS^{+}22, FK21, MGM23, WM22]. Motion [ACR21, DGZ20, EKN20, Fre22, FvMST23, HJN23, LLX20, LS20d, Lyc20, Ser20, vGvGR21]. Motions [DIM23, PT20]. Movement [FPT21]. Moving [SKM20]. Mpemba [BPR22]. Multi [AFKH^{+}20, ACCM21, BCM20, CGV23, DW21]. Multi-overlaps [BCM20]. Multi-proxy [AFKH^{+}20]. Multi-species [ACCM21, CGV23, DW21]. Multicolor [KL20b]. Multicomponent [FLNV23, YS21]. Multidimensional [BHL21, GN20]. Multifractal [ACM22]. Multifrequency [FW21]. Multilayer [PD22]. Multiple [Nas22]. Multiplicative [BHL21]. Multiscale [Gao20]. Multitype [FGR20]. Mutant [PKTM23]. Mutation [PKTM23, Yua20].

Nanoribbons [CN21]. Nanotube [SN22]. Natural [BDR23, BL21]. Navier [ALT22, FF20, Gao22, KR22c, LLY22, LYZ23]. Navigation [GS20]. Near [CJN20, DY21, KR22b, Shc20, TH22]. Near-Critical [CJN20, KR22b]. Nearest [CDHM23]. Nearly [TK21, Wen23]. Necessary [JK22]. Negative [ES22]. Negatively [BRO22a]. Neighbour [CDHM23]. Nematic [SND^{+}20]. Network [Iin21, PPR22, Wan22]. Networks [AAAB22, BDR23, BKN22, DHP20, DAR22, DG20, FS20a, GS20, HLP^{+}20b, Lep23, Li21, MC23, NRS20, RCF20, ST22]. Neumann [Ben23]. Neural [AAAB22, HLP^{+}20b]. Neurons [GLPP20, dSL23, Nas22, RNA23].

Neuroscience [Hep20]. Newtonian [JLY20]. Next [BEL22]. Next-to-Diagonal [BEL22]. Niño [TCND20]. Nishimori [ACCM21]. No [Wie20]. NoBLE [FvdH21, FvdH22]. Nodal [NPV23, PV20]. Nodes [YK20]. Noise [AOdLHT20, AKL20, BHP21, Bód20, CCH^{+}20b, FW21, GJ23, Lep23, MSG^{+}20, Nis23, ORD22, TZQY20, ZL21]. Noise-Induced [Bód20, FW21]. Noises [ZD21]. Noisy [HLW20]. Non [AMSY23, ABV22, AESW21, ALP21, BC20, BGJOS22, BR22c, CG23, CGP20b, CM20b, CT22, CD20a, DFP20, DMO20, Den22, EN21, EGN20, EM20, FLNV23, FGM23, FGK20, FFMV20, GL22a, GMS21, GMS23, GJ23, GM21b, HXX23, Jan22, JV22, JO20, Kep23, KK20b, KY23, LS22a, LD22, MSG^{+}20, Men20, MC23, NPT20, NIB21, Oto23, SM23b, TYM23, WLO20, ZXD22, ZM21, dZGFBC20, KM21, CM20a]. Non-commutative [CM20b, CM20a]. Non-compact [FGK20, TYM23, ZM21]. Non-conservative [EN21]. Non-constant [BC20]. Non-convex [GM21b]. Non-Cutoff [Den22, AMSY23, Kep23]. Non-equilibrium [BGJOS22, BR22c, FLNV23, GL22a, Men20, dZGFBC20]. Non-Gaussian [GJ23, LD22]. Non-Gibbs [KK20b]. Non-Hamiltonian [CG23]. Non-Hermitian [Jan22]. Non-homogeneous [AESW21]. Non-ideal [MC23]. Non-intersecting [GMS21, GMS23, JO20]. Non-linear [ABV22, WLO20]. Non-local [CD20a, Oto23]. Non-LTE [JV22]. Non-monotonicity [MSG^{+}20]. Non-Normal [SM23b]. Non-parametric [CT22]. Non-reversible [EGN20, LS22a]. Non-Static [ZXD22]. Non-uniform [HXX23]. Non-uniformity [KY23]. Non-unital [FGM23]. Non-universal [NPT20]. Non-zero [CGP20b]. Noncommutative [JLR23, Wir22]. Noncondensation [GLY21]. Nonconventional [BHL23]. Nonequilibrium [BR22b, BL22, DG23, Gal20, KMN22, Tan22, WFKM^{+}20, YS21, YS20, YQ21, ZW22, BR22a]. Nonidentical [SLZ23]. Noninteracting [Pec22]. Nonlinear [BH22, CLM20, CQ21, CSZ21, CD20a, DDD^{+}20, DSS20, LLM20, LY21, NB22, ZD21]. Nonlocality [BGH20]. Nonperiodic [LL20]. Nonrelativistic [Wre21]. Nonstationary [DP21a]. Norm [CLLL23]. Normal [HS20a, LS20a, Seo20, SM23b]. Note [BKN20, Cop22, Sha20]. Notion [Lyc20]. Nucleus [MPDH20]. Number [ABE23, CES21, God21, HLP^{+}20b, Kut23, MSG^{+}20, ZCD22]. Numbers [BR21, LLX20, LS20d]. Numerical [BR22c, MB21, NS21, RNA23, Xu21].

Observable [HS20b]. Observables [CH20, GAD20, RB20, Wie20]. Obstacle [FH22]. Obstacles [Pec22]. Occupation [ZCD22]. Ocean [Pie20]. Official [Nei21]. One [BOP23, CDN21, DHO20, FL18, FL20a, FS20b, KK22, KL20b, NP21, Pec22, Pig23, SH23, SSvB21, Sha20, XU22]. One-Dimensional [BOP23, CDN21, DHO20, FS20b, KK22, Pec22, SSvB21, XU22]. One-Half [Pig23]. One-Sided [KL20b]. Only [BGH20, Sam20a]. Onsager [YQ21]. Open [BR22a, BR22b, CGH22, CGJ21]. Operational [AS21b]. Operator [ZW23]. Operators [Due22, NT23]. Opinion [GZ20]. Optical [WLO20]. Optimal [BS13, CGP20a, CES21, CM20a, CM20b, LS20b, PR20, Shi21, Xu21, BS20b]. Optimality [VLC23]. Optimization [Zha20]. Optimizing [LMAC20]. Orbital [HR16, HR23]. Order [BS23, BPPS20, DMS23, GJJ21, HP21, MC16, MC21, MM21, Nis23, SND^{+}20, TR22]. Ordered [BCD22]. Orders [DAN20, WOK20]. organism [FPT21]. Ornstein [BHP21, GMW20]. Orthogonal [BBK^{+}23, MPS21]. Oscillation [TCND20]. Oscillations [CQ21, WFKM^{+}20]. Oscillator [HLW20, ORD22]. Oscillators [GKM21, HKR18, HKR23c, Hab23, Men20, TZQY20]. Oscillatory [DFP20]. Other [BDR23]. overlaps [BCM20]. Overs [Der23]. Own [CLTC15, CLTC23].

Packing [Mou23]. Packings [HNK22]. Pair [FM21, Sam20b]. Pairing [Gal21]. Palmer [AC21]. Parabolic [KKMS20, dHW22]. Paradox [Bou20, CDW21, BGH20]. Parallel [ADS^{+}22, DAT21, TBD^{+}20]. Paramagnetic [Ban20]. Parameter [GP23, PT20]. Parameters [Lim20]. Parametric [EK20b, CT22]. Parasitic [Phi20]. Parisi [BKL20, CCM20, KP20]. Parking [DT21]. Part [CTDN20, EGN20, TCDN20, TCND20]. Partial [Cri23]. Partially [BW21, LP22]. Particle [BR21, BKN20, CLLL23, FGR22, FL18, FL20a, FGK20, HG20, JK23, LT23, MM23, MGM23, MW20b, SKM20]. Particles [ACR21, CGV23, Gou23, GM21b, JKG21, KK21, KS22, LL23b, Mol20, Pig23, Ser20, VSG20]. Partite [Li21]. Partition [Ass22, LP22, ZCD22, Jun20b]. Partitions [CP20, FX21]. Passage [Bha20, CCS20, EGvdHN20]. Past [FH22]. Pasta [Gol23]. Pastur [EK20b]. Path [BDS20, BR22c, CDD20, FKSS20, Gre22, GBE22, LD22, SGG23]. Path-Integral [FKSS20]. Paths [BDNS20, EGvdHN20, EKN20]. Pathwise [BGM20]. Patterns [BDR23, EN21]. PDE [CD20a, Fac21a, Fac21b]. PDE/Statistical [Fac21a, Fac21b]. PDMP [MGM23]. Pearson [ALP21]. Perceptron [ES22]. Percolation [AK21, Bha20, BDNS20, Bro22b, CG22, Cha23, CN23, CMG20, EGvdHN20, GR22, Har22, HM20, Hut22, Jun20a, MPR20]. Perfect [Ben23, BGH20]. Periodic [AM21, FS20b, JLY20, Mas22, Roz22, Tak20, Wen23, ZXJL23, PV23]. Periodically [THFH20]. permeable [FK20]. Permutation [ZCD22]. Persistence [AFKH^{+}20, Mol22]. Perspective [Bou20, CH20, SS21]. Perturbation [AS21a, BTV21, CJN22, Fre22, HJ21]. Perturbations [CQ21, Yak21]. perturbative [Abh23]. Perturbed [PT20]. Perturbing [NT23]. Peru [AFKH^{+}20]. Phase [BS20c, BW21, CN23, GR20, GLY21, HR16, HR23, HKR23a, HOiS22, JCG20, LXZ23b, MQ20, MW20a, MC23, Mor20a, MS22a, Nas22, NRS20, SLZ23, TN18, TN21, TR22, WM22, ZCD22, LXZ23a]. Phase-Locked [BW21, HR16, HR23, HKR23a]. Phases [BO21, JP22, KM20]. Phenomena [Iin21]. Phenomenon [EFR23]. Photon [OMM22]. Physical [NT23]. Physics [KP20, Pel21]. Picture [BPP20]. Piecewise [DLDG21, MM23].

Piecewise-Tunneled [MM23]. Pinned [FL20b]. Pinning [GH20a]. Pitaevskii [CCS21]. Planar [DG23, JT20, JT21]. Planarian [CLTC15]. Planck [ADT21, BW20, Cop22, HLP20a, LLY22, LY21, MM20b]. Plank [FPT21]. Plasma [EK20a]. Plasmas [FB21]. Plasticity [RV21]. Playing [Ber22]. Point [FK21, FFMV20, KR22a, KTT22, MPDH20, MP22, RZ20, RZ22]. Points [AvdH20, GPS22, RJS22, Tak20]. Poisson [ADT21, CKL20, DY21, HLP20a, TTZ22, CS20, CDD20, MPDH20, McD20, Meo23, NT20, Pec22]. Poisson-Based [Meo23]. Polarization [PN23]. Polaron [LS20c, MS22b]. Pollicott [CTDN20, TCDN20, TCND20]. Polyatomic [jSY23]. Polygonal [BDR23, DLDG21]. Polygons [Mou23]. Polymer [CS22, DG20, Fia20, GK22b]. Polymers [BO21, JRA20, JO20, Mad23, WY22, Wu20]. Polynomial [BBK^{+}23, HP21]. Polynomials [MPS21, Shc20, TT21]. Polynuclear [Gat23]. Poor [Maj23]. Population [CH22, FPT21]. Populations [AM21, SBB^{+}21]. Porous [BdPGN20]. Posedness [CKL20]. Position [LT23]. Positive [LT21]. Positively [BRO22a]. Positivity [FGR20]. Posterior [LLV22]. Potential [Bód20, FL23, Hik22, Kim21, KLvR20, TH22, YQ21, WPK21]. Potentials [EM22, HS21, Li21, Pin23, Qi21, Sak21, SY20]. Potts [BdBBR23, BGN21, Iin21, IOVW20, JLS22, KM20, KM21, Lee22, Liu20]. Power [MBL20, Ste22]. Power-Law [MBL20, Ste22]. Precise [Mou23, Sco21]. Predictability [SSS21]. Prediction [HM23]. Predictive [Hiu21, MC16, MC21]. Prefactors [BR22c]. Preferential [GLY21, RCF20]. Preimage [CLW23]. Presence [BCD22, CWZ20, FK20, KG23, Pec22]. Preserving [CCH^{+}20b]. Pressure [LW20, Wu21]. Principal [SM23a]. Principle [BHL23, BC22, CB20, DHS23, Gao20, IOVW20, LT20, Tak23, XX20, YQ21]. Principles [BRO22a, Gao22, RV21]. Probabilistic [DAT21, Har22]. Probabilities [Yua20]. Probability [CH22, DGZ20, WFKM^{+}20]. Problem [HJ23, KK22, LMAC20, Mol20, Rue20]. Problems [BS20d]. Procacci [Fia20]. Process [BGJOS22, BBK^{+}23, CJN20, CDLP23, CDD20, DIM23, DHS21, GMW20, GN21, GL22b, HXX23, JCG20, JKG21, LLX20, MJB23, MBL20, NP21, OMM22, Pap20, PT20, PRD23, XX20, vGR20]. Processes [BDS20, CLV22, CH20, FP21, FGK20, FFMV20, Fur23, GJJ21, Gre22, Haf20, HLN22, Kut23, LXZ23a, LXZ23b, MC16, MM23, Meo23, MP22, RZ20, RZ22, Shi21, VLC23, MC21]. Processing [BPC22]. Product [Tex20]. Production [ABL21, BCJP21, FGM23, Tan22]. Profiles [Jan22, Kep23]. Projections [Doy22]. Proof [Fia20, Wat19, Wat20, WOK20]. Proofs [BGH20]. Propagation [Các21, Cao21, CT21, Cri23, GG22, GM21b, JPR21, JR21, LSW22, NR21, WZN20]. Properties [ABV22, BGJOS22, Bob22, BDS20, HM23, LT21, MGPCA22, Moh22, WLO20, Wre22]. Property [GL23b]. proxy [AFKH^{+}20]. Pruning [KZ20]. Pseudo [SY20]. Pseudo-Potentials [SY20]. Publicity [Nei21]. Puglisi [Pel21]. Pulled [Der23]. Punctures [BH20, HB21, Hik22]. Pure [ES22, Pap20, TTV23]. Purposes [KMN22].

Pursuit [Mol20]. Pushed [Der23].

Quadratic [LW20]. Quantifying [Wil23, ZL21]. Quantitative [Men20]. Quantities [MB21]. Quantity [TH22]. Quantization [FK21, GMP20]. Quantum [BY20, BL21, BPP20, BCJP21, BR22a, BR22b, BTV21, BB20, CGP20a, CG20, CGH22, CM20a, CM20b, CGJ21, DP21b, DGZ20, FGM23, FGK20, GM21a, GOS20, HJ21, Hab23, HSS20, LT21, LRRS21, LS23, LMAC20, Luc20b, Lyc20, MW20a, MW22, Mor20b, MS21, NR21, SBB^{+}21, SEG22, VLC23, Wre21, Wre23]. Quantum-State [VLC23]. Quasi [BGM20, CCH^{+}20b, Mas22, Oto23, PV23, iSHNY22, SH23]. Quasi-Equilibrium [Oto23]. Quasi-Geostrophic [CCH^{+}20b]. Quasi-One-Dimensional [SH23]. Quasi-Periodic [Mas22, PV23]. Quasi-static [iSHNY22]. Quasi-stationary [BGM20]. Quasistationarity [FGR20]. Quenched [AKL20, HS20a, Mad23, MPR20]. Quotients [Pin23].

Rabbit [Wie20]. Radially [Bob23b]. Radiation [Dem23, EK20a, JV22]. Radius [FLM22, Moh22, Seo20]. Rainfall [Wil23]. Random [AHP20, ABC23, AT21, BR21, BK21, BCM20, Bau20, BC20, Ber22, BRO22a, BTV21, BS20d, Các20a, Các20b, CDN21, CCH20a, CGL21, CQ21, Chu23, CW22, CDHM23, DIM23, CS22, Czu22, DMS23, DS21, DHS23, Due22, EFR23, FIS20, FZ20, FFMV20, FL20b, GK22a, GGM21, God21, GL22b, GN20, GNH21, GGGM23, Hab23, Haf20, HDS22, Har23, HS20a, HS20b, HK22, HLW20, IT20, Jan22, Kar20, KR22b, Kos23, LT20, Lau22, LJ20, LS21, LS22b, MW20a, MPDH20, MM21, MM23, MT20, Mol20, Mol22, MS22a, NRS20, Ngu22, PD22, PV20, Pel21, SH23, Seo20, Ser20, Shc20, Ste22, Tex20, WLB^{+}20, YK20, Yua20, ZW23, vdHvdHM23, Jun20b, WPK21]. Randomized [KL20b]. Randomly [KK21, Unt22]. Randomness [FH22]. Range [BGJOS22, CG23, EM22, FB22, GKM21, GM21a, HS21, HJZ20, SY20, TH22, WOK20, XTL20]. Rascle [DTZ22, DT20]. Rate [CH20, DE22, DT21, MC16, MC21, Son20, Unt22]. Rate-Distortion [MC16, MC21]. Rates [Men20]. Ratio [CH20, Har23]. Reacting [KLY21]. Reaction [CGV23, EK20a, FS20a, MM20a, MC23]. Reaction-Corrected [EK20a]. Reactions [GL22a]. Reactive [KG23]. Real [FK21, NT20, Wre22]. Real-Valued [Wre22]. Realizations [TBD^{+}20]. Record [AFKH^{+}20, GL22b]. Recovery [Li21, PV23]. Recursive [NT23]. Reduced [CTDN20, CLLL23, TCDN20, TCND20]. Reduction [PKG20]. Reeb [Fre22]. Refining [WBC21]. Reflected [PT20]. Reflection [LT21]. Regime [Ban20, CCS21, DW21, Due22, EGN20, GH20a, GMP20, GG20, HK22, KM21, Wu20]. Regimes [AFKH^{+}20, DP21a, NS22]. Region [Lim20, WM22]. Regular [BdBBR23, CDHR23, HPRS20, Har23, KTT22, MS22b]. Regularity [CCHS22]. Regularization [EGK21]. Regulation [VSG20]. Reinforced [Bau20, BRO22a, EFR23, GNH21, Lau22]. Related [CJN20]. Relation [Fac21a, Fac21b, FGM23, Gal20, NS20, NS21, NS22, Shi21, THFH20, Tan22, Wre22]. Relations [VPG20]. Relative [Mor20b]. Relativistic [EK20a, LSW22, MGPCA20, MGPCA22, PR20, Pig23, GPRR20]. Relaxation [CG23, HKR18, HKR23c, MGPCA20, NIB21]. Relevance [Wre22]. Remark [Lyc20]. Remarks [BG20, GLV21]. Renewal [HLN22, Zam22]. Renormalization [AKL20, CDS21, KC21, Abh23]. Renormalizing [CCM20]. Rényi [AK21, CCH20a, GGM21]. Repeated [BCJP21, BB20]. Replica [AAAB22, ABCM20, BS23, DW21, Har23, IISS23]. Replica-Symmetry [ABCM20, IISS23]. Replicated [GLV21]. Representability [JLM23]. Representation [WQ20]. Representations [KMN22]. Reproductive [CLTC15]. Repulsive [BM20, MP22, NF20b, Sak21]. Research [WZN20]. Reservoirs [BPP20, BdPGN20, BL22, HM23, XU22]. Reset [CH20, HJ21]. Resetting [CDN21]. Resistances [SPL20]. Resolved [Jon22]. Resonance [HLW20, ORD22, Rod21, TZQY20]. Resonances [CTDN20, TCDN20, TCND20]. Resources [VSG20]. Response [BB20, DS21, GL20, Hab23, NB22, ZD21, ZRW20]. Restricted [AS21b]. Result [RZ20, RZ22, Rya21]. Results [DW21, KLSS20, Sam20a]. Resummation [Abh23]. Retrieval [AS21b]. Reverse [LMAC20]. Reversibility [vGvGR21]. Reversibillity [Gal21]. Reversible [Bou20, JKG21, EGN20, LS22a]. review [Pel21]. Revisit [CCHS22, GL22a]. Revisited [CSZ21, KTT23, MPS21, dHW22]. Revisiting [BS23]. RG [KR22a]. Riccati [BR22c]. Richer [ABB20]. Riemannian [HKR18, HKR23c, vGR20]. Rigid [GL23b]. Rigorous [Vol22]. Ring [HPRS20, Hiu22, KTT22]. Robustness [CH20]. Role [BOP23, VSG20, vGvGR21]. Rooted [KZ20]. Rotation [MSG^{+}20]. Rotational [ABE23]. Rotationally [KP20]. Rotations [Czu22]. Rotor [IOS21]. Rough [KS22]. Roughening [XTL20]. Roughness [KBK23]. Rowlinson [KK20b]. RSA [WLB^{+}20]. Ruelle [CTDN20, TCDN20, TCND20]. Rumor [JRS20, WZN20]. Run [vGvGR21]. Run-and-Tumble [vGvGR21].

Sample [BDS20, EKN20, LD22, SGG23]. Sampling [GT20, PMACF20]. Sandpile [EN21, KW20, PR21, SSS21]. Satisfaction [BS20d]. Saturation [FR20, Mou23]. Scalar [FK21]. Scale [HDS22, TCVB20]. Scale-Free [HDS22]. Scaling [CDLP23, CDH21, FR20, FGR22, GH20b, JO20, KR22b, KL20b, NS22, Seo20, SSS21, SCDK23, YS21, vdHvdHM23]. Scatterers [Wen23]. Schmidtea [CLTC15]. Schrödinger [ABV22, BGH20, CLLL23, CSZ21, DSS20, Gao20, HP20b, ZW23]. Science [WZN20]. Second [GJJ21, LT23, MM21, Tan22, Hiu21, WQ20]. Second-Class [LT23]. Second-Order [GJJ21, MM21]. Seek [Ber22]. Selection [McD20, PKTM23]. Selective [CB20]. Self [Kep23, LP22, MSU20, SSS21]. Self-avoiding [LP22]. Self-decomposability [MSU20]. Self-similar [Kep23, SSS21]. Semi [BO21, FK20, MB21]. Semi- [MB21]. Semi-permeable [FK20]. Semi-stiff [BO21]. Semiclassical [CLL21]. Semiconductors [BR22a, BR22b].

Semigroup [RJS22, TYM23, ZM21]. Sensitivity [AvdH20, GL20]. Separable [HKP21]. Separation [MC23, TCVB20]. Sequences [FFMV20, HS20a, Zam22]. Sequential [FFMV20, PD22, WLB^{+}20]. Sequentially [Hiu21]. Series [AS21a, BJN21, CFVY20, CT22, GG20, Gre22, NB22]. Set [CN23]. Sets [Ara21, FX21, PV20, TYM23, ZM21, vEM20]. Shaken [ADS^{+}22]. Shallow [MQ20]. Shape [CP20, DGZ20]. Shapes [FX21]. Sharp [BN22, LP22]. Sharpness [Har22]. Sherrington [ABvSY21, AC21, Ban20, DW21, IISS23, LL23a, LRRS21, Liu21]. Shielded [BDNS20]. Shift [Der23]. Shifts [Pin23, Tak20, Tak23]. Shock [JPR21, JR21]. Short [GLPP20, GKM21, GM21a, HS21, Sam20b, SY20]. Short-Distance [Sam20b]. Short-Range [GM21a, HS21, SY20]. Shortcuts [BPC22]. Shortest [CDD20]. Shows [AFKH^{+}20]. Sided [KL20b]. Signals [HM23]. Significance [Pie20]. Silico [Mou23]. similar [Kep23, SSS21]. Simple [AS21a, DHS21, DZW^{+}20, Eld20, Fia20, TN18, TN21]. Simulated [GLV21]. Simulation [DAT21, FLM22, NP21, WLB^{+}20]. Simulations [GJ23, HG20, JPR21, JR21, Pie20]. Sinai [Bru21]. Single [LS21]. Singular [SS22, TH22]. Singularity [JKNV23, KM20]. SIP [ACR21]. SIS [FIS20]. Site [HM20, LT21]. Site-Monotonicity [LT21]. Sites [EN21]. Size [GK22a, KK21]. Skew [MPS21]. Skew-Orthogonal [MPS21]. Slab [KLY21]. SLE [FLM22]. Slow [BdPGN20, DHS21, EKN20, EGK21, EGN20, FR20, NR21, XZ21]. Slow-to-Start [FR20]. Smale [BM20]. Small [BHP21, HJ23, JP22, Pie20, iSHNY22, Ste22]. Smoluchowski [CWZ20, Son20, TD20]. Smooth [KLvR20]. Smoothed [Sos23]. Smoothness [DK23a, KC21]. Sneppen [GS22, KV21, Vol22]. Sobolev [Den22, JLR23, LS20b, Pap20]. Social [Phi20]. Soft [BGM20, KK20b, Pec22, Qi21]. Soft-Core [KK20b]. Solenoidal [CDS21]. Solutes [ZRW20]. Solution [DRB20, KTT23, Qi21]. Solutions [AMSY23, Bob23a, BW21, Cri23, CS22, FF20, FLNV23, FPV22, JV22, JLY20, Kep23, LLM20, ZXJL23]. Solvation [ZRW20]. Solvent [Maj23]. Some [Bob22, CH22, DIM23, Czu22, FZ20, GL22a, GPS22, GLV21, HKR17, HKR18, HKR23b, HKR23c, Haf20, Meo23, NR21, SPL20]. SOS [CKL23, DMO20, Roz22]. SOS-Models [CKL23]. Source [CW22, LS21]. Southern [TCND20]. Space [BKL20, CTDN20, Den22, DGZ20, Hep20, KP20, MGPCA22, TCDN20, TCND20, Wu20, Jun20b]. Space-Time [MGPCA22, Jun20b]. Spaced [KK21, CW22]. Spaces [LW20, LS20b]. Span [Wie20]. Spanning [GS20]. Sparse [ABC23, DHO20, HNK22]. Spatial [MP22, vdHvdHM23]. Spatially [AKL20, ABT23, LL23b]. SPDEs [KKMS20]. Special [ACF^{+}20, GH20b, LT23, Luc20a]. Species [BBK^{+}23, LT23, MJB23, ACCM21, CLTC15, CGV23, DW21]. Spectral [BS13, BS20b, LJ20, Moh22, Seo20, Sha20]. Spectrum [BTV21, CJN20, Moh22]. Speed [FvMST23, LSW22, Ngu22, Ser20]. Sphere [GG22, JKNV23, JR21, SCDK23]. Spheres [NIB21, Urr22]. Spherical [ES22, BCS22, DS20, Kos23, LL23a, Sub22]. Spherically [CG22, Cha23]. Spiking [dSL23, Nas22, RNA23]. Spin [ACCM21, AC21, BS23, BCM20, BKN22, BCS22, BR22a, BR22b, BH20, CK22, CCR21, CKL23, DDD^{+}20, DS20, Eld20, Fac21a, Fac21b, FGK20, HS20b, HB21, Hik22, KLSS20, LT21, LW20, LRRS21, LS23, MW22, NR21, SS21, Sha20, SBB^{+}21, Sub22, SHRE23]. Spin-Flip [CCR21]. Spin-Glass [ACCM21, LRRS21]. Spins [DFP20]. Split [IT20]. Split-and-Merge [IT20]. Splitting [FLM22]. Spontaneous [BC22, Tas20]. Sprays [BDD23]. Spread [LS21]. Spreading [AAVF23]. SRB [JL22]. SSEP [EGN20, XZ21]. Stability [Ara21, BDD23, CD20b, CD20a, GV20, HR16, HR23, JLR23, LY21, MM21]. Stable [CLW23, KM20, ZD21]. Stack [WBC21]. Start [FR20]. State [CG20, CGJ21, CTDN20, HKP21, Iin21, JKG21, Men20, Oto23, THFH20, TCDN20, TCND20, VLC23]. Statements [WQ20]. States [ES22, BJN21, CK22, CGP20b, CMRU20, CKL23, DMO20, DK23a, EM20, GMRS20, HR16, HR23, HKR23a, Kos23, LLM20, Mor20b, MS21, RRA23, Rue20, Tak23, TTV23, WOK20, Wre21, YS20, ZW23, vEKM20, vEM20]. Static [AKL20, ZRW20, ZXD22, iSHNY22]. Stationary [BY20, BKL20, BGJOS22, Bha20, CGP20b, CT22, CCHS22, CS22, DMO20, EM20, FLNV23, IT20, JRA20, KLY21, KK22, Mor20a, PYZ20, BGM20, FFMV20]. Statistical [AFKH^{+}20, Ara21, BS20a, BVG21, BL21, CB20, DG20, Fac21a, Fac21b, FF20, Luc20a, Mae21, MQ20, Pie20, Rue20, Wre23]. Statistics [BS20c, DMS23, FN20, GL22b, HK22, NT20, SM23b, THFH20, Yak21, ZCD22]. Steady [JKG21, LLM20, Men20]. Steady-State [JKG21]. Steps [BRO22a, Ngu22]. Stick [Bro22b]. Sticky [ACR21]. stiff [BO21]. Stiffness [HXX23]. Stirring [IT20]. Stochastic [AKL20, BM21, BDS20, CTDN20, CT22, CQ21, CCH^{+}20b, Cri23, DHL20, DK23b, EKN20, EGK21, FGM23, FGK20, FNS21, Gao20, Gao22, GHL20, GV20, Gol23, Gre22, GJ23, HOiS22, HLW20, JLY20, JMPR23, KK20a, KW20, dSL23, LD22, ORD22, PR21, Son20, SD21, TCDN20, TCND20, TZQY20, VLC23, XX20, Zha20, ZD21, ZXJL23]. Stokes [ALT22, LLY22, FF20, Gao22, KR22c, LYZ23]. Stops [Ber22]. Strange [HM23]. Strategies [CLTC15, Nei21]. Stress [KO20]. Stress-Energy [KO20]. Strong [CB20, FL23, LS20c, LS20d, MP22, MS22b]. Strongly [BN22]. Structural [SEG22]. Structure [GJJ21, JCG20, SCDK23]. Structures [MM20a, PV23]. Stubborn [MMR20]. Studies [WLB^{+}20]. Study [CW22, NS21, Pie20, RNA23]. Studying [SBB^{+}21]. Sturmian [vEKM20]. Sub [HS22]. Sub-Gaussian [HS22]. Subcritical [AK21, CD20b]. Subdiffusion [DHO20, KG23]. Subgraphs [Ste22]. Subject [FK21]. Subjected [dZGFBC20]. Subordinated [KdS21]. Subpartition [CP20]. Subpartition-Maximizing [CP20]. Subsets [CLW23, Wu21]. Subshifts [BHL21]. Subspace [TTV23]. Success [CLTC15]. Sufficient [JK22]. Sufficiently [EM22]. Sunklodas [LS20a]. Super [LS20d]. Super-Critical [LS20d]. Superadditivity [Wre22]. Superconductors [BR22a, BR22b]. Superdiffusion [Doy22]. Supersymmetric [Cra21, LS23]. Sure [DHS23, PV23]. Surface [ABV22, Iin21, KBK23, Mor20a]. Surfaces [BKN22, Bru21]. Surveying [SEG22]. Survival [GMW20, MPR20]. Switching [FGR22]. Symmetric [AS21a, Bob23b, CG22, Cha23, DHS21, DW21, KP20, KC21, vGR20]. Symmetries [SGG23]. Symmetry [AAAB22, ABCM20, BC22, IISS23, Jon22, PT20, Roz22, Sam20b, Tas20]. Symmetry-Resolved [Jon22]. Symplectic [MPS21]. Synaptic [GLPP20]. Synchronization [ST22, SLZ23]. System [ADT21, ALT22, BVG21, CG23, CDW21, CQ21, DY21, FL23, FF20, FPT21, GLPP20, Gao20, GG22, HS20b, HLP20a, KK21, KL20b, dSL23, LLY22, LYZ23, MW20b, Nas22, NP20, Pie20, SM23b, WFKM^{+}20]. Systems [BHP21, BR22a, BR22b, BB20, BL22, CM20b, CGJ21, CGV23, CMRU20, CTDN20, CLM20, CLL21, CLY21, Ciu22, DS21, DP21b, EKN20, EGK21, FB22, FLNV23, FGR22, FL18, FL20a, FW21, Fre22, FFMV20, GL23a, GPS22, GHL20, GGGM23, HG20, JK23, JLY20, Kar20, KY23, LT21, LS20a, LD22, MM20a, MM23, MGM23, RNA23, iSHNY22, SSvB21, SY20, SS21, SEG22, SHRE23, THFH20, TCDN20, TCND20, YS20, ZV20, CM20a].

Tail [GK22a]. Tailed [HLN22]. Tails [Bob23b, GRM20, HS22]. TAP [ABvSY21]. TASEP [DP21a, FN20]. Taylor [BS20b, AESW21, BS13, CH22, DE22]. Technique [Mou23]. Techniques [KL20a, RB20]. Telegraph [PRD23]. Teller [WM22]. Temperature [CS22, GK22b, KR22a, KY23, NT20, Pin23, SCDK23, dZGFBC20]. Temperatures [EM22]. Tempering [BDS20]. Temporal [TCVB20, WOK20, XTL20]. Tendency [GN20]. Tensile [HXX23]. Tension [Mor20a]. Tensor [Fen21, HP20a, HKP21, Iin21, KR22a]. Term [CD20a, GLPP20]. Terms [Mor20b]. Ternary [dZGFBC20]. Tessellation [MPDH20]. Tessellations [BDR23, JT20, JT21]. Their [CS22, EK20a]. Theorem [ABC23, CGH22, Czu22, DHP20, EM22, Gal21, Oto23, Wat19, Wat20]. Theorems [AT21, CDN21, GL22c, Liu20, LS22b, MT20]. Theoretic [NS20]. Theories [CK22, CDS21, Abh23]. Theory [AS21a, BGSRS20, BB20, CTDN20, EK20a, EM20, GHO21, GBE22, JLM23, KTT23, KM20, KO20, LW20, NB22, PR20, PMACF20, TBD^{+}20, Wre23, ZD21, ZRW20]. Thermal [BR22a, BR22b, DHP20, Gou23, HS21, MGPCA20, Mor20b]. Thermalisation [SHRE23]. Thermalization [BHP21, GAD20, Lep23]. Thermo [IOS21, Oto23]. Thermo-hydrodynamics [Oto23]. Thermo-mechanical [IOS21]. Thermodynamic [BH22, BPC22, GJJ21, HPRS20, LLV22, Lyc20, MB21, NS20, NS21, NS22, iSHNY22, ST22, Shi21, Wre22, Xu21, Zha20]. Thermodynamics [BL22, Jan20, YQ21, Jan21]. Thermostated [CT21]. Thick [BDD23]. Thompson [JRS20]. Thouless [AC21, MW22]. Three [AS21b, CLTC15, DAN20, EK20b, FN20, FL18, FL20a, KR22c, SPL20, WM22]. Three-Dimensional [WM22]. Three-Layer [AS21b]. Three-Parametric [EK20b]. Three-Towers [SPL20]. Threshold [GZ20, PR21, Shc20]. Thresholds [Cha23, JMPR23]. Tightness [NPV23]. tigrina [CLTC15]. Time [ADT21, AESW21, ALP21, BPP20, BJN21, Bha20, Bód20, CFVY20, CT22, CCS20, CSZ21, DGZ20, FIS20, GMV20, GN21, GG20, Gre22, GM21b, Hep20, HLN22, HM23, KP20, LS20a, LXZ23a, LXZ23b, MGPCA20, MGPCA22, NF20a, Pen20, RB20, RNA23, SS23, TTV23, XX20, Jun20b, Pel21]. Time-Dependent [LS20a, XX20]. Time-Inhomogeneous [GN21, LXZ23a, LXZ23b]. Time-Non-Local [ALP21]. Time-Series [CT22]. Time-Travels [Pel21]. Times [Meo23]. Tipping [AvdH20]. Tissue [DZW^{+}20]. Toda [BOP23, LL20, Spo20]. Toeplitz [BEL22]. Tools [AS21a, SBB^{+}21]. Toom [CK20]. Topological [BR22a, BR22b, SN22, TYM23, ZM21]. Topologies [Mad23]. Torus [IT20]. Total [GH20b, NPV23]. Towers [SPL20]. Traces [FLM22]. Tracy [MPS21]. Traffic [DT20, DTZ22, FR20]. Trajectorial [JK23]. Trajectories [BGN21, DGT^{+}20]. Transfer [NT23, Wan22]. Transform [Mol22]. transformation [Abh23]. Transformations [HS20a, JL22]. Transient [AAVF23, Gat23]. Transition [BJN21, BS20d, CN23, Der23, GMV20, KTT22, TN18, TN21]. Transitions [GR20, GG20, KK20b, KM21, LXZ23a, LXZ23b, MQ20, Mor20a, MS22a, Nas22, NRS20, TR22]. Transitive [FIS20]. Translation [SHRE23]. Transmission [GS20]. Transport [AOdLHT20, BCD22, BR22a, BR22b, CGP20a, CN21, CM20a, CM20b, CCH^{+}20b, DHL20, FS20a, GM21a, HG20, IOS21, JVG21, KTT22, Wir22]. Transverse [CG20, IISS23]. Trap [GK22a]. Traps [Gou23]. Traveling [KdS21]. Travels [Pel21]. Tree [ABB20, BdBBR23, CDHM23, GMRS20, RRA23, dHW22]. Trees [CDHR23, CKL23, DM22, FvdH21, FvdH22, GP23, JMPR23, JRS20, KMN22, KZ20, MPR20, MS21, Roz22]. Triangle [Hut22]. Triangles [GGM21]. Trick [BS23]. Tricritical [Iin21]. Trimodal [MS22a]. Triviality [BCS22]. Truncated [SM23b]. Tsingou [Gol23]. Tube [BGN21, WY22]. Tumble [vGvGR21]. Tumor [FGR20]. Tunneled [MM23]. Turbulence [DK23b, GBE22]. Twenty [DDG20]. Twenty-Vertex [DDG20]. Two [AAVF23, ABE23, BS20a, BPP20, Bla21, BBK^{+}23, CCS21, CLLL23, Cra21, DAT21, EM22, FK21, GHL20, GH20b, GL22c, JR21, KTT23, KL20a, KR22c, LT23, MJB23, MGPCA20, NS22, Sam20b, SKM20, SPL20, THFH20, WLB^{+}20]. Two-Body [EM22, KTT23]. Two-Component [WLB^{+}20]. Two-Dimensional [AAVF23, ABE23, BS20a, Bla21, DAT21, GL22c, Sam20b]. Two-Particle [CLLL23]. Two-Point [FK21]. Two-Species [BBK^{+}23, LT23, MJB23]. Two-State [THFH20]. Two-Time [BPP20]. Two-Vertex [SPL20]. Type [BM20, DT20, DTZ22, GV20, Hik22, KTT22, LW20, MT20, PKTM23, PN23, GN21]. Typical [BGN21, MPDH20, TTV23, vEM20].

Uhlenbeck [BHP21, GMW20]. Ulam [Gol23]. Unbounded [RZ20, RZ22]. Uncertainty [NS20, NS21, NS22, Shi21, VPG20]. Underdamped [TZQY20]. Unfolded [MM22]. Unified [SS21, ZCD22]. Uniform [BW20, BCD22, CG20, CT21, GM21b, PR21, Ste22, HXX23]. uniformity [KY23]. Uniformly [MPDH20]. Unimodality [MSU20]. Uniqueness [BdBBR23, CMRU20, JRA20, Tak20]. unital [FGM23]. Universal [DMS23, JLM23, LS23, NPT20]. Universality [ABE23, CW22, IISS23]. Unstable [NT23, Wre21]. Uphill [CGV23, FGR22]. Upper [KD20, TYM23, Vol22, ZM21]. Use [ABL21]. Using [BGH20, DAT21, GHO21, GG20, GL20, KR22b, KL20a, Lau22, MM20a, Mou23, NB22].

Value [PMACF20, SS22]. Valued [BCS20, DHS23, KM20, Qi21, Wre22]. Values [RB20]. Vanishing [GMV20, JKNV23, Mas22]. Variability [Pie20]. Variables [God21]. Variance [ABE23, CHR22, Jan22, Unt22]. Variant [KW20]. Variational [CLM20, LT20, Mor20b]. Varying [CWZ20, DT21]. Vector [DHS23, KM20]. Vector-Valued [DHS23, KM20]. Vehicles [DTZ22]. Velocity [ABT23, Các20b, CCHS22, DIM23, HHW21, KD20, Các20a]. Velocity-Dependent [HHW21]. Vergni [Pel21]. Version [KV21]. Versus [KR22c]. Vertex [DDG20, FNS21, SPL20]. Via [MJB23, BL21, CK20, EGK21, FLM22, KM20, LLV22, PMACF20, Sco21]. Vibrational [MM20b]. Viewpoint [GL22a, KTT23]. Violation [EN21, Tan22]. Viral [SBB^{+}21]. Virial [NF20b, Urr22]. Viscosity [Gal21]. Vlasov [LLY22, ADT21, BW20, CDW21, CKL20, CLLL23, DY21, EK20a, FL23, HLP20a, NP20, TTZ22]. Voltages [CHR22]. Volume [CGP20b, CLV22, Kos23]. Voronoi [GR22, MPDH20]. Vorticity [YQ21]. Voter [Bla21, MMR20, RF21]. Vries [Gao20]. Vulpiani [Pel21].

Waals [Mor20a, TN21, TN18]. Walk [Ber22, CDN21, EFR23, HS20b, HSS20, KK22, KKS23, Lau22, LP22, MT20, Pel21, ZXD22]. Walk/Zeta [KKS23]. Walkers [AJR21, LS21]. Walks [BR21, BK21, BBO21, Bau20, BRO22a, CGH22, CGL21, DMS23, FIS20, GL22b, GN20, GNH21, LS22b, Ngu22, SH23]. Wall [DDG20, IOVW20]. Walls [Sam20a]. Wasserstein [BHP21, BL21]. Water [MQ20, ZRW20]. Watson [DM22, MPR20, dHW22]. Wave [Các21, DK23b, SM23b, ZCD22]. Wave-Like [Các21]. Waves [KdS21, KR22c, MQ20, SCDK23]. Way [ADS^{+}22]. Weak [AS21a, BTV21, CCM20, GBE22, LRRS21, Sak21, WY22]. Weakly [CLY21, Cri23, LT20, Men20, RS22]. Wealth [ZL22]. Weather [AOdLHT20]. Web [PPR22, GS20, WZN20]. Weber [Mae21]. Wedge [Rod21]. Weight [Mas22]. Weighted [AMSY23, Den22, LS20b, RCF20]. Weiss [KM20, KM21, Lee22, Ban20, CJN22, LW20]. Weisskopf [JM23]. Well [CKL20, PYZ20]. Well-Posedness [CKL20]. Weyl [HKR22]. Whittaker [JO20]. Widom [KK20b, MPS21]. Wiener [XX20]. Wigner [JM23, LL23a]. Wind [ABB20]. Wind-Tree [ABB20]. Without [CGT21, Tak23]. Work [NB22, SD21]. Worlds [JP22].

XXZ [DDD^{+}20].

Y-Junctions [SN22]. Yang [HJN23].

Zebiak [TCND20]. Zeitlin [FPV22]. Zero [BGJOS22, BGN21, CS22, Kim21, Pin23, SGG23, SS21, SCDK23, CGP20b]. Zero-Freeness [SS21]. Zero-Range [BGJOS22]. Zero-Temperature [CS22]. Zeros [HKR22, HJN23]. Zeta [KKS23]. Zhang [BKL20, CCM20, DTZ22, KP20]. Zhang-Type [DTZ22].

References

Albanese:2022:RSB

[AAAB22] Linda Albanese, Francesco Alemanno, Andrea Alessandrelli, and Adriano Barra. Replica symmetry breaking in dense Hebbian neural networks. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-02966-8.

Acharyya:2023:TBD

[AAVF23] Ajanta Bhowal Acharyya, Muktish Acharyya, Erol Vatansever, and Nikolaos G. Fytas. Transient behavior of damage spreading in the two-dimensional Blume–Capel ferromagnet. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-03012-3.

Attarchi:2020:EWT

[ABB20] Hassan Attarchi, Mark Bolding, and Leonid A. Bunimovich. Ehrenfests’ wind-tree model is dynamically richer than the Lorentz gas. Journal of Statistical Physics, 180(1–6):440–458, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Amini:2023:CLT

[ABC23] Hamed Amini, Erhan Bayraktar, and Suman Chakraborty. A central limit theorem for diffusion in sparse random graphs. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-023-
03068-9.

Alberici:2020:ARS

[ABCM20] Diego Alberici, Adriano Barra, Pierluigi Contucci, and Emanuele Mingione. Annealing and replica-symmetry in deep Boltzmann machines. Journal of Statistical Physics, 180(1–6):665–677, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Akemann:2023:UNV

[ABE23] Gernot Akemann, Sung-Soo Byun, and Markus Ebke. Universality of the number variance in rotational invariant twodimensional Coulomb gases. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03005-2.

Abhignan:2023:EHF

[Abh23] Venkat Abhignan. Extrapolation from hypergeometric functions, continued functions and Borel–Leroy transformation; resummation of perturbative renormalization functions from field theories. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03108-4.

Alonso:2021:AUE

[ABL21] R. Alonso, V. Bagland, and B. Lods. About the use of entropy production for the Landau–Fermi–Dirac equation. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02751-z.

Ayot:2023:VAC

[ABT23] Valentin Ayot, Stéphane Brull, and Philippe Thieullen. A velocity alignment collision model for spatially homogeneous kinetic collective dynamics. Journal of Statistical Physics, 190 (2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03008-z.

Arezzo:2022:LDN

[ABV22] Claudio Arezzo, Federico Balducci, and Carlo Vanoni. Localization in the discrete non-linear Schrödinger equation and geometric properties of the microcanonical surface. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02870-7.

Adhikari:2021:DAT

[ABvSY21] Arka Adhikari, Christian Brennecke, Per von Soosten, and Horng-Tzer Yau. Dynamical approach to the TAP equations for the Sherrington–Kirkpatrick model. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSBo. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-021-02773-7.

Auffinger:2021:TAP

[AC21] Antonio Auffinger and Cathy Xi Chen. Thouless–Anderson– Palmer equations for the Ghatak–Sherrington mean field spin glass model. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02803-4.

Alberici:2021:MSM

[ACCM21] Diego Alberici, Francesco Camilli, Pierluigi Contucci, and Emanuele Mingione. The multi-species mean-field spin-glass on the Nishimori line. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aizenman:2020:ISI

[ACF^{+}20] Michael Aizenman, Ivan Corwin, Jürg Fröhlich, Giovanni Gallavotti, Shelly Goldstein, and Herbert Spohn. Introduction to the special issue in honor of Joel Lebowitz. Journal of Statistical Physics, 180(1–6):1–3, September 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02606-z.pdf.

Apolinario:2022:DFM

[ACM22] Gabriel B. Apolinário, Laurent Chevillard, and Jean-Christophe Mourrat. Dynamical fractional and multifractal fields. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02867-2.

Ayala:2021:CSP

[ACR21] Mario Ayala, Gioia Carinci, and Frank Redig. Condensation of SIP particles and sticky Brownian motion. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-021-02775-5;
 http://link.springer.com/content/pdf/10.1007/s10955-
021-02775-5.pdf.

Apollonio:2022:SDE

[ADS^{+}22] Valentina Apollonio, Roberto D’Autilia, Benedetto Scoppola, Elisabetta Scoppola, and Alessio Troiani. Shaken dynamics: an easy way to parallel Markov chain Monte Carlo. Journal of Statistical Physics, 189(3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-03004-3.

Addala:2021:TLH

[ADT21] Lanoir Addala, Jean Dolbeault, and M. Lazhar Tayeb. L^{2}- hypocoercivity and large time asymptotics of the linearized Vlasov–Poisson–Fokker–Planck system. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02784-4.

Arnold:2021:LTC

[AESW21] Anton Arnold, Amit Einav, Beatrice Signorello, and Tobias Wöhrer. Large time convergence of the non-homogeneous Goldstein–Taylor equation. Journal of Statistical Physics, 182 (2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-021-02702-8.pdf.

Ahn:2020:ASM

[AFKH^{+}20] Seonmin Ahn, Baylor Fox-Kemper, Timothy Herbert, Mark Altabet, and Charles E. Lawrence. Autoregressive statistical modeling of a Peru margin multi-proxy Holocene record shows correlation not causation, flickering regimes and persistence. Journal of Statistical Physics, 179(5–6):1553–1571, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Aizenman:2020:EDC

[AHP20] Michael Aizenman, Matan Harel, and Ron Peled. Exponential decay of correlations in the 2D random field Ising model. Journal of Statistical Physics, 180(1–6):304–331, September 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Andreys:2021:FWD

[AJR21] Simon Andréys, Alain Joye, and Renaud Raquépas. Fermionic walkers driven out of equilibrium. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02795-1.

Angel:2021:LDS

[AK21] Omer Angel and Brett Kolesnik. Large deviations for subcritical bootstrap percolation on the Erdős–Rényi graph. Journal of Statistical Physics, 185(2):??, November 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02819-w.

Antonov:2020:SAR

[AKL20] N. V. Antonov, P. I. Kakin, and N. M. Lebedev. Static approach to renormalization group analysis of stochastic models with spatially quenched noise. Journal of Statistical Physics, 178(2):392– 419, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ascione:2021:TNL

[ALP21] Giacomo Ascione, Nikolai Leonenko, and Enrica Pirozzi. Time-non-local Pearson diffusions. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-021-02786-2;
 http://link.springer.com/content/pdf/10.1007/s10955-
021-02786-2.pdf.

Alonso:2022:BEG

[ALT22] Ricardo J. Alonso, Bertrand Lods, and Isabelle Tristani. From Boltzmann equation for granular gases to a modified Navier– Stokes–Fourier system. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02915-5.

Aleandri:2021:DIP

[AM21] Michele Aleandri and Ida G. Minelli. Delay-induced periodic behaviour in competitive populations. Journal of Statistical Physics, 185(1):??, October 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02820-3.

Alonso:2023:GAW

[AMSY23] R. Alonso, Y. Morimoto, W. Sun, and T. Yang. De Giorgi argument for weighted L^{2} ∩ L^{∞} solutions to the non-cutoff Boltzmann equation. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03053-8.

Alonso-Oran:2020:MCW

[AOdLHT20] Diego Alonso-Orán, Aythami Bethencourt de León, Darryl D. Holm, and So Takao. Modelling the climate and weather of a 2D Lagrangian–averaged Euler–Boussinesq equation with transport noise. Journal of Statistical Physics, 179(5–6):1267–1303, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02443-9.pdf.

Araujo:2021:SSF

[Ara21] Vitor Araujo. On the statistical stability of families of attracting sets and the contracting Lorenz attractor. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Abhignan:2021:CFP

[AS21a] Venkat Abhignan and R. Sankaranarayanan. Continued functions and perturbation series: Simple tools for convergence of diverging series in O(n)-symmetric ϕ^{4} field theory at weak coupling limit. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02719-z.

Agliari:2021:LRO

[AS21b] Elena Agliari and Giulia Sebastiani. Learning and retrieval operational modes for three-layer restricted Boltzmann machines. Journal of Statistical Physics, 185(2):??, November 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02841-y.

Assiotis:2022:MPF

[Ass22] Theodoros Assiotis. On the moments of the partition function of the CβE field. Journal of Statistical Physics, 187(2):??, May 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02906-6.

Andres:2021:LLT

[AT21] Sebastian Andres and Peter A. Taylor. Local limit theorems for the random conductance model and applications to the Ginzburg–Landau ∇ϕ interface model. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-021-02705-5.
pdf.

Ashwin:2020:ESC

[AvdH20] Peter Ashwin and Anna S. von der Heydt. Extreme sensitivity and climate tipping points. Journal of Statistical Physics, 179 (5–6):1531–1552, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-019-02425-x.pdf.

Banerjee:2020:FFE

[Ban20] Debapratim Banerjee. Fluctuation of the free energy of Sherrington–Kirkpatrick model with Curie–Weiss interaction: The paramagnetic regime. Journal of Statistical Physics, 178 (1):211–246, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Baur:2020:CRW

[Bau20] Erich Baur. On a class of random walks with reinforced memory. Journal of Statistical Physics, 181(3):772–802, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bougron:2020:LRT

[BB20] Jean-François Bougron and Laurent Bruneau. Linear response theory and entropic fluctuations in repeated interaction quantum systems. Journal of Statistical Physics, 181(5):1636–1677, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Beck:2022:GCE

[BB22] Justin Beck and Federico Bonetto. Grand canonical evolution for the Kac model. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02932-4.

Barre:2020:FKF

[BBC^{+}20] J. Barré, C. Bernardin, R. Chétrite, Y. Chopra, and M. Mariani. From fluctuating kinetics to fluctuating hydrodynamics: a Γ-convergence of large deviations functionals approach. Journal of Statistical Physics, 180(1–6):1095–1127, September 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Blyschak:2023:OPD

[BBK^{+}23] Danyil Blyschak, Olivia Burke, Jeffrey Kuan, Dennis Li, Sasha Ustilovsky, and Zhengye Zhou. Orthogonal polynomial duality of a two-species asymmetric exclusion process. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03100-y.

Basile:2021:LDK

[BBO21] Giada Basile, Dario Benedetto, and Carlo Orrieri. Large deviations for Kac-like walks. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02794-2.

Benedetto:2020:ERM

[BC20] Dario Benedetto and Emanuele Caglioti. Euclidean random matching in 2D for non-constant densities. Journal of Statistical Physics, 181(3):854–869, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Buffoni:2022:SFS

[BC22] Lorenzo Buffoni and Michele Campisi. Spontaneous fluctuationsymmetry breaking and the Landauer principle. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02877-8.

Bhat:2022:HTO

[BCD22] Junaid Majeed Bhat, Gaëtan Cane, and Abhishek Dhar. Heat transport in an ordered harmonic chain in presence of a uniform magnetic field. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02848-5.

Benoist:2021:EPR

[BCJP21] T. Benoist, N. Cuneo, V. Jaksić, and C.-A. Pillet. On entropy production of repeated quantum measurements II. examples. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Barbier:2020:CMO

[BCM20] Jean Barbier, Chun Lam Chan, and Nicolas Macris. Concentration of multi-overlaps for random dilute ferromagnetic spin models. Journal of Statistical Physics, 180(1–6):534–557, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bronski:2020:MVK

[BCS20] Jared C. Bronski, Thomas E. Carty, and Sarah E. Simpson. A matrix-valued Kuramoto model. Journal of Statistical Physics, 178(2):595–624, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Belius:2022:TGM

[BCS22] David Belius, Jirí Cerný, and Marius A. Schmidt. Triviality of the geometry of mixed p-spin spherical Hamiltonians with external field. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02855-6.

Bencs:2023:UGM

[BdBBR23] Ferenc Bencs, David de Boer, Pjotr Buys, and Guus Regts. Uniqueness of the Gibbs measure for the anti-ferromagnetic Potts model on the infinite Δ-regular tree for large Δ. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB.

ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-023-03145-z.

Buet:2023:LST

[BDD23] C. Buet, B. Després, and L. Desvillettes. Linear stability of thick sprays equations. Journal of Statistical Physics, 190(3): ??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-03057-4.

Bock:2020:PFC

[BDNS20] Bounghun Bock, Michael Damron, C. M. Newman, and Vladas Sidoravicius. Percolation of finite clusters and shielded paths. Journal of Statistical Physics, 179(3):789–807, May 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bonorino:2020:HPM

[BdPGN20] L. Bonorino, R. de Paula, P. Gonçalves, and A. Neumann. Hydrodynamics of porous medium model with slow reservoirs. Journal of Statistical Physics, 179(3):748–788, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Balint:2023:EMP

[BDR23] Péter Bálint, Gábor Domokos, and Krisztina Regős. An evolution model for polygonal tessellations as models for crack networks and other natural patterns. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-023-03146-y.

Boniece:2020:FLP

[BDS20] B. Cooper Boniece, Gustavo Didier, and Farzad Sabzikar. On fractional Lévy processes: Tempering, sample path properties and stochastic integration. Journal of Statistical Physics, 178 (4):954–985, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Basor:2022:ABT

[BEL22] Estelle Basor, Torsten Ehrhardt, and Yuqi Li. Asymptotics of bordered Toeplitz determinants and next-to-diagonal Ising correlations. Journal of Statistical Physics, 187(1):??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02894-7.

BenPorat:2023:DEE

[Ben23] Immanuel Ben Porat. Derivation of Euler’s equations of perfect fluids from von Neumann’s equation with magnetic field. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-023-03131-5.

Bercu:2022:ERW

[Ber22] Bernard Bercu. On the elephant random walk with stops playing hide and seek with the Mittag-Leffler distribution. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02980-w.

Boccagna:2020:RIM

[BG20] Roberto Boccagna and Davide Gabrielli. Remarks on the interpolation method. Journal of Statistical Physics, 181(4):1218– 1238, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02624-x.pdf.

Bricmont:2020:SPP

[BGH20] Jean Bricmont, Sheldon Goldstein, and Douglas Hemmick. Schrödinger’s Paradox and proofs of nonlocality using only perfect correlations. Journal of Statistical Physics, 180(1–6):74–91, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bernardin:2022:NES

[BGJOS22] Cédric Bernardin, Patrícia Gonçalves, Byron Jiménez-Oviedo, and Stefano Scotta. Non-equilibrium stationary properties of the boundary driven zero-range process with long jumps. Journal of Statistical Physics, 189(3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02987-3.

Bianchi:2020:SCQ

[BGM20] A. Bianchi, A. Gaudillière, and P. Milanesi. On soft capacities, quasi-stationary distributions and the pathwise approach to metastability. Journal of Statistical Physics, 181(3):1052–1086, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bet:2021:CCT

[BGN21] Gianmarco Bet, Anna Gallo, and Francesca R. Nardi. Critical configurations and tube of typical trajectories for the Potts and Ising models with zero external field. Journal of Statistical Physics, 184(3):??, September 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02814-1.

Bodineau:2020:FTB

[BGSRS20] Thierry Bodineau, Isabelle Gallagher, Laure Saint-Raymond, and Sergio Simonella. Fluctuation theory in the Boltzmann-grad limit. Journal of Statistical Physics, 180(1–6):873–895, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Brezin:2020:PSC

[BH20] E. Brézin and S. Hikami. Punctures and p-spin curves from matrix models. Journal of Statistical Physics, 180(1–6):1031– 1060, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02581-5.pdf.

Barreira:2022:HDN

[BH22] Luis Barreira and Carllos Holanda. Higher-dimensional nonlinear thermodynamic formalism. Journal of Statistical Physics, 187(2):??, May 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02916-4.

Bhatia:2020:MDE

[Bha20] Manan Bhatia. Moderate deviation and exit time estimates for stationary last passage percolation. Journal of Statistical Physics, 181(4):1410–1432, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ban:2021:EMM

[BHL21] Jung-Chao Ban, Wen-Guei Hu, and Guan-Yu Lai. On the entropy of multidimensional multiplicative integer subshifts. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ban:2023:LDP

[BHL23] Jung-Chao Ban, Wen-Guei Hu, and Guan-Yu Lai. Large deviation principle of nonconventional ergodic averages. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-023-03073-y.

Barrera:2021:CTO

[BHP21] G. Barrera, M. A. Högele, and J. C. Pardo. Cutoff thermalization for Ornstein–Uhlenbeck systems with small Lévy noise in the Wasserstein distance. Journal of Statistical Physics, 184 (3):??, September 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02815-0.

Baldassarri:2023:MKD

[BJ23] Simone Baldassarri and Vanessa Jacquier. Metastability for Kawasaki dynamics on the hexagonal lattice. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-03061-8.

Bet:2021:EED

[BJN21] Gianmarco Bet, Vanessa Jacquier, and Francesca R. Nardi. Effect of energy degeneracy on the transition time for a series of metastable states. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02788-0.

Bailey:2021:MMB

[BK21] E. C. Bailey and J. P. Keating. Moments of moments and branching random walks. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02696-9.pdf.

Barraquand:2020:HSS

[BKL20] Guillaume Barraquand, Alexandre Krajenbrink, and Pierre Le Doussal. Half-space stationary Kardar–Parisi–Zhang equation. Journal of Statistical Physics, 181(4):1149–1203, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02622-z.pdf.

Betermin:2020:NCA

[BKN20] Laurent Bétermin, Hans Knüpfer, and Florian Nolte. Note on crystallization for alternating particle chains. Journal of Statistical Physics, 181(3):803–815, November 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02603-2.pdf.

Baskerville:2022:SGM

[BKN22] Nicholas P. Baskerville, Jonathan P. Keating, and Joseph Najnudel. A spin glass model for the loss surfaces of generative adversarial networks. Journal of Statistical Physics, 186 (2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02875-w.

Becker:2021:QSL

[BL21] Simon Becker and Wuchen Li. Quantum statistical learning via quantum Wasserstein natural gradient. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-020-02682-1.
pdf.

Bouley:2022:TND

[BL22] Angèle Bouley and Claudio Landim. Thermodynamics of nonequilibrium driven diffusive systems in mild contact with boundary reservoirs. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-02948-w.

Blanquicett:2021:FTD

[Bla21] Daniel Blanquicett. Fixation for two-dimensional U-Ising and Uvoter dynamics. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Butta:2020:CST

[BM20] Paolo Buttà and Carlo Marchioro. Cucker–Smale type dynamics of infinitely many individuals with repulsive forces. Journal of Statistical Physics, 181(6):2094–2108, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bianchi:2021:LSD

[BM21] Luigi Amedeo Bianchi and Francesco Morandin. Linear stochastic dyadic model. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02753-x.

Baldassarri:2022:CDS

[BN22] Simone Baldassarri and Francesca R. Nardi. Critical droplets and sharp asymptotics for Kawasaki dynamics with strongly anisotropic interactions. Journal of Statistical Physics, 186(3): ??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-02874-x.

Bradly:2021:ELI

[BO21] C. J. Bradly and A. L. Owczarek. Effect of lattice inhomogeneity on collapsed phases of semi-stiff ISAW polymers. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bobylev:2022:SPG

[Bob22] A. V. Bobylev. On some properties of generalized Burnett equations of hydrodynamics. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02927-1.

Bobylev:2023:SMB

[Bob23a] A. V. Bobylev. On solutions of the modified Boltzmann equation. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03010-5.

Bobylev:2023:RSM

[Bob23b] A. V. Bobylev. Radially symmetric models of the Landau kinetic equation and high energy tails. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03058-3.

Bodai:2020:EAE

[Bód20] Tamás Bódai. An efficient algorithm to estimate the potential barrier height from noise-induced escape time data. Journal of Statistical Physics, 179(5–6):1625–1636, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02574-4.pdf.

Benettin:2023:RIT

[BOP23] Giancarlo Benettin, Giuseppe Orsatti, and Antonio Ponno. On the role of the integrable Toda model in one-dimensional molecular dynamics. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03147-x.

Bouchet:2020:BER

[Bou20] Freddy Bouchet. Is the Boltzmann equation reversible? a large deviation perspective on the irreversibility paradox. Journal of Statistical Physics, 181(2):515–550, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Boyd:2022:STC

[BPC22] Alexander B. Boyd, Ayoti Patra, and James P. Crutchfield. Shortcuts to thermodynamic computing: The cost of fast and faithful information processing. Journal of Statistical Physics, 187(2):??, May 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02871-0.

Benoist:2020:HCF

[BPP20] T. Benoist, A. Panati, and Y. Pautrat. Heat conservation and fluctuations between quantum reservoirs in the two-time measurement picture. Journal of Statistical Physics, 178(4):893–925, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bossmann:2020:HOC

[BPPS20] Lea Boßmann, Natasa Pavlović, Peter Pickl, and Avy Soffer. Higher order corrections to the mean-field description of the dynamics of interacting bosons. Journal of Statistical Physics, 178 (6):1362–1396, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02500-8.pdf.

Biswas:2022:MEA

[BPR22] Apurba Biswas, V. V. Prasad, and R. Rajesh. Mpemba effect in anisotropically driven inelastic Maxwell gases. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02891-w.

Bai:2021:BRW

[BR21] Tianyi Bai and Pierre Rousselin. Branching random walks conditioned on particle numbers. Journal of Statistical Physics, 185(3):??, December 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02833-y.

Bondyopadhaya:2022:CNE

[BR22a] Nilanjan Bondyopadhaya and Dibyendu Roy. Correction to: Nonequilibrium Electrical, Thermal and Spin Transport in Open Quantum Systems of Topological Superconductors, Semiconductors and Metals. Journal of Statistical Physics, 187(2):??, May 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02907-5. See [BR22b].

Bondyopadhaya:2022:NET

[BR22b] Nilanjan Bondyopadhaya and Dibyendu Roy. Nonequilibrium electrical, thermal and spin transport in open quantum systems of topological superconductors, semiconductors and metals. Journal of Statistical Physics, 187(2):??, May 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02902-w. See correction [BR22a].

Bouchet:2022:PID

[BR22c] Freddy Bouchet and Julien Reygner. Path integral derivation and numerical computation of large deviation prefactors for non-equilibrium dynamics through matrix Riccati equations. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02983-7.

Bertenghi:2022:JIP

[BRO22a] Marco Bertenghi and Alejandro Rosales-Ortiz. Joint invariance principles for random walks with positively and negatively reinforced steps. Journal of Statistical Physics, 189(3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02993-5.

Broman:2022:HDS

[Bro22b] Erik I. Broman. Higher-dimensional stick percolation. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02849-4.

Bruin:2021:SBF

[Bru21] Henk Bruin. On Sinai billiards on flat surfaces with horns. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-021-
02746-w.

Bernard:2013:OES

[BS13] Étienne Bernard and Francesco Salvarani. Optimal estimate of the spectral gap for the degenerate Goldstein–Taylor model. Journal of Statistical Physics, 153(2):363–375, October 2013. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/article/10.1007/

s10955-013-0825-6. See correction [BS20b].

Bahri:2020:SMF

[BS20a] M. E. H. Bahri and Y. Sinai. Statistical mechanics of freely fluctuating two-dimensional elastic crystals. Journal of Statistical Physics, 180(1–6):739–748, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bernard:2020:COE

[BS20b] Étienne Bernard and Francesco Salvarani. Correction to: Optimal Estimate of the Spectral Gap for the Degenerate Goldstein– Taylor Model. Journal of Statistical Physics, 181(4):1470–1471, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL http://link.springer.com/
content/pdf/10.1007/s10955-020-02631-y.pdf. See [BS13].

Berry:2020:GPC

[BS20c] M. V. Berry and Pragya Shukla. Geometric phase curvature statistics. Journal of Statistical Physics, 180(1–6):297–303, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Budzynski:2020:ACT

[BS20d] Louise Budzynski and Guilhem Semerjian. The asymptotics of the clustering transition for random constraint satisfaction problems. Journal of Statistical Physics, 181(5):1490–1522, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Baldwin:2023:RRT

[BS23] Christopher L. Baldwin and Brian Swingle. Revisiting the replica trick: Competition between spin glass and conventional order. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03135-1.

Borisov:2021:QHW

[BTV21] Denis Borisov, Matthias Täufer, and Ivan Veselić. Quantum Hamiltonians with weak random abstract perturbation. II. Localization in the expanded spectrum. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Baldovin:2021:SMI

[BVG21] Marco Baldovin, Angelo Vulpiani, and Giacomo Gradenigo. Statistical mechanics of an integrable system. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/article/10.1007/s10955-021-02781-7.

Bedrossian:2020:LVV

[BW20] Jacob Bedrossian and Fei Wang. The linearized Vlasov and Vlasov–Fokker–Planck equations in a uniform magnetic field. Journal of Statistical Physics, 178(2):552–594, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Bronski:2021:PPL

[BW21] Jared C. Bronski and Lan Wang. Partially phase-locked solutions to the Kuramoto model. Journal of Statistical Physics, 183(3): ??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-021-02783-5.

Bae:2020:SQB

[BY20] Gi-Chan Bae and Seok-Bae Yun. Stationary quantum BGK model for bosons and fermions in a bounded interval. Journal of Statistical Physics, 178(4):845–868, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Caceres:2020:CFV

[Các20a] Manuel O. Cáceres. Correction to: Finite-Velocity Diffusion in Random Media. Journal of Statistical Physics, 181(3): 1087, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02616-x.pdf. See

[Các20b].

Caceres:2020:FVD

[Các20b] Manuel O. Cáceres. Finite-velocity diffusion in random media. Journal of Statistical Physics, 179(3):729–747, May 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [Các20a].

Caceres:2021:CWL

[Các21] Manuel O. Cáceres. Comments on wave-like propagation with binary disorder. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cao:2021:AAB

[Cao21] Fei Cao. K-averaging agent-based model: Propagation of chaos and convergence to equilibrium. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02807-0.

Conti:2020:SMS

[CB20] Giovanni Conti and Gualtiero Badin. Statistical measures and selective decay principle for generalized Euler dynamics: Formulation and application to the formation of strong fronts. Journal of Statistical Physics, 179(5–6):1046–1072, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chakrabarty:2020:EOB

[CCH20a] Arijit Chakrabarty, Sukrit Chakraborty, and Rajat Subhra Hazra. Eigenvalues outside the bulk of inhomogeneous Erdős– Rényi random graphs. Journal of Statistical Physics, 181(5): 1746–1780, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cotter:2020:DAQ

[CCH^{+}20b] Colin Cotter, Dan Crisan, Darryl Holm, Wei Pan, and Igor Shevchenko. Data assimilation for a quasi-geostrophic model with circulation-preserving stochastic transport noise. Journal of Statistical Physics, 179(5–6):1186–1221, June 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02524-0.pdf.

Chen:2022:RVA

[CCHS22] I.-Kun Chen, Ping-Han Chuang, Chun-Hsiung Hsia, and Jhe-Kuan Su. A revisit of the velocity averaging lemma: On the regularity of stationary Boltzmann equation in a bounded convex domain. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02977-5.

Comets:2020:RKP

[CCM20] Francis Comets, Clément Cosco, and Chiranjib Mukherjee. Renormalizing the Kardar–Parisi–Zhang equation in d ≥ 3 in weak disorder. Journal of Statistical Physics, 179(3):713–728, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chazottes:2021:ECU

[CCR21] Jean-René Chazottes, Pierre Collet, and Frank Redig. Evolution of concentration under lattice spin-flip dynamics. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02796-0.

Comtet:2020:LPT

[CCS20] Alain Comtet, Françoise Cornu, and Grégory Schehr. Lastpassage time for linear diffusions and application to the emptying time of a box. Journal of Statistical Physics, 181(5):1565–1602, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Caraci:2021:BEC

[CCS21] Cristina Caraci, Serena Cenatiempo, and Benjamin Schlein. Bose–Einstein condensation for two dimensional bosons in the Gross–Pitaevskii regime. Journal of Statistical Physics, 183(3): ??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/article/
10.1007/s10955-021-02766-6;   http://link.springer.com/
content/pdf/10.1007/s10955-021-02766-6.pdf.

Conlon:2020:GSC

[CD20a] Joseph G. Conlon and Michael Dabkowski. Global stability for a class of nonlinear PDE with non-local term. Journal of Statistical Physics, 178(2):420–471, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Conlon:2020:GAS

[CD20b] Joseph G. Conlon and Michael Dabkowski. On global asymptotic stability for the LSW model with subcritical initial data. Journal of Statistical Physics, 178(1):117–177, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chetlur:2020:SPD

[CDD20] Vishnu Vardhan Chetlur, Harpreet S. Dhillon, and Carl P. Dettmann. Shortest path distance in Manhattan Poisson line Cox process. Journal of Statistical Physics, 181(6):2109–2130, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cipriani:2021:SLM

[CDH21] Alessandra Cipriani, Biltu Dan, and Rajat Subhra Hazra. The scaling limit of the (∇+Δ)-model. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Collet:2023:GRT

[CDHM23] Pierre Collet, François Dunlop, Thierry Huillet, and Arif Mardin. A Gibbsian random tree with nearest neighbour interaction. Journal of Statistical Physics, 190(4):??, April 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03087-6.

Cipriani:2023:MMM

[CDHR23] Alessandra Cipriani, Biltu Dan, Rajat Subhra Hazra, and Rounak Ray. Maximum of the membrane model on regular trees. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03043-w.

Chakraborti:2023:BED

[CDK23] Subhadip Chakraborti, Abhishek Dhar, and Anupam Kundu. Boltzmann’s entropy during free expansion of an interacting gas. Journal of Statistical Physics, 190(4):??, April 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03088-5.

Chariker:2023:SLG

[CDLP23] Logan Chariker, Anna De Masi, Joel L. Lebowitz, and Errico Presutti. Scaling limit of a generalized contact process. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-03050-x.

Can:2021:LTO

[CDN21] Van Hao Can, Thai Son Doan, and Van Quyet Nguyen. Limit theorems for the one-dimensional random walk with random resetting to the maximum. Journal of Statistical Physics, 183 (2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-021-02754-w.

Cavagna:2021:DRG

[CDS21] Andrea Cavagna, Luca Di Carlo, and Mattia Scandolo. Dynamical renormalization group for mode-coupling field theories with solenoidal constraint. Journal of Statistical Physics, 184 (3):??, September 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02800-7.

Charles:2021:MVA

[CDW21] Frédérique Charles, Bruno Després, and Ricardo Weder. The magnetized Vlasov–Ampère system and the Bernstein–Landau paradox. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02755-9.

Caracciolo:2021:NOM

[CES21] Sergio Caracciolo, Vittorio Erba, and Andrea Sportiello. The number of optimal matchings for Euclidean assignment on the line. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02741-1.

Caby:2020:CEI

[CFVY20] Th. Caby, D. Faranda, S. Vaienti, and P. Yiou. On the computation of the extremal index for time series. Journal of Statistical Physics, 179(5–6):1666–1697, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Campanino:2020:UBE

[CG20] M. Campanino and M. Gianfelice. Uniform bound of the entanglement for the ground state of the quantum Ising model with large transverse magnetic field. Journal of Statistical Physics, 178(5):1248–1267, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chatterjee:2022:GPS

[CG22] Avik P. Chatterjee and Claudio Grimaldi. Geometric percolation of spherically symmetric fractal aggregates. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02952-0.

Campa:2023:RDL

[CG23] Alessandro Campa and Shamik Gupta. Relaxation dynamics in a long-range system with mixed Hamiltonian and non-Hamiltonian interactions. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03033-y.

Carbone:2022:GCL

[CGH22] Raffaella Carbone, Federico Girotti, and Anderson Melchor Hernandez. On a generalized central limit theorem and large deviations for homogeneous open quantum walks. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02938-y.

Carollo:2021:LDL

[CGJ21] Federico Carollo, Juan P. Garrahan, and Robert L. Jack. Large deviations at level 2.5 for Markovian open quantum systems: Quantum jumps and quantum state diffusion. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02799-x.

Chan:2021:RWL

[CGL21] Swee Hong Chan, Lila Greco, and Peter Li. Random walks with local memory. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02791-5.

Caglioti:2020:QOT

[CGP20a] E. Caglioti, F. Golse, and T. Paul. Quantum optimal transport is cheaper. Journal of Statistical Physics, 181(1):149–162, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Carinci:2020:SSI

[CGP20b] Gioia Carinci, Cristian Giardinà, and Errico Presutti. Stationary states in infinite volume with non-zero current. Journal of Statistical Physics, 180(1–6):366–397, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-019-02427-9.pdf.

Campa:2021:CME

[CGT21] Alessandro Campa, Giacomo Gori, and Andrea Trombettoni. Computation of microcanonical entropy at fixed magnetization without direct counting. Journal of Statistical Physics, 184(2): ??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-021-02809-y.

Casini:2023:URD

[CGV23] Francesco Casini, Cristian Giardinà, and Cecilia Vernia. Uphill in reaction–diffusion multi-species interacting particles systems. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-023-
03141-3.

Coghi:2020:LDP

[CH20] Francesco Coghi and Rosemary J. Harris. A large deviation perspective on ratio observables in reset processes: Robustness of rate functions. Journal of Statistical Physics, 179(1):131–154, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-020-02513-3.pdf.

Cohen:2022:TLS

[CH22] Joel E. Cohen and Thierry E. Huillet. Taylor’s law for some infinitely divisible probability distributions from population models. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02962-y.

Chatterjee:2023:PTS

[Cha23] Avik P. Chatterjee. Percolation thresholds for spherically symmetric fractal aggregates. Journal of Statistical Physics, 190 (6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-023-03124-4.

Conache:2022:VVL

[CHR22] Diana Conache, Markus Heydenreich, and Silke W. W. Rolles. Variance of voltages in a lattice Coulomb gas. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02857-4.

Chulaevsky:2023:ALD

[Chu23] Victor Chulaevsky. Anderson localization in discrete random displacements models. Journal of Statistical Physics, 190(1): ??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-03020-3.

Ciucu:2022:EMG

[Ciu22] Mihai Ciucu. The effect of microscopic gap displacement on the correlation of gaps in dimer systems. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02828-9.

Camia:2020:GPR

[CJN20] Federico Camia, Jianping Jiang, and Charles M. Newman. A Gaussian process related to the mass spectrum of the nearcritical Ising model. Journal of Statistical Physics, 179(4):885– 900, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Camia:2022:IMC

[CJN22] Federico Camia, Jianping Jiang, and Charles M. Newman. Ising model with Curie–Weiss perturbation. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02935-1.

Crawford:2020:TIC

[CK20] Nicholas Crawford and Gady Kozma. The Toom interface via coupling. Journal of Statistical Physics, 179(2):408–447, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Caetano:2022:CSI

[CK22] João Caetano and Shota Komatsu. Crosscap states in integrable field theories and spin chains. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02914-6.

Chen:2020:LWP

[CKL20] Hongxu Chen, Chanwoo Kim, and Qin Li. Local well-posedness of Vlasov–Poisson–Boltzmann equation with generalized diffuse boundary condition. Journal of Statistical Physics, 179(2):535– 631, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Coquille:2023:EIG

[CKL23] Loren Coquille, Christof Külske, and Arnaud Le Ny. Extremal inhomogeneous Gibbs states for SOS-models and finite-spin models on trees. Journal of Statistical Physics, 190(4):??, April 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03081-y.

Chen:2021:CMF

[CLL21] Li Chen, Jinyeop Lee, and Matthew Liew. Combined meanfield and semiclassical limits of large fermionic systems. Journal of Statistical Physics, 182(2):??, February 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-021-02700-w.pdf.

Chen:2023:MNE

[CLLL23] Li Chen, Jinyeop Lee, Yue Li, and Matthew Liew. A mixednorm estimate of the two-particle reduced density matrix of many-body Schrödinger dynamics for deriving the Vlasov equation. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03123-5.

Chekroun:2020:VAC

[CLM20] Mickaël D. Chekroun, Honghu Liu, and James C. McWilliams. Variational approach to closure of nonlinear dynamical systems: Autonomous case. Journal of Statistical Physics, 179(5–6):1073– 1160, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Carter:2015:EHO

[CLTC15] Jason A. Carter, Christine H. Lind, M. Phuong Truong, and Eva-Maria S. Collins. To each his own: Reproductive strategies and success of three common planarian species: Schmidtea mediterranea, Dugesia japonica, and Dugesia tigrina. Journal of Statistical Physics, 161(1):250–272, October 2015. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

015-1310-1. See corrections [CLTC23].

Carter:2023:CEH

[CLTC23] Jason A. Carter, Christine H. Lind, M. Phuong Truong, and Eva-Maria S. Collins. Correction: To each his own. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-03009-y. See

[CLTC15].

Chabane:2022:EHL

[CLV22] Lydia Chabane, Alexandre Lazarescu, and Gatien Verley. Effective Hamiltonians and Lagrangians for conditioned Markov processes at large volume. Journal of Statistical Physics, 187 (1):??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-02898-3.

Cheng:2023:PES

[CLW23] Dandan Cheng, Zhiming Li, and Weisheng Wu. Preimage entropy and stable entropy on subsets. Journal of Statistical Physics, 190(4):??, April 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03084-9.

Chen:2021:IMI

[CLY21] Xiaopeng Chen, Chang-Bing Li, and Yuan-Ling Ye. Invariant measure for infinite weakly hyperbolic iterated function systems.

Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Carlen:2020:CNC

[CM20a] Eric A. Carlen and Jan Maas. Correction to: Non-commutative Calculus, Optimal Transport and Functional Inequalities in Dissipative Quantum Systems. Journal of Statistical Physics, 181(6):2432–2433, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-020-02671-4.

pdf. See [CM20b].

Carlen:2020:NCC

[CM20b] Eric A. Carlen and Jan Maas. Non-commutative calculus, optimal transport and functional inequalities in dissipative quantum systems. Journal of Statistical Physics, 178(2):319–378, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02434-w.pdf. See correction [CM20a].

Coletti:2020:BPD

[CMG20] Cristian F. Coletti, Daniel Miranda, and Sebastian P. Grynberg. Boolean percolation on doubling graphs. Journal of Statistical Physics, 178(3):814–831, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Chazottes:2020:GCU

[CMRU20] J.-R. Chazottes, J. Moles, F. Redig, and E. Ugalde. Gaussian concentration and uniqueness of equilibrium states in lattice systems. Journal of Statistical Physics, 181(6):2131–2149, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Camiola:2021:HMC

[CN21] Vito Dario Camiola and Giovanni Nastasi. Hydrodynamical model for charge transport in graphene nanoribbons. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02802-5.

Chiarini:2023:PTL

[CN23] Alberto Chiarini and Maximilian Nitzschner. Phase transition for level-set percolation of the membrane model in dimensions d ≥ 5. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03072-z.

Coppini:2022:NFP

[Cop22] Fabio Coppini. A note on Fokker–Planck equations and graphons. Journal of Statistical Physics, 187(2):??, May 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02905-7.

Corwin:2020:LSS

[CP20] Ivan Corwin and Shalin Parekh. Limit shape of subpartitionmaximizing partitions. Journal of Statistical Physics, 180(1–6): 597–611, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cheng:2021:SLC

[CQ21] Yu-Chen Cheng and Hong Qian. Stochastic limit-cycle oscillations of a nonlinear system under random perturbations. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Crawford:2021:SHS

[Cra21] Nicholas Crawford. Supersymmetric hyperbolic σ-models and bounds on correlations in two dimensions. Journal of Statistical Physics, 184(3):??, September 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02817-y.

Criens:2023:PCW

[Cri23] David Criens. Propagation of chaos for weakly interacting mild solutions to stochastic partial differential equations. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-023-03121-7.

Chenn:2020:DPB

[CS20] Ilias Chenn and I. M. Sigal. On derivation of the Poisson– Boltzmann equation. Journal of Statistical Physics, 180(1–6): 954–1001, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Croydon:2022:SSR

[CS22] David A. Croydon and Makiko Sasada. On the stationary solutions of random polymer models and their zero-temperature limits. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02947-x.

Cong:2021:LTA

[CSZ21] Hongzi Cong, Yunfeng Shi, and Zhifei Zhang. Long-time Anderson localization for the nonlinear Schrödinger equation revisited. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Cortez:2021:UPC

[CT21] Roberto Cortez and Hagop Tossounian. Uniform propagation of chaos for the thermostated Kac model. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02763-9.

Chen:2022:NPE

[CT22] Xi Chen and Ilya Timofeyev. Non-parametric estimation of stochastic differential equations from stationary time-series. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-021-
02847-6.

Chekroun:2020:RPR

[CTDN20] Mickaël D. Chekroun, Alexis Tantet, Henk A. Dijkstra, and J. David Neelin. Ruelle–Pollicott resonances of stochastic systems in reduced state space. Part I: Theory. Journal of Statistical Physics, 179(5–6):1366–1402, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Claeys:2022:URM

[CW22] Tom Claeys and Dong Wang. Universality for random matrices with equi-spaced external source: a case study of a biorthogonal ensemble. Journal of Statistical Physics, 188(2):??, August 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02937-z.

Cerrai:2020:AAS

[CWZ20] Sandra Cerrai, Jan Wehr, and Yichun Zhu. An averaging approach to the Smoluchowski–Kramers approximation in the presence of a varying magnetic field. Journal of Statistical Physics, 181(1):132–148, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Czudek:2022:SCC

[Czu22] Klaudiusz Czudek. Some counterexamples to the central limit theorem for random rotations. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02975-7.

Darzi:2020:LEI

[DAN20] Rahmat Darzi, Bahram Agheli, and Juan J. Nieto. Langevin equation involving three fractional orders. Journal of Statistical Physics, 178(4):986–995, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Das:2022:DCD

[DAR22] Debankur Das, Pappu Acharya, and Kabir Ramola. Displacement correlations in disordered athermal networks. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-022-
02981-9.

DAutilia:2021:PST

[DAT21] Roberto D’Autilia, Louis Nantenaina Andrianaivo, and Alessio Troiani. Parallel simulation of two-dimensional Ising models using probabilistic cellular automata. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02792-4.

Das:2020:NFH

[DDD^{+}20] Avijit Das, Kedar Damle, Abhishek Dhar, David A. Huse, Manas Kulkarni, Christian B. Mendl, and Herbert Spohn. Nonlinear fluctuating hydrodynamics for the classical XXZ spin chain. Journal of Statistical Physics, 180(1–6):238–262, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Debin:2020:ACT

[DDG20] Bryan Debin, Philippe Di Francesco, and Emmanuel Guitter. Arctic curves of the twenty-vertex model with domain wall boundaries. Journal of Statistical Physics, 179(1):33–89, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dietert:2022:FJR

[DE22] Helge Dietert and Josephine Evans. Finding the jump rate for fastest decay in the Goldstein–Taylor model. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02925-3.

Dematte:2023:KED

[Dem23] Elena Demattè. On a kinetic equation describing the behavior of a gas interacting mainly with radiation. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03128-0.

Deng:2022:GEN

[Den22] Dingqun Deng. Global existence of non-cutoff Boltzmann equation in weighted Sobolev space. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02956-w.

Derrida:2023:COB

[Der23] Bernard Derrida. Cross-overs of Bramson’s shift at the transition between pulled and pushed fronts. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03077-8.

DaiPra:2020:OBM

[DFP20] Paolo Dai Pra, Marco Formentin, and Guglielmo Pelino. Oscillatory behavior in a model of non–Markovian mean field interacting spins. Journal of Statistical Physics, 179(3):690–712, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02544-w.pdf.

Duplantier:2020:SMC

[DG20] Bertrand Duplantier and Anthony J. Guttmann. Statistical mechanics of confined polymer networks. Journal of Statistical Physics, 180(1–6):1061–1094, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dobson:2023:CNL

[DG23] Matthew Dobson and Abdel Kader A. Geraldo. Convergence of nonequilibrium Langevin dynamics for planar flows. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-023-03109-3.

Durr:2020:BTH

[DGT^{+}20] Detlef Dürr, Sheldon Goldstein, Stefan Teufel, Roderich Tumulka, and Nino Zanghì. Bohmian trajectories for Hamiltonians with interior-boundary conditions. Journal of Statistical Physics, 180(1–6):34–73, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Durr:2020:QMS

[DGZ20] Detlef Dürr, Sheldon Goldstein, and Nino Zanghí. Quantum motion on shape space and the gauge dependent emergence of dynamics and probability in absolute space and time. Journal of Statistical Physics, 180(1–6):92–134, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Drivas:2020:LAS

[DHL20] Theodore D. Drivas, Darryl D. Holm, and James-Michael Leahy. Lagrangian averaged stochastic advection by Lie transport for fluids. Journal of Statistical Physics, 179(5–6):1304–1342, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02493-4.pdf.

DeRoeck:2020:SOD

[DHO20] Wojciech De Roeck, François Huveneers, and Stefano Olla. Subdiffusion in one-dimensional Hamiltonian chains with sparse interactions. Journal of Statistical Physics, 180(1–6):678–698, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Damak:2020:DFT

[DHP20] Mondher Damak, Mayssa Hammami, and Claude-Alain Pillet. A detailed fluctuation theorem for heat fluxes in harmonic networks out of thermal equilibrium. Journal of Statistical Physics, 180(1– 6):263–296, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Derrida:2021:LDS

[DHS21] Bernard Derrida, Ori Hirschberg, and Tridib Sadhu. Large deviations in the symmetric simple exclusion process with slow boundaries. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dragicevic:2023:VVA

[DHS23] D. Dragicević, Y. Hafouta, and J. Sedro. A vector-valued almost sure invariance principle for random expanding on average cocycles. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03067-w.

denHollander:2022:PAM

[dHW22] F. den Hollander and D. Wang. The parabolic Anderson model on a Galton–Watson tree revisited. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02951-1.

Crescenzo:2023:SFV

[DIM23] Antonio Di Crescenzo, Antonella Iuliano, and Verdiana Mustaro. On some finite-velocity random motions driven by the geometric counting process. Journal of Statistical Physics, 190(3): ??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-03045-8.

Dolai:2023:SID

[DK23a] Dhriti Ranjan Dolai and M. Krishna. Smoothness of integrated density of states of the Anderson model on Bethe lattice in high disorder. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03114-6.

Dymov:2023:FES

[DK23b] Andrey Dymov and Sergei Kuksin. Formal expansions in stochastic model for wave turbulence 2: Method of diagram decomposition. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02998-0.

DelMagno:2021:HPB

[DLDG21] Gianluigi Del Magno, João Lopes Dias, Pedro Duarte, and José Pedro Gaivão. Hyperbolic polygonal billiards close to 1dimensional piecewise expanding maps. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dunlop:2022:GWT

[DM22] François Dunlop and Arif Mardin. Galton–Watson trees with first ancestor interaction. Journal of Statistical Physics, 189 (3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03000-7.

DeMasi:2020:IFN

[DMO20] Anna De Masi, Immacolata Merola, and Stefano Olla. Interface fluctuations in non equilibrium stationary states: The SOS approximation. Journal of Statistical Physics, 180(1–6):414–426, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

DeBruyne:2023:UOS

[DMS23] Benjamin De Bruyne, Satya N. Majumdar, and Grégory Schehr. Universal order statistics for random walks & Lévy flights. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03027-w.

Doyon:2022:DSH

[Doy22] Benjamin Doyon. Diffusion and superdiffusion from hydrodynamic projections. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-021-02863-6.

Derbyshev:2021:NGT

[DP21a] A. E. Derbyshev and A. M. Povolotsky. Nonstationary generalized TASEP in KPZ and jamming regimes. Journal of Statistical Physics, 185(3):??, December 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02840-z.

Duboscq:2021:EMM

[DP21b] Romain Duboscq and Olivier Pinaud. Entropy minimization for many-body quantum systems. Journal of Statistical Physics, 185(1):??, October 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02824-z.

DiTerlizzi:2020:ESG

[DRB20] Ivan Di Terlizzi, Felix Ritort, and Marco Baiesi. Explicit solution of the generalised Langevin equation. Journal of Statistical Physics, 181(5):1609–1635, December 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02639-4.pdf.

Dembo:2020:DSS

[DS20] Amir Dembo and Eliran Subag. Dynamics for spherical spin glasses: Disorder dependent initial conditions. Journal of Statistical Physics, 181(2):465–514, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dey:2021:ERE

[DS21] Kajal Krishna Dey and Golam Ali Sekh. Effects of random excitations on the dynamical response of Duffing systems. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Laxa:2023:MSS

[dSL23] Kádmo de S. Laxa. Metastability in a stochastic system of spiking neurons with leakage. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03035-w.

Dodson:2020:NSE

[DSS20] Benjamin Dodson, Avraham Soffer, and Thomas Spencer. The nonlinear Schrödinger equation on Z and R with bounded initial data: Examples and conjectures. Journal of Statistical Physics, 180(1–6):910–934, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dimarco:2020:ART

[DT20] Giacomo Dimarco and Andrea Tosin. The Aw–Rascle traffic model: Enskog-type kinetic derivation and generalisations. Journal of Statistical Physics, 178(1):178–210, January 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dubovski:2021:MPV

[DT21] Pavel B. Dubovski and Michael Tamarov. Mathematics of parking: Varying parking rate. Journal of Statistical Physics, 182(2): ??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dimarco:2022:KDA

[DTZ22] Giacomo Dimarco, Andrea Tosin, and Mattia Zanella. Kinetic derivation of Aw–Rascle–Zhang-type traffic models with driverassist vehicles. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02862-7.

Duerinckx:2022:EFR

[Due22] Mitia Duerinckx. Eigenvalue fluctuations for random elliptic operators in homogenization regime. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02918-2.

Dey:2021:FRM

[DW21] Partha S. Dey and Qiang Wu. Fluctuation results for multispecies Sherrington–Kirkpatrick model in the replica symmetric regime. Journal of Statistical Physics, 185(3):??, December 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02835-w.

Duan:2021:VPL

[DY21] Renjun Duan and Hongjun Yu. The 3D Vlasov–Poisson–Landau system near 1D local Maxwellians. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

deZarate:2020:NEF

[dZGFBC20] José M. Ortiz de Zárate, Loreto García-Fernández, Henri Bataller, and Fabrizio Croccolo. Non-equilibrium fluctuations in a ternary mixture subjected to a temperature gradient. Journal of Statistical Physics, 181(1):1–18, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Dunlop:2020:ECS

[DZW^{+}20] J. W. C. Dunlop, G. A. Zickler, R. Weinkamer, F. D. Fischer, and P. Fratzl. The emergence of complexity from a simple model for tissue growth. Journal of Statistical Physics, 180(1–6): 459–473, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-019-02461-7.pdf.

Erhard:2023:DER

[EFR23] Dirk Erhard, Tertuliano Franco, and Guilherme Reis. The directed edge reinforced random walk: The ant mill phenomenon. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03031-0.

Engel:2021:HCF

[EGK21] Maximilian Engel, Marios Antonios Gkogkas, and Christian Kuehn. Homogenization of coupled fast–slow systems via intermediate stochastic regularization. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02765-7.

Erignoux:2020:HSN

[EGN20] C. Erignoux, P. Gonçalves, and G. Nahum. Hydrodynamics for SSEP with non-reversible slow boundary dynamics: Part I, the critical regime and beyond. Journal of Statistical Physics, 181 (4):1433–1469, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Eckhoff:2020:LPF

[EGvdHN20] Maren Eckhoff, Jesse Goodman, Remco van der Hofstad, and Francesca R. Nardi. Long paths in first passage percolation on the complete graph II. Global branching dynamics. Journal of Statistical Physics, 181(2):364–447, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02585-1.pdf.

Elskens:2020:MFK

[EK20a] Y. Elskens and M. K.-H. Kiessling. Microscopic foundations of kinetic plasma theory: The relativistic Vlasov–Maxwell equations and their radiation–reaction-corrected generalization. Journal of Statistical Physics, 180(1–6):749–772, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Endo:2020:TPM

[EK20b] Taiki Endo and Makoto Katori. Three-parametric Marcenko– Pastur density. Journal of Statistical Physics, 178(6):1397–1416, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Eichinger:2020:SPE

[EKN20] Katharina Eichinger, Christian Kuehn, and Alexandra Neamtu. Sample paths estimates for stochastic fast–slow systems driven by fractional Brownian motion. Journal of Statistical Physics, 179(5–6):1222–1266, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-020-02485-4.
pdf.

Eldan:2020:SAC

[Eld20] Ronen Eldan. A simple approach to chaos for p-spin models. Journal of Statistical Physics, 181(4):1266–1276, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Esposito:2020:SNE

[EM20] R. Esposito and R. Marra. Stationary non equilibrium states in kinetic theory. Journal of Statistical Physics, 180(1–6):773–809, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Endo:2022:LCL

[EM22] Eric O. Endo and Vlad Margarint. Local central limit theorem for long-range two-body potentials at sufficiently high temperatures. Journal of Statistical Physics, 189(3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02994-4.

Elyasizad:2021:LVC

[EN21] L. Elyasizad and M. D. Niry. Local violation of conservation in the Abelian sandpile model through fractal patterns of nonconservative sites. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02779-1.

Alaoui:2022:APS

[ES22] Ahmed El Alaoui and Mark Sellke. Algorithmic pure states for the negative spherical perceptron. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02976-6.

Fachechi:2021:CPS

[Fac21a] Alberto Fachechi. Correction to: PDE/statistical mechanics duality: Relation between Guerra’s interpolated p-spin ferromagnets and the Burgers hierarchy. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-021-02770-w;
 http://link.springer.com/content/pdf/10.1007/s10955-
021-02770-w.pdf.

Fachechi:2021:PSM

[Fac21b] Alberto Fachechi. PDE/statistical mechanics duality: Relation between Guerra’s interpolated p-spin ferromagnets and the Burgers hierarchy. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02747-9.

Feliachi:2021:DLD

[FB21] Ouassim Feliachi and Freddy Bouchet. Dynamical large deviations for plasmas below the Debye length and the Landau equation. Journal of Statistical Physics, 183(3):??, June 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
021-02771-9.

Feliachi:2022:DLD

[FB22] Ouassim Feliachi and Freddy Bouchet. Dynamical large deviations for homogeneous systems with long range interactions and the Balescu–Guernsey–Lenard equation. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02854-7.

Fendley:2021:IBT

[Fen21] Paul Fendley. Integrability and braided tensor categories. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fanelli:2020:SSB

[FF20] Francesco Fanelli and Eduard Feireisl. Statistical solutions to the barotropic Navier–Stokes system. Journal of Statistical Physics, 181(1):212–245, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Freitas:2020:PPN

[FFMV20] Ana Cristina Moreira Freitas, Jorge Milhazes Freitas, Mário Magalhães, and Sandro Vaienti. Point processes of non stationary sequences generated by sequential and random dynamical systems. Journal of Statistical Physics, 181(4):1365–1409, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Frassek:2020:NCQ

[FGK20] Rouven Frassek, Cristian Giardinà, and Jorge Kurchan. Noncompact quantum spin chains as integrable stochastic particle processes. Journal of Statistical Physics, 180(1–6):135–171, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fiorelli:2023:SEP

[FGM23] Eliana Fiorelli, Stefano Gherardini, and Stefano Marcantoni. Stochastic entropy production: Fluctuation relation and irreversibility mitigation in non-unital quantum dynamics. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-023-03118-2.

Ferrari:2020:TGP

[FGR20] Analía Ferrari, Pablo Groisman, and Krishnamurthi Ravishankar. Tumor growth, R-positivity, multitype branching and quasistationarity. Journal of Statistical Physics, 180(1–6):427– 439, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Floreani:2022:SIP

[FGR22] Simone Floreani, Cristian Giardinà, and Frank Redig. Switching interacting particle systems: Scaling limits, uphill diffusion and boundary layer. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02878-7.

Feireisl:2022:RCF

[FH22] Eduard Feireisl and Martina Hofmanová. Randomness in compressible fluid flows past an obstacle. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02879-6.

Fialho:2020:APG

[Fia20] Paula Mendes Soares Fialho. Abstract polymer gas: A simple inductive proof of the Fernández–Procacci criterion. Journal of Statistical Physics, 178(6):1354–1361, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Figueiredo:2020:ETS

[FIS20] Daniel Figueiredo, Giulio Iacobelli, and Seva Shneer. The end time of SIS epidemics driven by random walks on edge-transitive graphs. Journal of Statistical Physics, 179(3):651–671, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02547-7.pdf.

Freidlin:2020:DPC

[FK20] M. Freidlin and L. Koralov. Diffusion in the presence of cells with semi-permeable membranes. Journal of Statistical Physics, 178 (6):1417–1441, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fantoni:2021:MCE

[FK21] Riccardo Fantoni and John R. Klauder. Monte Carlo evaluation of the continuum limit of the two-point function of the Euclidean free real scalar field subject to affine quantization. Journal of Statistical Physics, 184(3):??, September 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02818-x.

Frohlich:2020:PIA

[FKSS20] Jürg Fröhlich, Antti Knowles, Benjamin Schlein, and Vedran Sohinger. A path-integral analysis of interacting Bose gases and loop gases. Journal of Statistical Physics, 180(1–6):810–831, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Foxall:2018:CTF

[FL18] Eric Foxall and Hanbaek Lyu. Clustering in the three and four color cyclic particle systems in one dimension. Journal of Statistical Physics, 171(3):470–483, May 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [FL20a].

Foxall:2020:CCT

[FL20a] Eric Foxall and Hanbaek Lyu. Correction to: Clustering in the Three and Four Color Cyclic Particle Systems in One Dimension. Journal of Statistical Physics, 181(5):2011–2014, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02651-8.pdf. See [FL18].

Fyodorov:2020:MPH

[FL20b] Yan V. Fyodorov and Pierre Le Doussal. Manifolds pinned by a high-dimensional random landscape: Hessian at the global energy minimum. Journal of Statistical Physics, 179(1):176–215, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-020-02522-2.pdf.

Fan:2023:VML

[FL23] Yingzhe Fan and Yuanjie Lei. The Vlasov–Maxwell–Landau system with Coulomb potential and strong background magnetic field. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03030-1.

Foster:2022:ARC

[FLM22] James Foster, Terry Lyons, and Vlad Margarint. An asymptotic radius of convergence for the Loewner equation and simulation of SLE_{κ} traces via splitting. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02979-3.

Ferreira:2023:NES

[FLNV23] Marina A. Ferreira, Jani Lukkarinen, Alessia Nota, and Juan J. L. Velázquez. Non-equilibrium stationary solutions for multicomponent coagulation systems with injection. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03107-5.

Fronk:2021:LPF

[FM21] Jacob Fronk and Andreas Mielke. Localised pair formation in bosonic flat-band Hubbard models. Journal of Statistical Physics, 185(2):??, November 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02842-x.

Ferrari:2020:STT

[FN20] Patrik L. Ferrari and Peter Nejjar. Statistics of TASEP with three merging characteristics. Journal of Statistical Physics, 180 (1–6):398–413, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Funaki:2021:SEV

[FNS21] Tadahisa Funaki, Yuto Nishijima, and Hayate Suda. Stochastic eight-vertex model, its invariant measures and KPZ limit. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-021-
02789-z.

Fontes:2021:ILG

[FP21] Luiz Renato Fontes and Gabriel R. C. Peixoto. Infinite level GREM-like K-processes existence and convergence. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fu:2021:FPS

[FPT21] Jingyi Fu, Benoit Perthame, and Min Tang. Fokker–Plank system for movement of micro-organism population in confined environment. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02760-y.

Flandoli:2022:IDL

[FPV22] Franco Flandoli, Umberto Pappalettera, and Milo Viviani. On the infinite dimension limit of invariant measures and solutions of Zeitlin’s 2D Euler equations. Journal of Statistical Physics, 189(3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03007-0.

Ferrari:2020:SST

[FR20] Pablo A. Ferrari and Leonardo T. Rolla. Slow-to-start traffic model: Traffic saturation and scaling limits. Journal of Statistical Physics, 180(1–6):935–953, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Freidlin:2022:PSF

[Fre22] Mark I. Freidlin. Perturbation of systems with a first integral: Motion on the Reeb graph. Journal of Statistical Physics, 189 (3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02995-3.

Favre:2020:HFR

[FS20a] Gianluca Favre and Christian Schmeiser. Hypocoercivity and fast reaction limit for linear reaction networks with kinetic transport. Journal of Statistical Physics, 178(6):1319–1335, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02503-5.pdf.

Friedrich:2020:COD

[FS20b] Manuel Friedrich and Ulisse Stefanelli. Crystallization in a onedimensional periodic landscape. Journal of Statistical Physics, 179(2):485–501, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Furtlehner:2023:FDF

[Fur23] Cyril Furtlehner. Free dynamics of feature learning processes. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-022-
03064-5.

Fitzner:2021:NLT

[FvdH21] Robert Fitzner and Remco van der Hofstad. NoBLE for lattice trees and lattice animals. Journal of Statistical Physics, 185(2):??, November 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02816-z.

Fitzner:2022:CNL

[FvdH22] Robert Fitzner and Remco van der Hofstad. Correction to: No-BLE for lattice trees and lattice animals. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02853-8.

Funaki:2023:MMC

[FvMST23] Tadahisa Funaki, Patrick van Meurs, Sunder Sethuraman, and Kenkichi Tsunoda. Motion by mean curvature from Glauber– Kawasaki dynamics with speed change. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-03044-9.

Freidlin:2021:DAN

[FW21] M. I. Freidlin and A. D. Wentzell. Diffusion approximation for noise-induced evolution of first integrals in multifrequency systems. Journal of Statistical Physics, 182(3):??, March 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Fatkullin:2021:LSG

[FX21] Ibrahim Fatkullin and Jianfei Xue. Limit shapes for Gibbs partitions of sets. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02756-8.

Forrester:2020:LES

[FZ20] P. J. Forrester and Jiyuan Zhang. Lyapunov exponents for some isotropic random matrix ensembles. Journal of Statistical Physics, 180(1–6):558–575, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ganapa:2020:TLO

[GAD20] Santhosh Ganapa, Amit Apte, and Abhishek Dhar. Thermalization of local observables in the α-FPUT chain. Journal of Statistical Physics, 180(1–6):1010–1030, September 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gallavotti:2020:NFR

[Gal20] Giovanni Gallavotti. Nonequilibrium and fluctuation relation. Journal of Statistical Physics, 180(1–6):172–226, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gallavotti:2021:VRC

[Gal21] Giovanni Gallavotti. Viscosity, reversibillity, chaotic hypothesis, fluctuation theorem and Lyapunov pairing. Journal of Statistical Physics, 185(3):??, December 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02830-1.

Gao:2020:APM

[Gao20] Peng Gao. Averaging principle for multiscale stochastic fractional Schrödinger–Korteweg–de Vries system. Journal of Statistical Physics, 181(5):1781–1816, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gao:2022:APS

[Gao22] Peng Gao. Averaging principles for stochastic 2D Navier–Stokes equations. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02876-9.

Gates:2023:ITG

[Gat23] David J. Gates. Initial transient growth in polynuclear models for crystal layers and epitaxial films. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03120-8.

Guioth:2022:PLD

[GBE22] Jules Guioth, Freddy Bouchet, and Gregory L. Eyink. Path large deviations for the kinetic theory of weak turbulence. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-022-
02971-x.

Gottwald:2020:DRT

[GG20] Georg A. Gottwald and Federica Gugole. Detecting regime transitions in time series using dynamic mode decomposition. Journal of Statistical Physics, 179(5–6):1028–1045, June 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gerasimenko:2022:PCH

[GG22] Viktor Gerasimenko and Igor Gapyak. Propagation of correlations in a hard-sphere system. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02958-8.

Grava:2023:DIS

[GGGM23] Tamara Grava, Massimo Gisonni, Giorgio Gubbiotti, and Guido Mazzuca. Discrete integrable systems and random Lax matrices. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03024-z.

Giardina:2021:ACG

[GGM21] Cristian Giardinà, Claudio Giberti, and Elena Magnanini. Approximating the cumulant generating function of triangles in the Erdös–Rényi random graph. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Giacomin:2020:LBJ

[GH20a] Giambattista Giacomin and Benjamin Havret. Localization, big-jump regime and the effect of disorder for a class of generalized pinning models. Journal of Statistical Physics, 181(6): 2015–2049, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Goncalves:2020:SFT

[GH20b] Branda Goncalves and Thierry Huillet. Scaling features of two special Markov chains involving total disasters. Journal of Statistical Physics, 178(2):499–531, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Geurts:2020:LET

[GHL20] Bernard J. Geurts, Darryl D. Holm, and Erwin Luesink. Lyapunov exponents of two stochastic Lorenz 63 systems. Journal of Statistical Physics, 179(5–6):1343–1365, June 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-019-02457-3.pdf.

Goddard:2021:MIG

[GHO21] B. D. Goddard, T. D. Hurst, and R. Ocone. Modelling inelastic granular media using dynamical density functional theory. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-020-
02675-0.

Gronbech-Jensen:2023:ANG

[GJ23] Niels Grønbech-Jensen. On the application of non-Gaussian noise in stochastic Langevin simulations. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03104-8.

Ge:2021:MSG

[GJJ21] Hao Ge, Chen Jia, and Xiao Jin. Martingale structure for general thermodynamic functionals of diffusion processes under secondorder averaging. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02798-y.

Gantert:2022:TLE

[GK22a] Nina Gantert and Achim Klenke. The tail of the length of an excursion in a trap of random size. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02957-9.

Gu:2022:HTB

[GK22b] Yu Gu and Tomasz Komorowski. High temperature behaviors of the directed polymer on a cylinder. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02899-2.

Grava:2021:CFC

[GKM21] T. Grava, T. Kriecherbauer, and K. D. T.-R. McLaughlin. Correlation functions for a chain of short range oscillators. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02735-z.

Gutierrez:2020:RSU

[GL20] Manuel Santos Gutiérrez and Valerio Lucarini. Response and sensitivity using Markov chains. Journal of Statistical Physics, 179(5–6):1572–1593, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-020-02504-4.
pdf.

Gao:2022:RMD

[GL22a] Yuan Gao and Jian-Guo Liu. Revisit of macroscopic dynamics for some non-equilibrium chemical reactions from a Hamiltonian viewpoint. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02985-5.

Godreche:2022:RSI

[GL22b] Claude Godrèche and Jean-Marc Luck. Record statistics of integrated random walks and the random acceleration process. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02852-9.

Gufler:2022:BTT

[GL22c] Stephan Gufler and Oren Louidor. Ballot theorems for the twodimensional discrete Gaussian free field. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02970-y.

Galli:2023:EEK

[GL23a] Daniele Galli and Marco Lenci. Extensions of exact and Kmixing dynamical systems. Journal of Statistical Physics, 190 (1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03039-6.

Gheissari:2023:ADM

[GL23b] Reza Gheissari and Eyal Lubetzky. Approximate domain Markov property for rigid Ising interfaces. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-023-03101-x.

Galves:2020:SIN

[GLPP20] A. Galves, E. Löcherbach, C. Pouzat, and E. Presutti. A system of interacting neurons with short term synaptic facilitation. Journal of Statistical Physics, 178(4):869–892, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gripon:2021:SRR

[GLV21] Vicent Gripon, Matthias Löwe, and Franck Vermet. Some remarks on replicated simulated annealing. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-021-02727-z.
pdf.

Grauer:2021:PAL

[GLY21] Arne Grauer, Lukas Lüchtrath, and Mark Yarrow. Preferential attachment with location-based choice: Degree distribution in the noncondensation phase. Journal of Statistical Physics, 184 (1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-021-02782-6.

Griffin:2021:QTC

[GM21a] Jory Griffin and Jens Marklof. Quantum transport in a crystal with short-range interactions: The Boltzmann–Grad limit. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-021-
02797-z.

Guillin:2021:ULT

[GM21b] Arnaud Guillin and Pierre Monmarché. Uniform long-time and propagation of chaos estimates for mean field kinetic particles in non-convex landscapes. Journal of Statistical Physics, 185 (2):??, November 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02839-6.

Giuliani:2020:QIH

[GMP20] Alessandro Giuliani, Vieri Mastropietro, and Marcello Porta. Quantization of the interacting Hall conductivity in the critical regime. Journal of Statistical Physics, 180(1–6):332– 365, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-019-02405-1.pdf.

Gandolfo:2020:GSF

[GMRS20] Daniel Gandolfo, Christian Maes, Jean Ruiz, and Senya Shlosman. Glassy states: The free Ising model on a tree. Journal of Statistical Physics, 180(1–6):227–237, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Grela:2021:NIB

[GMS21] Jacek Grela, Satya N. Majumdar, and Grégory Schehr. Nonintersecting Brownian bridges in the flat-to-flat geometry. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

021-02774-6. See comment [GMS23].

Grela:2023:CNI

[GMS23] Jacek Grela, Satya N. Majumdar, and Grégory Schehr. Comment on “Non-intersecting Brownian Bridges in the Flat-to-Flat Geometry”. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03038-7. See [GMS21].

Gaudilliere:2020:AEL

[GMV20] A. Gaudillière, P. Milanesi, and M. E. Vares. Asymptotic exponential law for the transition time to equilibrium of the metastable kinetic Ising model with vanishing magnetic field. Journal of Statistical Physics, 179(2):263–308, April 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Giorgini:2020:ASA

[GMW20] L. T. Giorgini, W. Moon, and J. S. Wettlaufer. Analytical survival analysis of the Ornstein–Uhlenbeck process. Journal of Statistical Physics, 181(6):2404–2414, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02669-y.pdf.

Gonzalez-Navarrete:2020:MWR

[GN20] Manuel González-Navarrete. Multidimensional walks with random tendency. Journal of Statistical Physics, 181(4):1138–1148, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Giorno:2021:TIF

[GN21] Virginia Giorno and Amelia G. Nobile. Time-inhomogeneous Feller-type diffusion process with absorbing boundary condition. Journal of Statistical Physics, 183(3):??, June 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-
021-02777-3;   http://link.springer.com/content/pdf/
10.1007/s10955-021-02777-3.pdf.

Gonzalez-Navarrete:2021:RRW

[GNH21] Manuel González-Navarrete and Ranghely Hernández. Reinforced random walks under memory lapses. Journal of Statistical Physics, 185(1):??, October 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02826-x.

Godreche:2021:CEF

[God21] Claude Godrèche. Condensation and extremes for a fluctuating number of independent random variables. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Goldfriend:2023:ESM

[Gol23] Tomer Goldfriend. Effective stochastic model for chaos in the Fermi–Pasta–Ulam–Tsingou chain. Journal of Statistical Physics, 190(4):??, April 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03080-z.

Grimmett:2020:BEE

[GOS20] Geoffrey R. Grimmett, Tobias J. Osborne, and Petra F. Scudo. Bounded entanglement entropy in the quantum Ising model. Journal of Statistical Physics, 178(1):281–296, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-019-02432-y.pdf.

Goudon:2023:MPI

[Gou23] Thierry Goudon. A model of particles interacting with thermal traps. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03055-6.

Gallo:2023:CPF

[GP23] Sandro Gallo and Caio Pena. Critical parameter of the frog model on homogeneous trees with geometric lifetime. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-03049-4.

Garcia-Perciante:2020:GIR

[GPRR20] Ana L. García-Perciante, Marcelo E. Rubio, and Oscar A. Reula. Generic instabilities in the relativistic Chapman–Enskog heat conduction law. Journal of Statistical Physics, 181(1):246–262, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gelfert:2022:EIP

[GPS22] Katrin Gelfert, Maria José Pacifico, and Diego Sanhueza. Entropy of irregular points for some dynamical systems. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02973-9.

Gabrielli:2020:DPT

[GR20] Davide Gabrielli and D. R. Michiel Renger. Dynamical phase transitions for flows on finite graphs. Journal of Statistical Physics, 181(6):2353–2371, December 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02667-0.pdf.

Ghosh:2022:CCA

[GR22] Partha Pratim Ghosh and Rahul Roy. Criticality and covered area fraction in confetti and Voronoi percolation. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02859-2.

Greenman:2022:TSP

[Gre22] Chris D. Greenman. Time series path integral expansions for stochastic processes. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02912-8.

Gebert:2020:LTF

[GRM20] Martin Gebert and Constanza Rojas-Molina. Lifshitz tails for the fractional Anderson model. Journal of Statistical Physics, 179(2):341–353, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ghosh:2020:TND

[GS20] Subhroshekhar Ghosh and Kumarjit Saha. Transmission and navigation on disordered lattice networks, directed spanning forests and Brownian Web. Journal of Statistical Physics, 180(1– 6):1167–1205, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Grigorescu:2022:HLB

[GS22] Ilie Grigorescu and Yishu Song. Hydrodynamic limit for the Bak–Sneppen branching diffusions. Journal of Statistical Physics, 187(2):??, May 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02900-y.

Guyader:2020:ELD

[GT20] Arnaud Guyader and Hugo Touchette. Efficient large deviation estimation based on importance sampling. Journal of Statistical Physics, 181(2):551–586, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gianfelice:2020:SSC

[GV20] Michele Gianfelice and Sandro Vaienti. Stochastic stability of the classical Lorenz flow under impulsive type forcing. Journal of Statistical Physics, 181(1):163–211, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Gartner:2020:TBD

[GZ20] Bernd Gärtner and Ahad N. Zehmakan. Threshold behavior of democratic opinion dynamics. Journal of Statistical Physics, 178 (6):1442–1466, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Haba:2023:RCE

[Hab23] Z. Haba. Response of a canonical ensemble of quantum oscillators to a random metric. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03047-6.

Hafouta:2020:AME

[Haf20] Yeor Hafouta. On the asymptotic moments and Edgeworth expansions for some processes in random dynamical environment. Journal of Statistical Physics, 179(4):945–971, May 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hartarsky:2022:BPP

[Har22] Ivailo Hartarsky. Bootstrap percolation, probabilistic cellular automata and sharpness. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02922-6.

Harangi:2023:IRB

[Har23] Viktor Harangi. Improved replica bounds for the independence ratio of random regular graphs. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03062-7.

Hikami:2021:PSC

[HB21] S. Hikami and E. Brézin. Punctures and p-spin curves from matrix models II. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-021-02776-4.

Haig:2022:ACS

[HDS22] Alastair Haig, Fraser Daly, and Seva Shneer. Asymptotics for cliques in scale-free random graphs. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02982-8.

Hepp:2020:STC

[Hep20] K. Hepp. Space, time, categories, mechanics, and consciousness: On Kant and neuroscience. Journal of Statistical Physics, 180(1– 6):896–909, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hurtado:2020:STH

[HG20] Pablo I. Hurtado and Pedro L. Garrido. Simulations of transport in hard particle systems. Journal of Statistical Physics, 180(1– 6):474–533, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Haack:2021:CBM

[HHW21] J. Haack, C. Hauck, and S. Warnecke. A consistent BGK model with velocity-dependent collision frequency for gas mixtures. Journal of Statistical Physics, 184(3):??, September 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02821-2.

Hikami:2022:PSC

[Hik22] Shinobu Hikami. Punctures and p-spin curves from matrix models III. D_{l} type and logarithmic potential. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02950-2.

Hiura:2021:GDS

[Hiu21] Ken Hiura. Gibbs distribution from sequentially predictive form of the Second Law. Journal of Statistical Physics, 185(1):??, October 2021. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-021-02825-y.

Hiura:2022:MDF

[Hiu22] Ken Hiura. Macroscopic dynamical fluctuations in Kac ring model. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02861-8.

Haack:2021:PAQ

[HJ21] Géraldine Haack and Alain Joye. Perturbation analysis of quantum reset models. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02752-y.

He:2023:CPC

[HJ23] Ling-Bing He and Jin-Cheng Jiang. On the Cauchy problem for the cutoff Boltzmann equation with small initial data. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-023-03065-y.

Hou:2023:MLY

[HJN23] Qi Hou, Jianping Jiang, and Charles M. Newman. Motion of Lee–Yang zeros. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03066-x.

He:2020:CAB

[HJZ20] Ling-Bing He, Jin-Cheng Jiang, and Yu-Long Zhou. On the cutoff approximation for the Boltzmann equation with long-range interaction. Journal of Statistical Physics, 181(5):1817–1905, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hislop:2022:LES

[HK22] Peter D. Hislop and M. Krishna. On the local eigenvalue statistics for random band matrices in the localization regime. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02923-5.

Ha:2021:CSS

[HKP21] Seung-Yeal Ha, Dohyun Kim, and Hansol Park. On the completely separable state for the Lohe tensor model. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02750-0.

Ha:2017:EDG

[HKR17] Seung-Yeal Ha, Dongnam Ko, and Sang Woo Ryoo. Emergent dynamics of a generalized Lohe model on some class of Lie groups. Journal of Statistical Physics, 168(1):171–207, July 2017. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [HKR23b].

Ha:2018:RDL

[HKR18] Seung-Yeal Ha, Dongnam Ko, and Sang Woo Ryoo. On the relaxation dynamics of Lohe oscillators on some Riemannian manifolds. Journal of Statistical Physics, 172(5):1427–1478, September 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [HKR23c].

Haimi:2022:ZGW

[HKR22] Antti Haimi, Günther Koliander, and José Luis Romero. Zeros of Gaussian Weyl–Heisenberg functions and hyperuniformity of charge. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02917-3.

Ha:2023:CFC

[HKR23a] Seung-Yeal Ha, Hwa Kil Kim, and Seung-Yeon Ryoo. Correction to: On the finiteness of collisions and phase-locked states for the Kuramoto model. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03149-9.

Ha:2023:CED

[HKR23b] Seung-Yeal Ha, Dongnam Ko, and Seung-Yeon Ryoo. Correction to: Emergent dynamics of a generalized Lohe model on some class of Lie groups. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03150-2. See [HKR17].

Ha:2023:CRD

[HKR23c] Seung-Yeal Ha, Dongnam Ko, and Seung-Yeon Ryoo. Correction to: On the relaxation dynamics of Lohe oscillators on some Riemannian manifolds. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03151-1. See [HKR18].

Henheik:2022:BEG

[HL22] Joscha Henheik and Asbjørn Bækgaard Lauritsen. The BCS energy gap at high density. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02965-9.

Horii:2022:LTA

[HLN22] Hiroshi Horii, Raphaël Lefevere, and Takahiro Nemoto. Large time asymptotic of heavy tailed renewal processes. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02856-5.

Huang:2020:MFL

[HLP20a] Hui Huang, Jian-Guo Liu, and Peter Pickl. On the meanfield limit for the Vlasov–Poisson–Fokker–Planck system. Journal of Statistical Physics, 181(5):1915–1965, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02648-3.pdf.

Hwang:2020:NLC

[HLP^{+}20b] Sungmin Hwang, Enrico Lanza, Giorgio Parisi, Jacopo Rocchi, Giancarlo Ruocco, and Francesco Zamponi. On the number of limit cycles in diluted neural networks. Journal of Statistical Physics, 181(6):2304–2321, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Huang:2020:GSR

[HLW20] Xipei Huang, Lifeng Lin, and Huiqi Wang. Generalized stochastic resonance for a fractional noisy oscillator with random mass and random damping. Journal of Statistical Physics, 178(5): 1201–1216, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Heydenreich:2020:CSP

[HM20] Markus Heydenreich and Kilian Matzke. Critical site percolation in high dimension. Journal of Statistical Physics, 181(3): 816–853, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02607-y.pdf.

Hsu:2023:SPL

[HM23] Alexander Hsu and Sarah E. Marzen. Strange properties of linear reservoirs in the infinitely large limit for prediction of continuoustime signals. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03040-z.

Hollmer:2022:SHD

[HNK22] Philipp Höllmer, Nicolas Noirault, and Werner Krauth. Sparse hard-disk packings and local Markov chains. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02908-4.

Hiraizumi:2022:PGH

[HOiS22] Mao Hiraizumi, Hiroki Ohta, and Shin ichi Sasa. Phase growth with heat diffusion in a stochastic lattice model. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02990-8.

Ha:2020:EBL

[HP20a] Seung-Yeal Ha and Hansol Park. Emergent behaviors of Lohe tensor flocks. Journal of Statistical Physics, 178(5):1268–1292, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ha:2020:SLH

[HP20b] Seung-Yeal Ha and Hansol Park. On the Schrödinger–Lohe hierarchy for aggregation and its emergent dynamics. Journal of Statistical Physics, 181(6):2150–2190, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ha:2021:GFF

[HP21] Seung-Yeal Ha and Hansol Park. On the gradient flow formulation of the Lohe matrix model with high-order polynomial couplings. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02804-3.

Ha:2020:EBT

[HPRS20] Seung-Yeal Ha, Hansol Park, Tommaso Ruggeri, and Woojoo Shim. Emergent behaviors of thermodynamic Kuramoto ensemble on a regular ring lattice. Journal of Statistical Physics, 181 (3):917–943, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ha:2016:EOS

[HR16] Seung-Yeal Ha and Sang Woo Ryoo. On the emergence and orbital stability of phase-locked states for the Lohe model. Journal of Statistical Physics, 163(2):411–439, April 2016. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

016-1481-4. See correction [HR23].

Ha:2023:CEO

[HR23] Seung-Yeal Ha and Seung-Yeon Ryoo. Correction to: On the emergence and orbital stability of phase-locked states for the Lohe model. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03148-w. See [HR16].

Hella:2020:QNA

[HS20a] Olli Hella and Mikko Stenlund. Quenched normal approximation for random sequences of transformations. Journal of Statistical Physics, 178(1):1–37, January 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-019-02390-5.pdf.

Helmuth:2020:LER

[HS20b] Tyler Helmuth and Assaf Shapira. Loop-erased random walk as a spin system observable. Journal of Statistical Physics, 181(4): 1306–1322, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hasler:2021:TIS

[HS21] David Hasler and Oliver Siebert. Thermal ionization for shortrange potentials. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02688-9.pdf.

Hairer:2022:PMS

[HS22] Martin Hairer and Rhys Steele. The Φ_{3}^{4} measure has sub-Gaussian tails. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02866-3.

Higuchi:2020:ECI

[HSS20] Yusuke Higuchi, Mohamed Sabri, and Etsuo Segawa. Electric circuit induced by quantum walk. Journal of Statistical Physics, 181(2):603–617, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Hutchcroft:2022:DMF

[Hut22] Tom Hutchcroft. On the derivation of mean-field percolation critical exponents from the triangle condition. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02967-7.

Hao:2023:ENU

[HXX23] Da-Peng Hao, Zhi-Peng Xun, and Hui Xia. Effect of nonuniform stiffness on tensile avalanche process of fibrous materials. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03021-2.

Iino:2021:BCT

[Iin21] Shumpei Iino. Boundary CFT and tensor network approach to surface critical phenomena of the tricritical 3-state Potts model. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Itoi:2023:URS

[IISS23] C. Itoi, H. Ishimori, K. Sato, and Y. Sakamoto. Universality of replica-symmetry breaking in the transverse field Sherrington– Kirkpatrick model. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-023-03078-7.

Iacobucci:2021:TMT

[IOS21] Alessandra Iacobucci, Stefano Olla, and Gabriel Stoltz. Thermomechanical transport in rotor chains. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02748-8.

Ioffe:2020:IPP

[IOVW20] Dmitry Ioffe, Sébastien Ott, Yvan Velenik, and Vitali Wachtel. Invariance principle for a Potts interface along a wall. Journal of Statistical Physics, 180(1–6):832–861, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Sasa:2022:QSD

[iSHNY22] Shin ichi Sasa, Ken Hiura, Naoko Nakagawa, and Akira Yoshida. Quasi-static decomposition and the Gibbs factorial in small thermodynamic systems. Journal of Statistical Physics, 189 (2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02991-7.

Ioffe:2020:SMS

[IT20] Dmitry Ioffe and Bálint Tóth. Split-and-merge in stationary random stirring on lattice torus. Journal of Statistical Physics, 180(1–6):630–653, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10955-020-02487-2.
pdf.

Jansen:2020:THM

[Jan20] Sabine Jansen. Thermodynamics of a hierarchical mixture of cubes. Journal of Statistical Physics, 179(2):309–340, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [Jan21].

Jansen:2021:CTH

[Jan21] Sabine Jansen. Correction to: Thermodynamics of a Hierarchical Mixture of Cubes. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02742-0. See [Jan20].

Jana:2022:CNH

[Jan22] Indrajit Jana. CLT for non-Hermitian random band matrices with variance profiles. Journal of Statistical Physics, 187(2):??, May 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02892-9.

Jatuviriyapornchai:2020:SCP

[JCG20] Watthanan Jatuviriyapornchai, Paul Chleboun, and Stefan Grosskinsky. Structure of the condensed phase in the inclusion process. Journal of Statistical Physics, 178(3):682–710, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02451-9.pdf.

Jansen:2022:CEN

[JK22] Sabine Jansen and Leonid Kolesnikov. Cluster expansions: Necessary and sufficient convergence conditions. Journal of Statistical Physics, 189(3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02992-6.

Jahnel:2023:TDP

[JK23] Benedikt Jahnel and Jonas Köppl. Trajectorial dissipation of Φ-entropies for interacting particle systems. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03136-0.

Jindal:2021:SSD

[JKG21] Akriti Jindal, A. B. Kolomeisky, and Arvind Kumar Gupta. Steady-state dynamics of exclusion process with local reversible association of particles. Journal of Statistical Physics, 185 (3):??, December 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02843-w.

Jang:2023:VAS

[JKNV23] Jin Woo Jang, Bernhard Kepka, Alessia Nota, and Juan J. L. Velázquez. Vanishing angular singularity limit to the hardsphere Boltzmann equation. Journal of Statistical Physics, 190 (4):??, April 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-023-03089-4.

Jiang:2022:SEM

[JL22] Miaohua Jiang and Marco Lopez. SRB entropy of Markov transformations. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02954-y.

Jex:2023:CDF

[JLM23] Michal Jex, Mathieu Lewin, and Peter S. Madsen. Classical density functional theory: Representability and universal bounds. Journal of Statistical Physics, 190(4):??, April 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03086-7.

Junge:2023:SLS

[JLR23] Marius Junge, Nicholas Laracuente, and Cambyse Rouzé. Stability of logarithmic Sobolev inequalities under a noncommutative change of measure. Journal of Statistical Physics, 190 (2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03026-x.

Jalowy:2022:FMB

[JLS22] Jonas Jalowy, Matthias Löwe, and Holger Sambale. Fluctuations of the magnetization in the block Potts model. Journal of Statistical Physics, 187(1):??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02889-4.

Jiang:2020:EPS

[JLY20] Xiaomeng Jiang, Yong Li, and Xue Yang. Existence of periodic solutions in distribution for stochastic Newtonian systems. Journal of Statistical Physics, 181(2):329–363, October 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Joye:2023:AWW

[JM23] Alain Joye and Marco Merkli. The adiabatic Wigner– Weisskopf model. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03116-4.

Junge:2023:SCM

[JMPR23] Matthew Junge, Zoe McDonald, Jean Pulla, and Lily Reeves. A stochastic combustion model with thresholds on trees. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-023-03102-w.

Johnston:2020:SLN

[JO20] Samuel G. G. Johnston and Neil O’Connell. Scaling limits for non-intersecting polymers and Whittaker measures. Journal of Statistical Physics, 179(2):354–407, April 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02534-y.pdf.

Jones:2022:SRE

[Jon22] Nick G. Jones. Symmetry-resolved entanglement entropy in critical free-fermion chains. Journal of Statistical Physics, 188 (3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02941-3.

Jackson:2022:PSW

[JP22] Andrew D. Jackson and Subodh P. Patil. Phases of small worlds: a mean field formulation. Journal of Statistical Physics, 189 (3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02997-1.

Joy:2021:SPF

[JPR21] Jilmy P. Joy, Sudhir N. Pathak, and R. Rajesh. Shock propagation following an intense explosion: Comparison between hydrodynamics and simulations. Journal of Statistical Physics, 182(2): ??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Joy:2021:SPH

[JR21] Jilmy P. Joy and R. Rajesh. Shock propagation in the hard sphere gas in two dimensions: Comparison between simulations and hydrodynamics. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02790-6.

Janjigian:2020:UES

[JRA20] Christopher Janjigian and Firas Rassoul-Agha. Uniqueness and ergodicity of stationary directed polymers on Z^{2}. Journal of Statistical Physics, 179(3):672–689, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Junior:2020:MTR

[JRS20] Valdivino V. Junior, Pablo M. Rodriguez, and Adalto Speroto. The Maki–Thompson rumor model on infinite Cayley trees. Journal of Statistical Physics, 181(4):1204–1217, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Son:2023:BPM

[jSY23] Sung jun Son and Seok-Bae Yun. The ES–BGK for the polyatomic molecules with infinite energy. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03139-x.

Jahnel:2020:EMP

[JT20] Benedikt Jahnel and András Tóbiás. Exponential moments for planar tessellations. Journal of Statistical Physics, 179(1):90– 109, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See corrections [JT21, JT21].

Jahnel:2021:CEM

[JT21] Benedikt Jahnel and András Tóbiás. Correction to: Exponential Moments for Planar Tessellations. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02743-z. See [JT20].

Junge:2020:CPD

[Jun20a] Matthew Junge. Critical percolation and A + B → 2A dynamics. Journal of Statistical Physics, 181(2):738–751, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Junk:2020:CPF

[Jun20b] Stefan Junk. Comparison of partition functions in a space-time random environment. Journal of Statistical Physics, 181(1):95– 115, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Jang:2022:LNL

[JV22] Jin Woo Jang and Juan J. L. Velázquez. LTE and non-LTE solutions in gases interacting with radiation. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02888-5.

Jindal:2021:CDB

[JVG21] Akriti Jindal, Atul Kumar Verma, and Arvind Kumar Gupta. Cooperative dynamics in bidirectional transport on flexible lattice. Journal of Statistical Physics, 182(1):??, January 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kargin:2020:CRM

[Kar20] Vladislav Kargin. Cycles in random meander systems. Journal of Statistical Physics, 181(6):2322–2345, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kratzer:2023:KLB

[KBK23] Matthew M. Kratzer, Suresh K. Bhatia, and Alexander Y. Klimenko. Knudsen layer behaviour and momentum accommodation from surface roughness modelling. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03075-w.

Kumar:2021:RSB

[KC21] Rohit Kumar and V. V. M. S. Chandramouli. Renormalization of symmetric bimodal maps with low smoothness. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02764-8.

Kumar:2020:IUB

[KD20] Aanjaneya Kumar and Deepak Dhar. Improved upper bounds on the asymptotic growth velocity of Eden clusters. Journal of Statistical Physics, 180(1–6):710–720, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kondratiev:2021:ABS

[KdS21] Yuri Kondratiev and José Luís da Silva. Asymptotic behavior of the subordinated traveling waves. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02745-x.

Kepka:2023:SSP

[Kep23] Bernhard Kepka. Self-similar profiles for homoenergetic solutions of the Boltzmann equation for non-cutoff Maxwell molecules. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03034-x.

Kay:2023:SPR

[KG23] Toby Kay and Luca Giuggioli. Subdiffusion in the presence of reactive boundaries: a generalized Feynman–Kac approach. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03105-7.

Kim:2021:MBC

[Kim21] Seonwoo Kim. Metastability of Blume–Capel model with zero chemical potential and zero external field. Journal of Statistical Physics, 184(3):??, September 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02823-0.

Kim:2020:SLM

[KK20a] Dohyun Kim and Jeongho Kim. Stochastic Lohe matrix model on the Lie group and mean-field limit. Journal of Statistical Physics, 178(6):1467–1514, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kissel:2020:DGN

[KK20b] Sascha Kissel and Christof Külske. Dynamical Gibbs–non-Gibbs transitions in lattice Widom–Rowlinson models with hard-core and soft-core interactions. Journal of Statistical Physics, 178 (3):725–762, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-019-02478-y.pdf.

Khokonov:2021:CSD

[KK21] M. Kh. Khokonov and A. Kh. Khokonov. Cluster size distribution in a system of randomly spaced particles. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Komatsu:2022:SMI

[KK22] Takashi Komatsu and Norio Konno. Stationary measure induced by the eigenvalue problem of the one-dimensional Hadamard walk. Journal of Statistical Physics, 187(1):??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02901-x.

Khoshnevisan:2020:DPS

[KKMS20] Davar Khoshnevisan, Kunwoo Kim, Carl Mueller, and Shang-Yuan Shiu. Dissipation in parabolic SPDEs. Journal of Statistical Physics, 179(2):502–534, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Komatsu:2023:WZC

[KKS23] Takashi Komatsu, Norio Konno, and Iwao Sato. Walk/zeta correspondence. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03052-9.

Koskinen:2020:ELM

[KL20a] Kalle Koskinen and Jani Lukkarinen. Estimation of local microcanonical averages in two lattice mean-field models using coupling techniques. Journal of Statistical Physics, 180(1–6): 1206–1251, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02612-1.pdf.

Kuniba:2020:LDO

[KL20b] Atsuo Kuniba and Hanbaek Lyu. Large deviations and one-sided scaling limit of randomized multicolor box–ball system. Journal of Statistical Physics, 178(1):38–74, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Knopfel:2020:FRG

[KLSS20] Holger Knöpfel, Matthias Löwe, Kristina Schubert, and Arthur Sinulis. Fluctuation results for general block spin Ising models. Journal of Statistical Physics, 178(5):1175–1200, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02489-0.pdf.

Konarovskyi:2020:DKD

[KLvR20] Vitalii Konarovskyi, Tobias Lehmann, and Max von Renesse. On Dean–Kawasaki dynamics with smooth drift potential. Journal of Statistical Physics, 178(3):666–681, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kim:2021:SBM

[KLY21] Doheon Kim, Myeong-Su Lee, and Seok-Bae Yun. Stationary BGK models for chemically reacting gas in a slab. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02810-5.

Kulske:2020:SMP

[KM20] Christof Külske and Daniel Meißner. Stable and metastable phases for the Curie–Weiss–Potts model in vector-valued fields via singularity theory. Journal of Statistical Physics, 181(3): 968–989, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02615-y.pdf.

Kulske:2021:DGN

[KM21] Christof Külske and Daniel Meißner. Dynamical Gibbs–non– Gibbs transitions in the Curie–Weiss Potts model in the regime β < 3. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02793-3.

Khodabandehlou:2022:TFN

[KMN22] Faezeh Khodabandehlou, Christian Maes, and Karel Netocný. Trees and forests for nonequilibrium purposes: an introduction to graphical representations. Journal of Statistical Physics, 189 (3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03003-4.

Kupiainen:2020:SEL

[KO20] Antti Kupiainen and Joona Oikarinen. Stress-energy in Liouville conformal field theory. Journal of Statistical Physics, 180(1–6): 1128–1166, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02601-4.pdf.

Koskinen:2023:IVG

[Kos23] Kalle Koskinen. Infinite volume Gibbs states and metastates of the random field mean-field spherical model. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03076-9.

Krajnik:2020:KPZ

[KP20] Ziga Krajnik and Tomaz Prosen. Kardar–Parisi–Zhang physics in integrable rotationally symmetric dynamics on discrete space– time lattice. Journal of Statistical Physics, 179(1):110–130, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Kennedy:2022:TRA

[KR22a] Tom Kennedy and Slava Rychkov. Tensor RG approach to high-temperature fixed point. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02924-4.

Klausen:2022:MSN

[KR22b] Frederik Ravn Klausen and Aran Raoufi. Mass scaling of the near-critical 2D Ising model using random currents. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-022-
02939-x.

Kumar:2022:BWT

[KR22c] Amit Kumar and R. Rajesh. Blast waves in two and three dimensions: Euler versus Navier–Stokes equations. Journal of Statistical Physics, 188(2):??, August 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02933-3.

Khain:2021:DEC

[KS21] Evgeniy Khain and John Straetmans. Dynamics of an expanding cell monolayer. Journal of Statistical Physics, 184(2):??, August 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02812-3.

Kremer:2022:GGI

[KS22] Gilberto M. Kremer and Andrés Santos. Granular gas of inelastic and rough Maxwell particles. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02984-6.

Koike:2022:DLT

[KTT22] Hajime Koike, Hideki Takayasu, and Misako Takayasu. Diffusion– localization transition point of gravity type transport model on regular ring lattices and Bethe lattices. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02882-x.

Kanazawa:2023:EST

[KTT23] Kiyoshi Kanazawa, Hideki Takayasu, and Misako Takayasu. Exact solution to two-body financial dealer model: Revisited from the viewpoint of kinetic theory. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03022-1.

Kutsenko:2023:AND

[Kut23] Anton A. Kutsenko. Approximation of the number of descendants in branching processes. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-023-03079-6.

Kennerberg:2020:CC

[KV20] Philip Kennerberg and Stanislav Volkov. Convergence in the pcontest. Journal of Statistical Physics, 178(5):1096–1125, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02491-6.pdf.

Kennerberg:2021:LBV

[KV21] Philip Kennerberg and Stanislav Volkov. A local barycentric version of the Bak–Sneppen model. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-021-02718-0.pdf.

Kim:2020:SVA

[KW20] Seungki Kim and Yuntao Wang. A stochastic variant of the Abelian sandpile model. Journal of Statistical Physics, 178(3): 711–724, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Konishi:2023:TDN

[KY23] Tetsuro Konishi and Tatsuo Yanagita. Temperature dependence of non-uniformity in energy of systems with holonomic constraints. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03029-8.

Kovchegov:2020:DPR

[KZ20] Yevgeniy Kovchegov and Ilya Zaliapin. Dynamical pruning of rooted trees with applications to 1-D ballistic annihilation. Journal of Statistical Physics, 181(2):618–672, October 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Laulin:2022:NIR

[Lau22] Lucile Laulin. New insights on the reinforced elephant random walk using a martingale approach. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02834-x.

Li:2022:EGL

[LD22] Yang Li and Jinqiao Duan. Extracting governing laws from sample path data of non-Gaussian stochastic dynamical systems. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02873-y.

Lee:2022:ELM

[Lee22] Jungkyoung Lee. Energy landscape and metastability of Curie– Weiss–Potts model. Journal of Statistical Physics, 187(1): ??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-02897-4.

Lepri:2023:TIH

[Lep23] Stefano Lepri. Thermalization of isolated harmonic networks under conservative noise. Journal of Statistical Physics, 190 (1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03032-z.

Li:2021:ERC

[Li21] Zhongyang Li. Exact recovery of community detection in k - partite graph models with applications to learning electric potentials in electric networks. Journal of Statistical Physics, 182 (1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lima:2020:BMD

[Lim20] Paulo C. Lima. The BEG model in the disordered region and at the antiquadrupolar-disordered line of parameters. Journal of Statistical Physics, 178(1):265–280, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lin:2021:LEH

[Lin21] Yier Lin. Lyapunov exponents of the half-line SHE. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-021-02772-8.

Liu:2020:LTB

[Liu20] Qun Liu. Limit theorems for the bipartite Potts model. Journal of Statistical Physics, 181(6):2071–2093, December 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Liu:2021:FBS

[Liu21] Qun Liu. Fluctuations for the bipartite Sherrington–Kirkpatrick model. Journal of Statistical Physics, 184(1):??, July 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s10955-021-02801-6.

Leonenko:2020:SAF

[LJ20] Nikolai Leonenko and Jayme Vaz Jr. Spectral analysis of fractional hyperbolic diffusion equations with random data. Journal of Statistical Physics, 179(1):155–175, April 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02520-4.pdf.

Li:2020:EFN

[LL20] Jian Li and Chuanzhong Li. Extensions of the finite nonperiodic Toda lattices with indefinite metrics. Journal of Statistical Physics, 179(4):901–919, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lanchier:2022:CHK

[LL22] Nicolas Lanchier and Hsin-Lun Li. Consensus in the Hegselmann– Krause model. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02920-8.

Lee:2023:SSK

[LL23a] Ji Oon Lee and Yiting Li. Spherical Sherrington–Kirkpatrick model for deformed Wigner matrix with fast decaying edges. Journal of Statistical Physics, 190(2):??, February 2023. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic).  URL https://link.springer.com/article/10.
1007/s10955-022-03048-5.

Liu:2023:CES

[LL23b] Bocheng Liu and Xuguang Lu. On the convergence to equilibrium for the spatially homogeneous Boltzmann equation for Fermi–Dirac particles. Journal of Statistical Physics, 190(8): ??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-023-03152-0.

Lemou:2020:NII

[LLM20] M. Lemou, A. M. Luz, and F. Méhats. Nonlinear instability of inhomogeneous steady states solutions to the HMF model. Journal of Statistical Physics, 178(3):645–665, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lopes:2022:BPC

[LLV22] Artur O. Lopes, Silvia R. C. Lopes, and Paulo Varandas. Bayes posterior convergence for loss functions via almost additive thermodynamic formalism. Journal of Statistical Physics, 186(3): ??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-02885-8.

Li:2020:LLN

[LLX20] Songzi Li, Xiang-Dong Li, and Yong-Xiao Xie. On the law of large numbers for the empirical measure process of generalized Dyson Brownian motion. Journal of Statistical Physics, 181(4): 1277–1305, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Li:2022:NSV

[LLY22] Hailiang Li, Shuangqian Liu, and Tong Yang. The Navier– Stokes–Vlasov–Fokker–Planck system in bounded domains. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-022-
02886-7.

Loomis:2020:OQM

[LMAC20] Samuel P. Loomis, John R. Mahoney, Cina Aghamohammadi, and James P. Crutchfield. Optimizing quantum models of classical channels: The reverse Holevo problem. Journal of Statistical Physics, 181(5):1966–1985, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Legrand:2022:SAP

[LP22] Alexandre Legrand and Nicolas Pétrélis. A sharp asymptotics of the partition function for the collapsed interacting partially directed self-avoiding walk. Journal of Statistical Physics, 186(3): ??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-02890-x.

Leschke:2021:FEQ

[LRRS21] Hajo Leschke, Sebastian Rothlauf, Rainer Ruder, and Wolfgang Spitzer. The free energy of a quantum Sherrington–Kirkpatrick spin-glass model for weak disorder. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02689-8.pdf.

Leppanen:2020:SAN

[LS20a] Juho Leppänen and Mikko Stenlund. Sunklodas’ approach to normal approximation for time-dependent dynamical systems. Journal of Statistical Physics, 181(5):1523–1564, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Li:2020:OED

[LS20b] Fucai Li and Baoyan Sun. Optimal exponential decay for the linearized ellipsoidal BGK model in weighted Sobolev spaces. Journal of Statistical Physics, 181(2):690–714, October 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lieb:2020:DEM

[LS20c] Elliott H. Lieb and Robert Seiringer. Divergence of the effective mass of a polaron in the strong coupling limit. Journal of Statistical Physics, 180(1–6):23–33, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-019-02322-3.pdf.

Louidor:2020:SLL

[LS20d] Oren Louidor and Santiago Saglietti. A strong law of large numbers for super-critical branching Brownian motion with absorption. Journal of Statistical Physics, 181(4):1112–1137, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Levine:2021:HFD

[LS21] Lionel Levine and Vittoria Silvestri. How far do activated random walkers spread from a single source? Journal of Statistical Physics, 185(3):??, December 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02836-9.

Lee:2022:NRM

[LS22a] Jungkyoung Lee and Insuk Seo. Non-reversible metastable diffusions with Gibbs invariant measure II: Markov chain convergence. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02986-4.

Luczynska:2022:LTR

[LS22b] Gabriela Luczyńska and Tomasz Szarek. Limits theorems for random walks on Homeo(S^{1}). Journal of Statistical Physics, 187(1):??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02903-9.

Li:2023:SQS

[LS23] Chuanzhong Li and Bao Shou. Supersymmetric quantum spin chains and modified universal characters. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-03063-6.

Lyu:2022:FSP

[LSW22] Ming-Jiea Lyu, Baoyan Sun, and Kung-Chien Wu. Finite speed of propagation of the relativistic Landau and Boltzmann equations. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02850-x.

Lammers:2020:VPW

[LT20] Piet G. Lammers and Martin Tassy. Variational principle for weakly dependent random fields. Journal of Statistical Physics, 179(4):846–870, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Lees:2021:SMP

[LT21] Benjamin Lees and Lorenzo Taggi. Site-monotonicity properties for reflection positive measures with applications to quantum spin systems. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/article/10.
1007/s10955-021-02778-2;   http://link.springer.com/
content/pdf/10.1007/s10955-021-02778-2.pdf.

Lee:2023:DSC

[LT23] Eunghyun Lee and Zhanibek Tokebayev. Distribution of a second-class particle’s position in the two-species ASEP with a special initial configuration. Journal of Statistical Physics, 190(4):??, April 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-023-03085-8.

Lucarini:2020:ISI

[Luc20a] Valerio Lucarini. Introduction to the special issue on the statistical mechanics of climate. Journal of Statistical Physics, 179 (5–6):997–1009, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02605-0.pdf.

Luczka:2020:QCC

[Luc20b] Jerzy Luczka. Quantum counterpart of classical equipartition of energy. Journal of Statistical Physics, 179(4):839–845, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02557-5.pdf.

Leplaideur:2020:CWT

[LW20] Renaud Leplaideur and Frédérique Watbled. Curie–Weiss type models for general spin spaces and quadratic pressure in ergodic theory. Journal of Statistical Physics, 181(1):263–292, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Liu:2023:CPT

[LXZ23a] Yao Liu, Yingchao Xie, and Mengge Zhang. Correction: Phase transitions for a class of time-inhomogeneous diffusion processes. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03070-1. See [LXZ23b].

Liu:2023:PTC

[LXZ23b] Yao Liu, Yingchao Xie, and Mengge Zhang. Phase transitions for a class of time-inhomogeneous diffusion processes. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-03054-7.

See affiliation correction [LXZ23a].

Liao:2021:SGM

[LY21] Jie Liao and Xiongfeng Yang. Stability of global Maxwellian for fully nonlinear Fokker–Planck equations. Journal of Statistical Physics, 185(3):??, December 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02844-9.

Lychkovskiy:2020:RNI

[Lyc20] Oleg Lychkovskiy. A remark on the notion of independence of quantum integrals of motion in the thermodynamic limit. Journal of Statistical Physics, 178(4):1028–1038, February 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Li:2023:CBC

[LYZ23] Ruo Li, Yichen Yang, and Yizhou Zhou. Construction of boundary conditions for Navier–Stokes equations from the moment system. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03115-5.

Madras:2023:ALP

[Mad23] Neal Madras. Adsorption of lattice polymers with quenched topologies. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03006-1.

Maes:2021:SMF

[Mae21] Christian Maes. Statistical mechanical foundation of Weber– Fechner laws. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Majumdar:2023:DMP

[Maj23] Debjyoti Majumdar. DNA melting in poor solvent. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-03025-y.

Mastropietro:2022:VDW

[Mas22] Vieri Mastropietro. Vanishing of Drude weight in interacting fermions on Z^{d} with quasi-periodic disorder. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02868-1.

Mendl:2021:ENE

[MB21] Christian B. Mendl and Folkmar Bornemann. Efficient numerical evaluation of thermodynamic quantities on infinite (semi- )classical chains. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-021-02736-y.pdf.

Meoli:2020:FBP

[MBL20] Alessandra Meoli, Niko Beerenwinkel, and Mykola Lebid. The fractional birth process with power-law immigration. Journal of Statistical Physics, 178(3):775–799, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Marzen:2016:PRD

[MC16] Sarah E. Marzen and James P. Crutchfield. Predictive ratedistortion for infinite-order Markov processes. Journal of Statistical Physics, 163(6):1312–1338, June 2016. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

URL http://link.springer.com/article/10.1007/s10955-

016-1520-1. See correction [MC21].

Marzen:2021:CPR

[MC21] Sarah E. Marzen and James P. Crutchfield. Correction to: Predictive Rate-Distortion for Infinite-Order Markov Processes. Journal of Statistical Physics, 182(2):??, February 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-021-02698-1.pdf. See [MC16].

Miangolarra:2023:NIC

[MC23] Ander Movilla Miangolarra and Michele Castellana. On nonideal chemical-reaction networks and phase separation. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-022-
03037-8.

McDonald:2020:FGS

[McD20] N. R. McDonald. Finger growth and selection in a Poisson field. Journal of Statistical Physics, 178(3):763–774, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02454-6.pdf.

Menegaki:2020:QRC

[Men20] Angeliki Menegaki. Quantitative rates of convergence to nonequilibrium steady state for a weakly anharmonic chain of oscillators. Journal of Statistical Physics, 181(1):53–94, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02565-5.pdf.

Meoli:2023:SPB

[Meo23] Alessandra Meoli. Some Poisson-based processes at geometric times. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03117-3.

Monemvassitis:2023:PCE

[MGM23] Athina Monemvassitis, Arnaud Guillin, and Manon Michel. PDMP characterisation of event-chain Monte Carlo algorithms for particle systems. Journal of Statistical Physics, 190(3): ??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-023-03069-8.

Mendez:2020:TDT

[MGPCA20] A. R. Méndez, A. L. García-Perciante, and G. Chacón-Acosta. Thermal dissipation in two dimensional relativistic Fermi gases with a relaxation time model. Journal of Statistical Physics, 178 (4):936–953, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mendez:2022:DPD

[MGPCA22] A. R. Méndez, A. L. García-Perciante, and G. Chacón-Acosta. Dissipative properties of degenerate relativistic gases: The complete kernel calculation in a (d + 1) flat space-time. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02883-w.

Mason:2023:MBT

[MJB23] James Mason, Robert L. Jack, and Maria Bruna. Macroscopic behaviour in a two-species exclusion process via the method of matched asymptotics. Journal of Statistical Physics, 190(3): ??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-03036-9.

Maas:2020:MCR

[MM20a] Jan Maas and Alexander Mielke. Modeling of chemical reaction systems with detailed balance using gradient structures. Journal of Statistical Physics, 181(6):2257–2303, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02663-4.pdf.

Mathiaud:2020:BFP

[MM20b] J. Mathiaud and L. Mieussens. BGK and Fokker–Planck models of the Boltzmann equation for gases with discrete levels of vibrational energy. Journal of Statistical Physics, 178(5):1076– 1095, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Medvedev:2021:SCS

[MM21] Georgi S. Medvedev and Mathew S. Mizuhara. Stability of clusters in the second-order Kuramoto model on random graphs. Journal of Statistical Physics, 182(2):??, February 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Medvedev:2022:CU

[MM22] Georgi S. Medvedev and Matthew S. Mizuhara. Chimeras unfolded. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02881-y.

Menguturk:2023:PTC

[MM23] Levent Ali Mengütürk and Murat Cahit Mengütürk. Piecewisetunneled captive processes and corridored random particle systems. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02996-2.

Mukhopadhyay:2020:VMD

[MMR20] Arpan Mukhopadhyay, Ravi R. Mazumdar, and Rahul Roy. Voter and majority dynamics with biased and stubborn agents. Journal of Statistical Physics, 181(4):1239–1265, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-020-02625-w.pdf.

Mohammadpour:2022:LSP

[Moh22] Reza Mohammadpour. Lyapunov spectrum properties and continuity of the lower joint spectral radius. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02910-w.

Molchan:2020:LEP

[Mol20] G. Molchan. Leadership exponent in the pursuit problem for 1-D random particles. Journal of Statistical Physics, 181(3): 952–967, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Molchan:2022:PEG

[Mol22] G. Molchan. The persistence exponents of Gaussian random fields connected by the Lamperti transform. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02864-5.

Morfe:2020:STC

[Mor20a] Peter S. Morfe. Surface tension and Γ-convergence of van der Waals–Cahn–Hilliard phase transitions in stationary ergodic media. Journal of Statistical Physics, 181(6):2225–2256, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Moriya:2020:GVF

[Mor20b] Hajime Moriya. Gibbs variational formula for thermal equilibrium states in terms of quantum relative entropy density. Journal of Statistical Physics, 181(3):761–771, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Moud:2023:PDS

[Mou23] Aref Abbasi Moud. Precise determination of the saturation coverage of polygons in silico using exclusion assisted packing technique. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03097-4.

Michelen:2022:SSM

[MP22] Marcus Michelen and Will Perkins. Strong spatial mixing for repulsive point processes. Journal of Statistical Physics, 189 (1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02969-5.

Mankar:2020:DNU

[MPDH20] Praful D. Mankar, Priyabrata Parida, Harpreet S. Dhillon, and Martin Haenggi. Distance from the nucleus to a uniformly random point in the 0-cell and the typical cell of the Poisson– Voronoi tessellation. Journal of Statistical Physics, 181(5):1678– 1698, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Michelen:2020:QSB

[MPR20] Marcus Michelen, Robin Pemantle, and Josh Rosenberg. Quenched survival of Bernoulli percolation on Galton–Watson trees. Journal of Statistical Physics, 181(4):1323–1364, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mays:2021:TWD

[MPS21] Anthony Mays, Anita Ponsaing, and Grégory Schehr. Tracy– Widom distributions for the Gaussian orthogonal and symplectic ensembles revisited: a skew-orthogonal polynomials approach. Journal of Statistical Physics, 182(2):??, February 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Majda:2020:SPT

[MQ20] Andrew J. Majda and Di Qi. Statistical phase transitions and extreme events in shallow water waves with an abrupt depth change. Journal of Statistical Physics, 179(5–6):1718–1741, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mukhamedov:2021:DQM

[MS21] Farrukh Mukhamedov and Abdessatar Souissi. Diagonalizability of quantum Markov states on trees. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mukherjee:2022:PTB

[MS22a] Soheli Mukherjee and Sumedha. Phase transitions in the Blume– Capel model with trimodal and Gaussian random fields. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-022-
02949-9.

Mysliwy:2022:PMR

[MS22b] Krzysztof Myśliwy and Robert Seiringer. Polaron models with regular interactions at strong coupling. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02851-w.

Marangio:2020:AMN

[MSG^{+}20] L. Marangio, J. Sedro, S. Galatolo, A. Di Garbo, and M. Ghil. Arnold maps with noise: Differentiability and non-monotonicity of the rotation number. Journal of Statistical Physics, 179(5– 6):1594–1624, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Mlotkowski:2020:FSD

[MSU20] Wojciech Młotkowski, Noriyoshi Sakuma, and Yuki Ueda. Free self-decomposability and unimodality of the Fuss–Catalan distributions. Journal of Statistical Physics, 178(5):1055–1075, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Miyazaki:2020:LTL

[MT20] Tatsuya Miyazaki and Masato Takei. Limit theorems for the ‘laziest’ minimal random walk model of elephant type. Journal of Statistical Physics, 181(2):587–602, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Manai:2020:PDQ

[MW20a] Chokri Manai and Simone Warzel. Phase diagram of the quantum random energy model. Journal of Statistical Physics, 180(1– 6):654–664, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02492-5.pdf.

Mustafa:2020:BEL

[MW20b] Dawan Mustafa and Bernt Wennberg. The BGK equation as the limit of an N-particle system. Journal of Statistical Physics, 181 (2):715–737, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02596-y.pdf.

Manai:2022:ATL

[MW22] Chokri Manai and Simone Warzel. The de Almeida–Thouless line in hierarchical quantum spin glasses. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02860-9.

Nascimento:2022:MPT

[Nas22] A. M. B. Nascimento. Multiple phase transitions for an infinite system of spiking neurons. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02931-5.

Naze:2022:SEE

[NB22] Pierre Nazé and Marcus V. S. Bonança. Series expansion of the excess work using nonlinear response theory. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02869-0.

Neirotti:2021:SEO

[Nei21] Juan Neirotti. Strategies for an efficient official publicity campaign. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02761-x.

Nascimento:2020:CTE

[NF20a] A. M. B. Nascimento and L. R. Fontes. Convergence time to equilibrium of the Metropolis dynamics for the GREM. Journal of Statistical Physics, 178(1):297–317, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nguyen:2020:CCV

[NF20b] Tong Xuan Nguyen and Roberto Fernández. Convergence of cluster and virial expansions for repulsive classical gases. Journal of Statistical Physics, 179(2):448–484, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nguyen:2022:SER

[Ngu22] Tuan-Minh Nguyen. Speed of excited random walks with long backward steps. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02926-2.

Nishikawa:2021:RDN

[NIB21] Yoshihiko Nishikawa, Atsushi Ikeda, and Ludovic Berthier. Relaxation dynamics of non–Brownian spheres below jamming. Journal of Statistical Physics, 182(2):??, February 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nisoli:2023:HDN

[Nis23] Isaia Nisoli. How does noise induce order? Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-03041-y.

Neiss:2020:MFL

[NP20] R. A. Neiss and P. Pickl. A mean field limit for the Hamiltonian Vlasov system. Journal of Statistical Physics, 178(2):472–498, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Nandi:2021:ODE

[NP21] Riya Nandi and Priyanka. One dimensional exclusion process with dynein inspired hops: Simulation and mean field analysis. Journal of Statistical Physics, 182(2):??, February 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Napiorkowski:2020:NUC

[NPT20] M. Napiórkowski, J. Piasecki, and J. W. Turner. Non-universal Casimir forces at approach to Bose–Einstein condensation of an ideal gas: Effect of Dirichlet boundary conditions. Journal of Statistical Physics, 181(3):944–951, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02613-0.pdf.

Notarnicola:2023:FCB

[NPV23] Massimo Notarnicola, Giovanni Peccati, and Anna Vidotto. Functional convergence of Berry’s nodal lengths: Approximate tightness and total disorder. Journal of Statistical Physics, 190 (5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-023-03111-9.

Nachtergaele:2021:SPS

[NR21] Bruno Nachtergaele and Jake Reschke. Slow propagation in some disordered quantum spin chains. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Neeman:2020:PTF

[NRS20] Joe Neeman, Charles Radin, and Lorenzo Sadun. Phase transitions in finite random networks. Journal of Statistical Physics, 181(1):305–328, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Niggemann:2020:FTT

[NS20] Oliver Niggemann and Udo Seifert. Field-theoretic thermodynamic uncertainty relation. Journal of Statistical Physics, 178 (5):1142–1174, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-019-02479-x.pdf.

Niggemann:2021:NST

[NS21] Oliver Niggemann and Udo Seifert. Numerical study of the thermodynamic uncertainty relation for the KPZ–Equation. Journal of Statistical Physics, 182(2):??, February 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02692-z.pdf.

Niggemann:2022:TSR

[NS22] Oliver Niggemann and Udo Seifert. The two scaling regimes of the thermodynamic uncertainty relation for the KPZ-equation. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-021-
02845-8.

Nakano:2020:PSB

[NT20] Fumihiko Nakano and Khanh Duy Trinh. Poisson statistics for beta ensembles on the real line at high temperature. Journal of Statistical Physics, 179(2):632–649, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ni:2023:RDF

[NT23] Angxiu Ni and Yao Tong. Recursive divergence formulas for perturbing unstable transfer operators and physical measures. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03138-y.

Oliveira:2022:PFD

[OMM22] Guilherme Eduardo Freire Oliveira, Christian Maes, and Kasper Meerts. Photon frequency diffusion process. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02960-0.

Ou:2022:SRF

[ORD22] Hong-Lei Ou, Ruibin Ren, and Ke Deng. Stochastic resonance in a fractional oscillator with cross-correlation noise. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02934-2.

Otomo:2023:NLQ

[Oto23] Hiroshi Otomo. A non-local quasi-equilibrium state in the Bhatnagar–Gross–Krook Boltzmann equation for thermohydrodynamics: Conservation laws, the Boltzmann H-theorem, and the fluctuation–dissipation theorem. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03119-1.

Papageorgiou:2020:MLS

[Pap20] Ioannis Papageorgiou. Modified log-Sobolev inequality for a compact pure jump Markov process with degenerate jumps. Journal of Statistical Physics, 178(6):1293–1318, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Parida:2022:MRS

[PD22] Priyabrata Parida and Harpreet S. Dhillon. Multilayer random sequential adsorption. Journal of Statistical Physics, 187(1): ??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-02896-5.

Pechmann:2022:BEC

[Pec22] Maximilian Pechmann. On Bose–Einstein condensation in onedimensional noninteracting Bose gases in the presence of soft Poisson obstacles. Journal of Statistical Physics, 189(3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-022-03001-6.

Peliti:2021:RMC

[Pel21] Luca Peliti. A review of: Massimo Cencini, Andrea Puglisi, Davide Vergni, Angelo Vulpiani: A Random Walk in Physics: Beyond Black Holes and Time-Travels. Journal of Statistical Physics, 185(2):??, November 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02808-z.

Penrose:2020:MIL

[Pen20] Oliver Penrose. Microscopic irreversibility: Looking for a microscopic description of time asymmetry. Journal of Statistical Physics, 180(1–6):862–872, September 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02548-6.pdf.

Phillips:2020:HES

[Phi20] Michael Phillips. Hysteresis effects in social behavior with parasitic infection. Journal of Statistical Physics, 181(1):293–304, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pierini:2020:SSS

[Pie20] Stefano Pierini. Statistical significance of small ensembles of simulations and detection of the internal climate variability: an excitable ocean system case study. Journal of Statistical Physics, 179(5–6):1475–1495, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Piguet:2023:BBC

[Pig23] Olivier Piguet. De Broglie–Bohm cycles. free relativistic onehalf particles. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03137-z.

Pinto:2023:ZTL

[Pin23] Nicolás Pinto. Zero temperature limits for quotients of potentials in countable Markov shifts. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-023-03106-6.

Pavelka:2020:GDL

[PKG20] Michal Pavelka, Václav Klika, and Miroslav Grmela. Generalization of the dynamical lack-of-fit reduction from GENERIC to GENERIC. Journal of Statistical Physics, 181(1):19–52, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Park:2023:BSM

[PKTM23] Su-Chan Park, Joachim Krug, Léo Touzo, and Peter Mörters. Branching with selection and mutation I: Mutant fitness of Fréchet type. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03125-3.

Pons:2020:SHE

[PMACF20] Flavio Maria Emanuele Pons, Gabriele Messori, M. Carmen Alvarez-Castro, and Davide Faranda. Sampling hyperspheres via extreme value theory: Implications for measuring attractor dimensions. Journal of Statistical Physics, 179(5–6):1698–1717, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pohl:2023:BDT

[PN23] Lorena Pohl and Barbara Niethammer. A Becker–Döring type model for cell polarization. Journal of Statistical Physics, 190 (8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-023-03144-0.

Parvaneh:2022:DND

[PPR22] Azadeh Parvaneh, Afshin Parvardeh, and Rahul Roy. A drainage network with dependence and the Brownian web. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02978-4.

Pennisi:2020:CLR

[PR20] Sebastiano Pennisi and Tommaso Ruggeri. Classical limit of relativistic moments associated with Boltzmann–Chernikov equation: Optimal choice of moments in classical theory. Journal of Statistical Physics, 179(1):231–246, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Podder:2021:UTF

[PR21] Moumanti Podder and Leonardo T. Rolla. Uniform threshold for fixation of the stochastic sandpile model on the line. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Pogorui:2023:TPH

[PRD23] Anatoliy A. Pogorui and Ramón M. Rodríguez-Dagnino. Telegraph process on a hyperbola. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-023-03133-3.

Petrov:2020:PSP

[PT20] Leonid Petrov and Mikhail Tikhonov. Parameter symmetry in perturbed GUE corners process and reflected drifted Brownian motions. Journal of Statistical Physics, 181(5):1996–2010, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Peccati:2020:GRM

[PV20] Giovanni Peccati and Anna Vidotto. Gaussian random measures generated by Berry’s nodal sets. Journal of Statistical Physics, 178(4):996–1027, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Petrache:2023:ASR

[PV23] Mircea Petrache and Rodolfo Viera. Almost sure recovery in quasi-periodic structures. Journal of Statistical Physics, 190 (2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03059-2.

Procaccia:2020:SDW

[PYZ20] Eviatar B. Procaccia, Jiayan Ye, and Yuan Zhang. Stationary DLA is well defined. Journal of Statistical Physics, 181(4): 1089–1111, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Qi:2021:MVS

[Qi21] Kunlun Qi. On the measure valued solution to the inelastic Boltzmann equation with soft potentials. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02762-w.

Ragone:2020:CEV

[RB20] Francesco Ragone and Freddy Bouchet. Computation of extreme values of time averaged observables in climate models with large deviation techniques. Journal of Statistical Physics, 179 (5–6):1637–1665, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ruiz:2020:DAW

[RCF20] Diego Ruiz, Juan Campos, and Jorge Finke. Dynamics in affinity–weighted preferential attachment networks. Journal of Statistical Physics, 181(2):673–689, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Ross:2021:BCD

[RF21] Robert J. H. Ross and Walter Fontana. Balancing conservative and disruptive growth in the voter model. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02749-7.

Rodrigues:2022:EPA

[RJS22] Fagner B. Rodrigues, Thomas Jacobus, and Marcus V. Silva. Entropy points and applications for free semigroup actions. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-021-
02858-3.

Romaro:2023:NST

[RNA23] C. Romaro, F. A. Najman, and M. André. A numerical study of the time of extinction in a class of systems of spiking neurons. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03060-9.

Rodrigues:2021:DRW

[Rod21] Alexandre A. P. Rodrigues. Dissecting a resonance wedge on heteroclinic bifurcations. Journal of Statistical Physics, 184 (3):??, September 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02811-4.

Rozikov:2022:MSH

[Roz22] U. A. Rozikov. Mirror symmetry of height-periodic gradient Gibbs measures of an SOS model on Cayley trees. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02953-z.

Rahmatullaev:2023:GSG

[RRA23] Muzaffar M. Rahmatullaev, Muhayyo A. Rasulova, and Javohir N. Asqarov. Ground states and Gibbs measures of Ising model with competing interactions and an external field on a Cayley tree. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03129-z.

Rademacher:2022:LDE

[RS22] Simone Rademacher and Robert Seiringer. Large deviation estimates for weakly interacting bosons. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02940-4.

Ruelle:2020:ICF

[Rue20] David Ruelle. Identities for correlation functions in classical statistical mechanics and the problem of crystal states. Journal of Statistical Physics, 180(1–6):1002–1009, September 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Robert:2021:APM

[RV21] Philippe Robert and Gaëtan Vignoud. Averaging principles for Markovian models of plasticity. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/article/10.1007/s10955-021-02785-3.

Ryan:2021:MLM

[Rya21] Kieran Ryan. The Manhattan and Lorentz mirror models: a result on the cylinder with low density of mirrors. Journal of Statistical Physics, 185(2):??, November 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02837-8.

Roelly:2020:MGP

[RZ20] Sylvie Roelly and Alexander Zass. Marked Gibbs point processes with unbounded interaction: an existence result. Journal of Statistical Physics, 179(4):972–996, May 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02559-3.pdf. See correction [RZ22].

Roelly:2022:CMG

[RZ22] Sylvie Roelly and Alexander Zass. Correction to: Marked Gibbs point processes with unbounded interaction: an existence result. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02972-w. See [RZ20].

Sakagawa:2021:BLG

[Sak21] Hironobu Sakagawa. Behavior of the lattice Gaussian free field with weak repulsive potentials. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Samaj:2020:ALC

[Sam20a] Ladislav Samaj. Attraction of like-charged walls with counterions only: Exact results for the 2D cylinder geometry. Journal of Statistical Physics, 181(5):1699–1729, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Samaj:2020:SDS

[Sam20b] Ladislav Samaj. Short-distance symmetry of pair correlations in two-dimensional jellium. Journal of Statistical Physics, 178 (1):247–264, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Shivam:2021:SVP

[SBB^{+}21] Saumya Shivam, Christopher L. Baldwin, John Barton, Mehran Kardar, and S. L. Sondhi. Studying viral populations with tools from quantum spin chains. Journal of Statistical Physics, 182(2): ??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Singh:2023:BWZ

[SCDK23] Sahil Kumar Singh, Subhadip Chakraborti, Abhishek Dhar, and P. L. Krapivsky. Blast waves in the zero temperature hard sphere gas: Double scaling structure. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03127-1.

Scola:2021:LMP

[Sco21] Giuseppe Scola. Local moderate and precise large deviations via cluster expansions. Journal of Statistical Physics, 183(1): ??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-021-02740-2.

Sone:2021:JEC

[SD21] Akira Sone and Sebastian Deffner. Jarzynski equality for conditional stochastic work. Journal of Statistical Physics, 183(1): ??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print),

1572-9613 (electronic). URL https://link.springer.com/
article/10.1007/s10955-021-02720-6.

Suen:2022:SSC

[SEG22] Whei Yeap Suen, Thomas J. Elliott, and Mile Gu. Surveying structural complexity in quantum many-body systems. Journal of Statistical Physics, 187(1):??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02895-6.

Seo:2020:ESL

[Seo20] Seong-Mi Seo. Edge scaling limit of the spectral radius for random normal matrix ensembles at hard edge. Journal of Statistical Physics, 181(5):1473–1489, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Serva:2020:RML

[Ser20] Maurizio Serva. Random motion of light-speed particles. Journal of Statistical Physics, 181(5):1603–1608, December 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Schorlepp:2023:SZM

[SGG23] Timo Schorlepp, Tobias Grafke, and Rainer Grauer. Symmetries and zero modes in sample path large deviations. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-03051-w.

Song:2020:LMD

[SGNS20] Eric Yilun Song, Reza Gheissari, Charles M. Newman, and Daniel L. Stein. Local minima in disordered mean-field ferromagnets. Journal of Statistical Physics, 180(1–6):576–596, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Sasom:2023:RWQ

[SH23] Srawut Sasom and Varagorn Hengpunya. Random walks on quasi-one-dimensional lattices. Journal of Statistical Physics, 190(5):??, May 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-023-03113-7.

Shapira:2020:NSG

[Sha20] Assaf Shapira. A note on the spectral gap of the Fredrickson– Andersen one spin facilitated model. Journal of Statistical Physics, 181(6):2346–2352, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

http://link.springer.com/content/pdf/10.1007/s10955-
020-02666-1.pdf.

Shcherbina:2020:CPR

[Shc20] Tatyana Shcherbina. Characteristic polynomials for random band matrices near the threshold. Journal of Statistical Physics, 179(4):920–944, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Shiraishi:2021:OTU

[Shi21] Naoto Shiraishi. Optimal thermodynamic uncertainty relation in Markov jump processes. Journal of Statistical Physics, 185 (3):??, December 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02829-8.

Sugimoto:2023:ETH

[SHRE23] Shoki Sugimoto, Joscha Henheik, Volodymyr Riabov, and László Erdős. Eigenstate thermalisation hypothesis for translation invariant spin systems. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03132-4.

Sampat:2020:DPM

[SKM20] Pranay Bimal Sampat, Sameer Kumar, and Shradha Mishra. Dynamics of a particle moving in a two dimensional Lorentz lattice gas. Journal of Statistical Physics, 181(5):1986–1995, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Shi:2023:CPS

[SLZ23] Yushi Shi, Ting Li, and Jiandong Zhu. Complete phase synchronization of nonidentical high-dimensional Kuramoto model. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03023-0.

Sanchez-Mendoza:2023:PEL

[SM23a] Daniel Sánchez-Mendoza. Principal eigenvalue and landscape function of the Anderson model on a large Box. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03130-6.

Sun:2023:NNN

[SM23b] Hui Sun and Nicholas J. Moore. On normal and non-normal wave statistics implied by a canonical–microcanonical Gibbs ensemble of the truncated KdV system. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02999-z.

Shabbir:2022:CAT

[SN22] Ayesha Shabbir and Muhammad Faisal Nadeem. Computational analysis of topological index-based entropies of carbon nanotube Y-junctions. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02955-x.

Santhosh:2020:AIN

[SND^{+}20] Sreejith Santhosh, Mehrana R. Nejad, Amin Doostmohammadi, Julia M. Yeomans, and Sumesh P. Thampi. Activity induced nematic order in isotropic liquid crystals. Journal of Statistical Physics, 180(1–6):699–709, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Son:2020:RCS

[Son20] Ta Cong Son. The rate of convergence for the Smoluchowski– Kramers approximation for stochastic differential equations with FBM. Journal of Statistical Physics, 181(5):1730–1745, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Soshnikov:2023:GFS

[Sos23] Alexander Soshnikov. Gaussian fluctuation for smoothed local correlations in CUE. Journal of Statistical Physics, 190 (1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-03028-9.

Sardar:2020:STV

[SPL20] Muhammad Shoaib Sardar, Xiang-Feng Pan, and Yun-Xiang Li. Some two-vertex resistances of the three-towers Hanoio graph formed by a fractal graph. Journal of Statistical Physics, 181 (1):116–131, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Spohn:2020:GGE

[Spo20] Herbert Spohn. Generalized Gibbs ensembles of the classical Toda chain. Journal of Statistical Physics, 180(1–6):4–22, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Shao:2021:CUP

[SS21] Shuai Shao and Yuxin Sun. Contraction: a unified perspective of correlation decay and zero-freeness of 2-spin systems. Journal of Statistical Physics, 185(2):??, November 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02831-0.

Shcherbina:2022:LSV

[SS22] Mariya Shcherbina and Tatyana Shcherbina. The least singular value of the general deformed Ginibre ensemble. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02989-1.

Stanislavova:2023:LTD

[SS23] Milena Stanislavova and Atanas G. Stefanov. On the long time dynamics of the Landau–De Gennes gradient flow. Journal of Statistical Physics, 190(1):??, January 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-03046-7.

Shapoval:2021:PSB

[SSS21] Alexander Shapoval, Dayana Savostianova, and Mikhail Shnirman. Predictability and scaling in a BTW sandpile on a selfsimilar lattice. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02744-y.

Schmidt:2021:LGM

[SSvB21] J. Schmidt, G. M. Schütz, and H. van Beijeren. A lattice gas model for generic one-dimensional Hamiltonian systems. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-021-02709-1.

Selley:2022:SNG

[ST22] Fanni M. Sélley and Matteo Tanzi. Synchronization for networks of globally coupled maps in the thermodynamic limit. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02968-6.

Stegehuis:2022:DPL

[Ste22] Clara Stegehuis. Distinguishing power-law uniform random graphs from inhomogeneous random graphs through small subgraphs. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02884-9.

Subag:2022:CFE

[Sub22] Eliran Subag. Convergence of the free energy for spherical spin glasses. Journal of Statistical Physics, 189(2):??, November 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02988-2.

Seiringer:2020:EHP

[SY20] Robert Seiringer and Jakob Yngvason. Emergence of Haldane pseudo-potentials in systems with short-range interactions. Journal of Statistical Physics, 181(2):448–464, October 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02586-0.pdf.

Takahasi:2020:UMC

[Tak20] Hiroki Takahasi. Uniqueness of minimizer for countable Markov shifts and equidistribution of periodic points. Journal of Statistical Physics, 181(6):2415–2431, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Takahasi:2023:LLD

[Tak23] Hiroki Takahasi. Level-2 large deviation principle for countable Markov shifts without Gibbs states. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-023-03126-2.

Tanaka:2020:ECC

[Tan20] Akinori Tanaka. An extension of the cell-construction method for the flat-band ferromagnetism. Journal of Statistical Physics, 181(3):897–916, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tanogami:2022:VSF

[Tan22] Tomohiro Tanogami. Violation of the second fluctuationdissipation relation and entropy production in nonequilibrium medium. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02921-7.

Tasaki:2020:SSB

[Tas20] Hal Tasaki. Spontaneous symmetry breaking in coupled Bose– Einstein condensates. Journal of Statistical Physics, 178(2):379– 391, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tel:2020:TPC

[TBD^{+}20] T. Tél, T. Bódai, G. Drótos, T. Haszpra, M. Herein, B. Kaszás, and M. Vincze. The theory of parallel climate realizations. Journal of Statistical Physics, 179(5–6):1496–1530, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-019-02445-7.pdf.

Tantet:2020:RPRa

[TCDN20] Alexis Tantet, Mickaël D. Chekroun, Henk A. Dijkstra, and J. David Neelin. Ruelle–Pollicott resonances of stochastic systems in reduced state space. Part II: Stochastic Hopf bifurcation. Journal of Statistical Physics, 179(5–6):1403–1448, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tantet:2020:RPRb

[TCND20] Alexis Tantet, Mickaël D. Chekroun, J. David Neelin, and Henk A. Dijkstra. Ruelle–Pollicott resonances of stochastic systems in reduced state space. Part III: Application to the Cane– Zebiak model of the El Niño–Southern Oscillation. Journal of Statistical Physics, 179(5–6):1449–1474, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tondeur:2020:TSS

[TCVB20] Maxime Tondeur, Alberto Carrassi, Stephane Vannitsem, and Marc Bocquet. On temporal scale separation in coupled data assimilation with the ensemble Kalman filter. Journal of Statistical Physics, 179(5–6):1161–1185, June 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02525-z.pdf.

Tan:2020:BEB

[TD20] Nguyen Van Tan and Nguyen Tien Dung. A Berry–Esseen bound in the Smoluchowski–Kramers approximation. Journal of Statistical Physics, 179(4):871–884, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Texier:2020:FPR

[Tex20] Christophe Texier. Fluctuations of the product of random matrices and generalized Lyapunov exponent. Journal of Statistical Physics, 181(3):990–1051, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Takata:2022:SBM

[TH22] Shigeru Takata and Masanari Hattori. Singular behavior of the macroscopic quantity near the boundary for a Lorentz-gas model with the infinite-range potential. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02959-7.

Takahashi:2020:FCS

[THFH20] Kazutaka Takahashi, Yuki Hino, Keisuke Fujii, and Hisao Hayakawa. Full counting statistics and fluctuation–dissipation relation for periodically driven two-state systems. Journal of Statistical Physics, 181(6):2206–2224, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Tamura:2021:FDH

[TK21] Kensuke Tamura and Hosho Katsura. Ferromagnetism in ddimensional SU(n) Hubbard models with nearly flat bands. Journal of Statistical Physics, 182(1):??, January 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02687-w.pdf.

Takata:2018:SKM

[TN18] Shigeru Takata and Takashi Noguchi. A simple kinetic model for the phase transition of the van der Waals fluid. Journal of Statistical Physics, 172(3):880–903, August 2018. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). See correction [TN21].

Takata:2021:CSK

[TN21] Shigeru Takata and Takashi Noguchi. Correction to: A Simple Kinetic Model for the Phase Transition of the van der Waals Fluid. Journal of Statistical Physics, 185(1):??, October 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02827-w. See [TN18].

Trugilho:2022:MCF

[TR22] L. F. Trugilho and L. G. Rizzi. Microcanonical characterization of first-order phase transitions in a generalized model for aggregation. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02880-z.

Trinh:2021:BJE

[TT21] Hoang Dung Trinh and Khanh Duy Trinh. Beta Jacobi ensembles and associated Jacobi polynomials. Journal of Statistical Physics, 185(1):??, October 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02832-z.

Teufel:2023:TET

[TTV23] Stefan Teufel, Roderich Tumulka, and Cornelia Vogel. Time evolution of typical pure states from a macroscopic Hilbert subspace. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03074-x.

Tong:2022:DLB

[TTZ22] Leilei Tong, Zhong Tan, and Xu Zhang. The diffusive limit of the bipolar Vlasov–Poisson–Boltzmann equations. Journal of Statistical Physics, 188(1):??, July 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-02928-0.

Tang:2023:UCT

[TYM23] Yanjie Tang, Xiaojiang Ye, and Dongkui Ma. The upper capacity topological entropy of free semigroup actions for certain noncompact sets, II. Journal of Statistical Physics, 190(4):??, April 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03083-w.

Tu:2020:SRC

[TZQY20] Zhe Tu, Dazhi Zhao, Fei Qiu, and Tao Yu. Stochastic resonance in coupled underdamped harmonic oscillators with fluctuating frequency driven by dichotomous noise. Journal of Statistical Physics, 179(1):247–262, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Unterberger:2022:ECG

[Unt22] Jérémie Unterberger. Exact computation of growth-rate variance in randomly fluctuating environment. Journal of Statistical Physics, 189(3):??, December 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-022-03002-5.

Urrutia:2022:FVC

[Urr22] Ignacio Urrutia. The fourth virial coefficient for hard spheres in even dimension. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02913-7.

vanderHofstad:2023:SCF

[vdHvdHM23] Remco van der Hofstad, Pim van der Hoorn, and Neeladri Maitra. Scaling of the clustering function in spatial inhomogeneous random graphs. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03122-6.

vanEnter:2020:SGS

[vEKM20] Aernout van Enter, Henna Koivusalo, and Jacek Miȩkisz. Sturmian ground states in classical lattice-gas models. Journal of Statistical Physics, 178(3):832–844, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

vanEnter:2020:TGS

[vEM20] Aernout van Enter and Jacek Miȩkisz. Typical ground states for large sets of interactions. Journal of Statistical Physics, 181(5): 1906–1914, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10955-020-02647-4.pdf.

vanGinkel:2020:HLS

[vGR20] Bart van Ginkel and Frank Redig. Hydrodynamic limit of the symmetric exclusion process on a compact Riemannian manifold. Journal of Statistical Physics, 178(1):75–116, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02420-2.pdf.

vanGinkel:2021:RTM

[vGvGR21] Bart van Ginkel, Bart van Gisbergen, and Frank Redig. Runand-tumble motion: The role of reversibility. Journal of Statistical Physics, 183(3):??, June 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL http://

link.springer.com/article/10.1007/s10955-021-02787-1;
 http://link.springer.com/content/pdf/10.1007/s10955-
021-02787-1.pdf.

Venegas-Li:2023:OCM

[VLC23] Ariadna Venegas-Li and James P. Crutchfield. Optimality and complexity in measured quantum-state stochastic processes. Journal of Statistical Physics, 190(6):??, June 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-023-
03112-8.

Volkov:2022:RUB

[Vol22] Stanislav Volkov. Rigorous upper bound for the discrete Bak– Sneppen model. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02838-7.

Vroylandt:2020:IUR

[VPG20] Hadrien Vroylandt, Karel Proesmans, and Todd R. Gingrich. Isometric uncertainty relations. Journal of Statistical Physics, 178(4):1039–1053, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Verma:2020:RIP

[VSG20] Atul Kumar Verma, Natasha Sharma, and Arvind Kumar Gupta. On the role of interacting particles and limited resources in the regulation of lattice length dynamics. Journal of Statistical Physics, 179(1):216–230, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wand:2022:AFT

[Wan22] Tobias Wand. Analysis of the football transfer market network. Journal of Statistical Physics, 187(3):??, June 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-022-
02919-1.

Watanabe:2019:PBT

[Wat19] Haruki Watanabe. A proof of the Bloch theorem for lattice models. Journal of Statistical Physics, 177(4):717–726, November 2019. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL http://link.springer.com/content/pdf/
10.1007/s10955-019-02386-1.pdf. See correction [Wat20].

Watanabe:2020:CPB

[Wat20] Haruki Watanabe. Correction to: A Proof of the Bloch Theorem for Lattice Models. Journal of Statistical Physics, 178(6):1515, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-

9613 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10955-020-02502-6.pdf. See [Wat19].

Wimsatt:2021:RLS

[WBC21] Gregory W. Wimsatt, Alexander B. Boyd, and James P. Crutchfield. Refining Landauer’s stack: Balancing error and dissipation when erasing information. Journal of Statistical Physics, 183(1):??, April 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-021-02733-1.

Wennberg:2023:LGN

[Wen23] Bernt Wennberg. The Lorentz gas with a nearly periodic distribution of scatterers. Journal of Statistical Physics, 190(7):??, July 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03134-2.

Weiss:2020:NOP

[WFKM^{+}20] Jeffrey B. Weiss, Baylor Fox-Kemper, Dibyendu Mandal, Arin D. Nelson, and R. K. P. Zia. Nonequilibrium oscillations, probability angular momentum, and the climate system. Journal of Statistical Physics, 179(5–6):1010–1027, June 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wiese:2020:SOW

[Wie20] Kay Jörg Wiese. Span observables: “when is a foraging rabbit no longer hungry?”. Journal of Statistical Physics, 178(2):625– 643, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wilkinson:2023:QLD

[Wil23] Michael Wilkinson. Quantifying the lucky droplet model for rainfall. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03056-5.

Wirth:2022:DFN

[Wir22] Melchior Wirth. A dual formula for the noncommutative transport distance. Journal of Statistical Physics, 187(2):??, May 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02911-9.

Wagaskar:2020:SSR

[WLB^{+}20] K. V. Wagaskar, Ravikiran Late, A. G. Banpurkar, A. V. Limaye, and Pradip B. Shelke. Simulation studies of random sequential adsorption (RSA) of mixture of two-component circular discs. Journal of Statistical Physics, 181(6):2191–2205, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Watanabe:2020:GPN

[WLO20] Haruki Watanabe, Yankang Liu, and Masaki Oshikawa. On the general properties of non-linear optical conductivities. Journal of Statistical Physics, 181(6):2050–2070, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02654-5.pdf.

Wojtkowiak:2022:BTD

[WM22] Zbigniew Wojtkowiak and Grzegorz Musiał. The behavior of the three-dimensional Askin–Teller model at the mixed phase region by a new Monte Carlo approach. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02964-w.

Watanabe:2020:PAL

[WOK20] Haruki Watanabe, Masaki Oshikawa, and Tohru Koma. Proof of the absence of long-range temporal orders in Gibbs states.

Journal of Statistical Physics, 178(4):926–935, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-019-02471-5.pdf.

Wilkinson:2021:FDD

[WPK21] Michael Wilkinson, Marc Pradas, and Gerhard Kling. Flooding dynamics of diffusive dispersion in a random potential. Journal of Statistical Physics, 182(3):??, March 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-021-02721-5.pdf.

Wang:2020:MRC

[WQ20] Yue Wang and Hong Qian. Mathematical representation of Clausius’ and Kelvin’s statements of the Second Law and irreversibility. Journal of Statistical Physics, 179(3):808–837, May 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wreszinski:2021:USM

[Wre21] Walter F. Wreszinski. Unstable states in a model of nonrelativistic quantum electrodynamics: Corrections to the Lorentzian distribution. Journal of Statistical Physics, 182(2):??, February 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wreszinski:2022:RTR

[Wre22] Walter F. Wreszinski. A relation of thermodynamic relevance between the superadditivity, concavity and homogeneity properties of real-valued functions. Journal of Statistical Physics, 186(2):??, February 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02872-z.

Wreszinski:2023:TQS

[Wre23] Walter F. Wreszinski. A theory of quantum (statistical) measurement. Journal of Statistical Physics, 190(3):??, March 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03071-0.

Wu:2020:IDR

[Wu20] Xuan Wu. Intermediate disorder regime for half-space directed polymers. Journal of Statistical Physics, 181(6):2372–2403, December 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Wu:2021:LPS

[Wu21] Weisheng Wu. Local pressure of subsets and measures. Journal of Statistical Physics, 185(2):??, November 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL

https://link.springer.com/article/10.1007/s10955-021-
02822-1.

Wei:2022:WCL

[WY22] Ran Wei and Jinjiong Yu. Weak coupling limits for directed polymers in tube environments. Journal of Statistical Physics, 186(3):??, March 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02887-6.

Wang:2020:BAR

[WZN20] Zhiying Wang, Hongli Zhao, and Huifang Nie. Bibliometric analysis of rumor propagation research through Web of science from 1989 to 2019. Journal of Statistical Physics, 178(2):532–551, January 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Xia:2020:LRT

[XTL20] Hui Xia, Gang Tang, and Yueheng Lan. Long-range temporal correlations in kinetic roughening. Journal of Statistical Physics, 178(3):800–813, February 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Xu:2021:NMF

[Xu21] Rongxing Xu. A numerical method to find the optimal thermodynamic cycle in microscopic heat engine. Journal of Statistical Physics, 184(3):??, September 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.

springer.com/article/10.1007/s10955-021-02813-2.

XU:2022:HOD

[XU22] Lu XU. Hydrodynamics for one-dimensional ASEP in contact with a class of reservoirs. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02963-x.

Xu:2020:APT

[XX20] Wenjing Xu and Wei Xu. An averaging principle for the timedependent abstract stochastic evolution equations with infinite delay and Wiener process. Journal of Statistical Physics, 178 (5):1126–1141, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Xue:2021:MDS

[XZ21] Xiaofeng Xue and Linjie Zhao. Moderate deviations for the SSEP with a slow bond. Journal of Statistical Physics, 182(3): ??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Yakir:2021:FLS

[Yak21] Oren Yakir. Fluctuations of linear statistics for Gaussian perturbations of the lattice z^{d}. Journal of Statistical Physics, 182 (3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Yousuf:2020:GGC

[YK20] Muhammad Irfan Yousuf and Suhyun Kim. Generating graphs by creating associative and random links between existing nodes. Journal of Statistical Physics, 179(1):1–32, April 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Young:2020:TMM

[You20] Lai-Sang Young. Towards a mathematical model of the brain. Journal of Statistical Physics, 180(1–6):612–629, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Yang:2021:BNT

[YQ21] Ying-Jen Yang and Hong Qian. Bivectorial nonequilibrium thermodynamics: Cycle affinity, vorticity potential, and Onsager’s principle. Journal of Statistical Physics, 182(3):??, March 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Yang:2020:DML

[YS20] Fang Yang and Xu Sun. On dynamics of the maximum likelihood states in nonequilibrium systems. Journal of Statistical Physics, 181(3):753–760, November 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Yamamoto:2021:DML

[YS21] Keiichi Yamamoto and Takeshi Seta. Derivation of multicomponent lattice Boltzmann equations by introducing a nonequilibrium distribution function into the Maxwell iteration based on the convective scaling. Journal of Statistical Physics, 182(1): ??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Yuan:2020:KMR

[Yua20] Linglong Yuan. Kingman’s model with random mutation probabilities: Convergence and condensation II. Journal of Statistical Physics, 181(3):870–896, November 2020. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (elec-

tronic). URL http://link.springer.com/content/pdf/10.
1007/s10955-020-02609-w.pdf.

Zamparo:2022:RMD

[Zam22] Marco Zamparo. Renewal model for dependent binary sequences. Journal of Statistical Physics, 187(1):??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-02893-8.

Zhou:2022:UFG

[ZCD22] Chi-Chun Zhou, Yu-Zhu Chen, and Wu-Sheng Dai. Unified framework for generalized statistics: Canonical partition function, maximum occupation number, and permutation phase of wave function. Journal of Statistical Physics, 186(1):??, January 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02865-4.

Zhang:2021:LRT

[ZD21] Qi Zhang and Jinqiao Duan. Linear response theory for nonlinear stochastic differential equations with α-stable Lévy noises. Journal of Statistical Physics, 182(2):??, February 2021. CO-DEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zhang:2020:OST

[Zha20] Yunxin Zhang. Optimization of stochastic thermodynamic machines. Journal of Statistical Physics, 178(6):1336–1353, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zhang:2021:QDG

[ZL21] Yue Zhang and Shunlong Luo. Quantifying decoherence of Gaussian noise channels. Journal of Statistical Physics, 183(2):??, May 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-021-02758-6.

Zhou:2022:KDI

[ZL22] Xia Zhou and Shaoyong Lai. A kinetic description of individual wealth growth and control. Journal of Statistical Physics, 188(3):??, September 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02961-z.

Zhu:2021:UCT

[ZM21] Li Zhu and Dongkui Ma. The upper capacity topological entropy of free semigroup actions for certain non-compact sets. Journal of Statistical Physics, 182(1):??, January 2021. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zhao:2020:RTS

[ZRW20] Renjie Zhao, Richard C. Remsing, and John D. Weeks. Response theory for static and dynamic solvation of ionic and dipolar solutes in water. Journal of Statistical Physics, 180(1–6):721–738, September 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zhu:2020:GLE

[ZV20] Yuanran Zhu and Daniele Venturi. Generalized Langevin equations for systems with local interactions. Journal of Statistical Physics, 178(5):1217–1247, March 2020. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic).

Zeng:2022:NEC

[ZW22] Qian Zeng and Jin Wang. Nonequilibrium enhanced classical measurement and estimation. Journal of Statistical Physics, 189(1):??, October 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://link.springer.

com/article/10.1007/s10955-022-02974-8.

Zhang:2023:RSO

[ZW23] Lixia Zhang and Caishi Wang. Random Schrödinger operator on infinite-dimensional hypercube (I): Ergodicity and density of states. Journal of Statistical Physics, 190(8):??, August 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-023-03153-z.

Zhou:2022:LWD

[ZXD22] Tian Zhou, Pengbo Xu, and Weihua Deng. Lévy walk dynamics in an external constant force field in non-static media. Journal of Statistical Physics, 187(1):??, April 2022. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613 (electronic). URL https://

link.springer.com/article/10.1007/s10955-022-02904-8.

Zhou:2023:PSD

[ZXJL23] Xinping Zhou, Jiamin Xing, Xiaomeng Jiang, and Yong Li. Periodic solutions in distribution of mean-field stochastic differential equations. Journal of Statistical Physics, 190(2):??, February 2023. CODEN JSTPSB. ISSN 0022-4715 (print), 1572-9613

(electronic). URL https://link.springer.com/article/10.
1007/s10955-022-03042-x.